### v1

# LoopServe: An Adaptive Dual-phase LLM Inference Acceleration System for Multi-Turn Dialogues

Abstract.
Multi-turn dialogues are essential in many real-world applications of large language models, such as chatbots and virtual assistants. As conversation histories become longer, existing large language models face increasing computational and memory challenges, which hinder their ability to provide efficient and responsive interactions. Most current acceleration methods either compress the context or optimize key value caching, but they often rely on fixed or position-based heuristics that do not adapt well to the dynamic and unpredictable patterns found in actual multi-turn conversations. In this paper, we present LoopServe, an adaptive dual-phase inference acceleration framework for large language models in multi-turn dialogues. LoopServe introduces two main innovations. First, it performs online sparsification during the prefilling phase by dynamically selecting the most important parts of the attention matrix for each new input. Second, it uses progressive key value compression during decoding by adaptively maintaining a relevant and efficient cache based on the most recently generated output tokens. We also propose a new benchmark with eleven multi-turn datasets that reflect realistic query positions and conversational dependencies. Extensive experiments demonstrate that LoopServe consistently achieves superior effectiveness compared to existing baselines and significantly accelerates LLM inference across a wide range of long-context dialogue tasks.
PVLDB Reference Format:
PVLDB, 18(1): XXX-XXX, 2025. doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097. doi:XX.XX/XXX.XX
1. Introduction
Multi-turn dialogues are at the core of numerous real-world applications, from customer service chatbots to virtual assistants and collaborative agents. These scenarios demand that large language models (LLMs) (Hadi et al., 2023; Deng et al., 2025; Li et al., 2024c; Zhou et al., 2024; van Renen et al., 2024) not only generate coherent responses but also maintain contextual consistency across lengthy, evolving conversations. As the number of dialogue turns increases, so does the computational workload. For instance, processing a multi-turn dialogue comprising 10,000 tokens with Llama-3.1-70B (Grattafiori et al., 2024) can demand trillions of floating-point operations (FLOPs), quickly challenging the limits of real-time inference. Despite the remarkable progress of LLMs such as GPT (Brown et al., 2020; Radford et al., 2018, 2019), Llama (Grattafiori et al., 2024; Touvron et al., 2023), and DeepSeek (DeepSeek-AI, 2024, 2025), their inefficiency in handling multi-turn dialogues remains largely unaddressed.
In a typical multi-turn setting, each new user turn expands the conversation history, requiring the LLM to process ever-growing input sequences. The model’s self-attention mechanism, which lies at the heart of the Transformer (Vaswani, 2017), needs to compute pairwise attention scores between every pair of tokens in the input. Specifically, given an input sequence of length $n$, an LLM with $P$ parameters and a hidden dimension $d$, the time complexity of generating $m$ tokens is $\mathcal{O}(m((n+m)^{2}d+P))$. For instance, in a 3-turn conversation with 5000 tokens per turn, the effective context length for the model reaches 15,000 tokens, resulting in quadratic growth in both computational cost and memory usage. This compounding effect of context accumulation makes real-time, cost-efficient inference increasingly difficult as the conversation progresses.
This challenge is distinct from single-turn tasks, where the input context is static and the user query is typically placed at the end of the sequence. In contrast, multi-turn dialogues feature dynamic context accumulation and variable query positions: user questions may appear at the beginning, middle, or end of the input. This variability causes the attention patterns within the model to change unpredictably across turns and inputs, requiring the model to flexibly maintain and retrieve relevant information from any point in the conversation. The resulting attention matrices not only grow in size with each turn but also exhibit highly dynamic and input-dependent sparsity patterns, further exacerbating the inefficiency of existing inference methods.
Recent research on LLM inference acceleration has primarily focused on two directions: context compression and Key-Value (KV) compression. Firstly, Context compression techniques, such as LLMLingua (Pan et al., 2024), CompAct (Yoon et al., 2024), and RAG-based approaches (Zhao et al., 2024; Jiang et al., 2024b; Wang et al., 2025; Chen et al., 2025; Edge et al., 2024; Wang et al., 2025), reduce input length by filtering, summarizing, or structuring the context. While effective in lowering computational costs, these strategies often sacrifice fine-grained details, which can degrade output fidelity. Secondly, KV compression approaches aim to reduce the computational burden of attention weight calculations during both the prefilling and decoding stages. In the prefilling stage, where the attention matrix is computed for all token pairs, methods (Jiang et al., 2024a; Lv et al., 2024; Lai et al., 2025), such as Minference (Jiang et al., 2024a), sparsify the attention matrix to reduce quadratic computation (See Table 2). In the decoding stage, KV caches store precomputed Key and Value vectors for reuse, minimizing redundant computations during autoregressive text generation. For example, H2O (Zhang et al., 2023b), PQCache (Zhang et al., 2025), SnapKV (Li et al., 2024b), and AdaKV (Feng et al., 2024) select or cache only the most relevant tokens based on attention weights or fixed token positions. However, these methods predominantly rely on static or position-based heuristics, which are ill-suited for the dynamic, input-dependent patterns observed in real-world multi-turn dialogues.
Moreover, current evaluation benchmarks (Li et al., 2024c, 2025; Hsieh et al., 2024; Kim et al., 2025; An et al., 2023) for LLM acceleration misrepresent real-world dialogue scenarios. Most benchmarks (Bai et al., 2024a; Li et al., 2025; Dacheng Li* and Zhang, 2023) assume queries are always placed at the end of the input and focus on single-turn tasks, which oversimplifies the problem and favors acceleration methods that exploit positional biases. As a result, approaches (Zhang et al., 2023b; Li et al., 2024b; Feng et al., 2024) that perform well on these benchmarks often fail to generalize to realistic dialogue scenarios where queries may appear at arbitrary positions and contextual dependencies span multiple turns.
To address these challenges, we conduct experiments and summarize three key findings: (1) The sparsity patterns of attention matrices are highly dynamic across inputs, heads, and layers. A small fraction of vertical and slash lines in the attention matrix accounts for most attention weights, but these patterns vary significantly and cannot be effectively captured by static or offline sparsification methods. (2) The position of questions in user queries significantly affects the performance of existing LLM acceleration techniques, which assume queries are always at the end of the input sequence. However, queries often appear in the middle or beginning of the context. (3) Token importance evolves dynamically during multi-turn dialogues as new outputs are generated. Fixed token selection strategies fail to capture these changing dependencies, leading to inefficiencies and reduced conversational quality.
Based on these observations, we propose LoopServe, an adaptive dual-phase LLM inference acceleration framework specifically designed for multi-turn dialogues. LoopServe introduces two key innovations: online prefilling sparsification and progressive KV compression. In the prefilling phase, LoopServe dynamically identifies and selects the most critical components of the attention matrix, specifically the vertical and slash lines that account for the majority of attention weights. Unlike fixed sparsification methods, LoopServe adapts online, ensuring computational efficiency while maintaining high attention fidelity. In the decoding phase, LoopServe employs a progressive KV compression strategy. It dynamically selects and compresses relevant input tokens based on the most recently generated output tokens. This approach ensures the KV cache remains efficient and relevant throughout the decoding process, significantly reducing computational overhead without compromising output quality. Also, we propose a multi-turn long-context benchmark, including 11 datasets. This benchmark captures diverse query positions and multi-turn dependencies, providing a more realistic evaluation framework for multi-turn dialogue scenarios. The contributions of this paper are summarized as follows:
  • We empirically demonstrate that attention patterns and key token positions in multi-turn dialogues are highly dynamic and unpredictable, revealing the limitations of static sparsification and KV selection methods.
  • We introduce LoopServe, a novel dual-phase LLM inference acceleration system that integrates online attention sparsification and progressive KV compression, substantially improving efficiency and robustness for multi-turn dialogue inference.
  • We propose a new benchmark with 11 diverse multi-turn datasets. This benchmark incorporates diverse query positions and multi-turn dependencies, offering a comprehensive evaluation framework for LLMs in realistic long-context conversational tasks.
  • Extensive experiments conducted on 11 long-context multi-turn datasets demonstrate that LoopServe consistently outperforms state-of-the-art LLM acceleration models.
2. Preliminary and Related Work
In this section, we first introduce the preliminary of LLMs in Section 2.1, and then introduce various techniques that enhance the efficiency of long-context LLM inference in Section 2.2. Finally, we analyze existing benchmarks for LLM acceleration in Section 2.3.
2.1. Large Language Models
LLMs, such as GPT (Brown et al., 2020), Llama (Grattafiori et al., 2024), and DeepSeek (DeepSeek-AI, 2024, 2025), demonstrate exceptional abilities in context understanding and reasoning, achieving significant success in various tasks, such as data cleaning (Feuer et al., 2024; Sun et al., 2024; Ni et al., 2024; Zhu et al., 2024c; Eltabakh et al., 2024; Shankar et al., 2024; Ang et al., 2024; Yan et al., 2024), data management (Li et al., 2024e; Zhang et al., 2024; Halevy et al., 2024; Lou et al., 2024; Huang et al., 2024; Le et al., 2024; Tang et al., 2024; Nagrecha and Kumar, 2023), data visualization (Arora et al., 2023; Wu et al., 2024), data security (Li et al., 2024a), recommendation system (Yao et al., 2023; Zheng et al., 2024), text to SQL (van Renen et al., 2024; Fan et al., 2024; Li et al., 2024d; Zhu et al., 2024a; Ren et al., 2024; Luoma and Kumar, 2025; Zhang et al., 2023a), and information extraction (Zhu et al., 2024b; Xue et al., 2024; Singh et al., 2023; Hu et al., 2024; Tang et al., 2024; Dargahi Nobari and Rafiei, 2024; Chen and Koudas, 2024; Omar et al., 2025; Yu et al., 2025; Huo et al., 2024). Their impressive performance stems from extensive training datasets and a transformer architecture, which effectively captures long-range dependencies and enables advanced autoregressive generation.
2.1.1. Transformer Decoder
The effectiveness of Transformer decoders stems from the Multi-Head Self-Attention (MHSA) mechanism, which efficiently captures both local and global dependencies among input tokens. For clarity, we analyze a single representative Transformer block, as the operational principles are consistent across all blocks in the architecture. Given an input token sequence ${X}=[x_{1},x_{2},\cdots,x_{n}]$ with embeddings $\mathbf{X}\in\mathbb{R}^{n\times d}$, the self-attention mechanism computes weighted aggregations over the token sequence, where the weights are derived from inter-token similarity scores. Specifically, the MHSA mechanism generates query vectors $\mathbf{Q}^{i}\in\mathbb{R}^{n\times d_{k}}$, key vectors $\mathbf{K}^{i}\in\mathbb{R}^{n\times d_{k}}$, and value vectors $\mathbf{V}^{i}\in\mathbb{R}^{n\times d_{v}}$ for the $i$-th attention head through learned linear transformations:
(1) $\mathbf{Q}^{i}=\mathbf{X}\mathbf{W}_{Q^{i}},\quad\mathbf{K}^{i}=\mathbf{X}\mathbf{W}_{K^{i}},\quad\mathbf{V}^{i}=\mathbf{X}\mathbf{W}_{V^{i}},$
where $\mathbf{W}_{Q^{i}}\in\mathbb{R}^{d_{x}\times d_{k}}$, $\mathbf{W}_{K_{i}}\in\mathbb{R}^{d_{x}\times d_{k}}$ and $\mathbf{W}_{V_{i}}\in\mathbb{R}^{d_{x}\times d_{v}}$ are learnable parameters. The self-attention operation is subsequently applied to each triple $(\mathbf{Q}^{i},\mathbf{K}^{i},\mathbf{V}^{i})$, yielding the output of the $i$-th attention head $\mathbf{Z}^{i}$ as follows:
(2) $\mathbf{Z}^{i}=\textsf{Attention}(\mathbf{Q}^{i},\mathbf{K}^{i},\mathbf{V}^{i})=\textsf{Softmax}\left(\frac{\mathbf{Q}^{i}(\mathbf{K}^{i})^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}^{i},$
where the scaling factor $\sqrt{d_{k}}$ ensures numerical stability (Vaswani, 2017). To capture diverse semantic relationships, $h$ parallel attention heads process $\mathbf{X}$ concurrently. Then, $\mathbf{Z}=\textsf{Concat}(\mathbf{Z}^{1},\mathbf{Z}^{2},\dots,\mathbf{Z}^{h})\mathbf{W}_{O},$ where Concat is the concatenation operation, and $\mathbf{W}_{O}\in\mathbb{R}^{h\cdot d_{v}\times d_{o}}$ is the trainable projection matrix. Following the self-attention mechanism, the output is passed through a Feed Forward Network, whose result serves as the input to the next Transformer block.
2.1.2. Auto-regressive Generation Mechanism
LLMs utilize an autoregressive generation paradigm, where text is produced sequentially with each token conditioned on the tokens generated before it. This step-by-step approach maintains semantic coherence and ensures contextual relevance across the output. Formally, given an input sequence $X=[x_{1},x_{2},\cdots,x_{n}]$, LLM predicts the subsequent token $x_{n+1}$ by estimating the conditional probability distribution:
(3) $P(x_{n+1}|x_{1},x_{2},\cdots,x_{n})=\textsf{Softmax}(\mathbf{h}_{n}\mathbf{W}_{\text{out}}+\mathbf{b}_{\text{out}}),$
where $\mathbf{h}_{n}\in\mathbb{R}^{d_{h}}$ represents the hidden state encoding for sequence $X$ at step $n$, $\mathbf{W}_{\text{out}}\in\mathbb{R}^{d_{h}\times n_{vocal}}$ denotes the output projection matrix mapping to vocabulary size $n_{vocal}$. The softmax function transforms logits into a probability distribution across the vocabulary space. During generation, the model samples the next token $x_{n+1}$ from the computed probability distribution: $x_{n+1}\sim P(x_{n+1}|x_{1},x_{2},\cdots,x_{n})$. The newly generated token $x_{n+1}$ is appended to the sequence $X$ to obtain resulting $X=[x_{1},\cdots,x_{n},x_{n+1}]$. This process iterates until either a designated end-of-sequence (EOS) token is produced or a predetermined maximum sequence length is reached.
2.2. Efficient Long-Context Inference
The impressive abilities of LLMs across various tasks weaken significantly when handling long contexts. This happens mainly because of basic limits in computation and memory. The drop in performance, especially efficiency, is due to the quadratic complexity of self-attention mechanisms in Equation (2), which require calculating interactions between every pair of tokens in the input sequence. Specifically, for an LLM architecture comprising $L$ layers with $h$ attention heads per layer, processing an input sequence of length $n$ requires computational operations scaling at $O(Lhn^{2})$. This quadratic growth in both computational demand and memory utilization renders the handling of extensive sequences prohibitively expensive (Li et al., 2024c). To overcome these efficiency challenges, researchers have created specialized techniques to speed up long-context LLM processing. These methods typically fall into two main categories: context compression and KV-based optimization.
Table 1. Summary of important notations.
Symbol Definition
$X_{i}$ Input sequence of tokens
$X_{i,j}$ The $j$-turn input of $X_{i}$
$Y_{i}$ Output sequence of tokens
$y_{i,j}$ The $j$-turn output of $X_{i}$
$n_{i},n_{i,j}$ Length of input sequence $X_{i}$ and $X_{i,j}$
$m_{i},m_{i,j}$ Length of output sequence $Y_{i}$ and $y_{i,j}$
$M_{\theta}$ LLM model
$n_{h}$ The total number of attention head of $M_{\theta}$
$\mathbf{Q}^{k}_{i},\mathbf{K}^{k}_{i},\mathbf{V}^{k}_{i}$ Query, Key, and Value matrices
$\mathbf{A}^{k}_{i}$ The $k$-th attention head $X_{i}$
$\mathcal{S}^{k}_{i}$, $\mathcal{V}^{k}_{i}$ Slash and vertical lines of head $\mathbf{A}^{k}_{i}$
$\mathcal{\hat{S}}^{k}_{i},\mathcal{\hat{V}}^{k}_{i}$ Selected slash lines and vertical lines
$n_{d}$ Decoding interval
$B$ Budget for input tokens
$\hat{X}^{k}_{i}$ Selected important tokens for attention head $k$
2.2.1. Context Compression Methods
Context compression strategies aim to reduce effective sequence length by systematically eliminating non-essential content or distilling key information from verbose inputs. These approaches effectively transform lengthy inputs into more computationally manageable representations. Firstly, the filtering-based methods, such as LLMLingua (Jiang et al., 2023), LLMLingua-v2 (Pan et al., 2024), and CompAct (Yoon et al., 2024), selectively identify and preserve high-relevance content, enabling models to process critical information without engaging the entire sequence at maximum resolution. Alternatively, RAG-based techniques (Zhao et al., 2024; Jiang et al., 2024b; Wang et al., 2025; Chen et al., 2025; Edge et al., 2024) employ a two-stage process: initially constructing knowledge graphs through extraction of essential semantic triples from input text, followed by synthesizing this structured knowledge into condensed representations that serve as reduced-dimension inputs for LLM processing. While these compression strategies substantially reduce computational and memory requirements during inference, they inevitably sacrifice some degree of granular detail, potentially compromising the fidelity and precision of LLM-generated outputs.
2.2.2. KV-based Optimization
To mitigate the computational burden, KV-based approaches focus on reducing the number of attention weight calculations during both prefilling and decoding stages, as summarized in Table 2.
Prefilling-stage Optimization. Given a input $X_{i}$ with $n_{i}$ length, the prefilling stage requires $O(n^{2})$ computations to compute each attention matrix across all token pairs, which is a significant computational bottleneck. Recent approaches such as Minference (Jiang et al., 2024a), FlexPrefill (Lai et al., 2025), and CritiPrefill (Lv et al., 2024) address this by selectively discarding negligible attention values, effectively sparsifying each attention head matrix:
$\min\,\left|\mathbf{{A}}^{j}_{i}-\mathbf{\hat{A}}^{j}_{i}\right|,s.t.\ ,\mathbf{\hat{A}}^{j}_{i}=\textsf{Softmax}\left(\frac{\mathbf{Q}^{j}(\mathbf{K}^{j})^{\top}}{\sqrt{d_{k}}}-c(1-\mathbf{M})\right),$
where $\mathbf{M}\in\{0,1\}^{n\times n}$ represents a binary sparsity mask, where $\mathbf{M}[a][b]=1$ indicates that the attention weight $\mathbf{A}^{j}_{i}[a][b]$ should be preserved. The constant $c=10^{8}$ ensures that masked positions ($\mathbf{M}[a][b]=0$) receive negligible attention weight after softmax normalization. This optimization reduces the computational complexity from $O(n^{2})$ to $O(\alpha n^{2})$ per head, where the sparsity factor $\alpha\ll 1$. Specifically, Minference (Jiang et al., 2024a) improves efficiency by selecting key vertical and slash attention patterns or important blocks, while FlexPrefill (Lai et al., 2025) enhances this approach by dynamically adjusting the sparsity rate based on attention recovery rates. In contrast, CritiPrefill (Lv et al., 2024) partitions keys and queries into blocks and simply selects the top-$k$ most important key blocks for computation. However, as shown by our motivational experiments (Section 3), the sparsity and importance patterns in attention matrices are highly dynamic and input-dependent, varying significantly across turns, heads, and layers. Consequently, fixed-pattern selection approaches often fail to achieve robust performance in real multi-turn scenarios.
Decoding-stage Optimization. During autoregressive text generation, LLMs utilize KV caches to store and reuse partial previously computed Key and Value embeddings, thereby reducing computational redundancy. Firstly, token-level selection approaches (Ge et al., 2024; Li et al., 2024b), such as , H2O (Zhang et al., 2023b), SnapKV (Li et al., 2024b), and AdaKV (Feng et al., 2024), aim to select important tokens from inputs before output generation. These methods assume that critical input tokens required for generating outputs can be identified beforehand and that these tokens significantly overlap with those in the last observation window of the inputs. As analyzed in Section 3, these approaches perform well on existing benchmarks like LongBench (Bai et al., 2024a, b), because such benchmarks place user queries at the end of the input sequence. This assumption holds in existing benchmarks, where user queries are always placed at the end of the input sequence. In such settings, token-level selection methods perform well: by focusing on the most recent tokens (which include the user query), these methods can accurately preserve the necessary context for high-quality output, while significantly reducing memory and computation.
However, this benchmark design does not match real-world multi-turn dialogues, where user queries can appear anywhere in the input. When queries are not always at the end, token-level selection methods like SnapKV (Li et al., 2024b) and AdaKV (Feng et al., 2024) often miss the critical tokens needed for accurate output, resulting in lower performance and poor generalization. As shown in Section 3.2, their effectiveness drops significantly when queries are in the middle or at the front of the context. This highlights the need for adaptive, context-aware KV selection methods that can handle dynamic query positions and dependencies in practical dialogue scenarios.
2.3. Long-Context Benchmarks
Recently, various long-context benchmarks (Li et al., 2024c, 2025; Hsieh et al., 2024; Kim et al., 2025; An et al., 2023; Qin et al., 2024) have been proposed to evaluate LLMs, such as NumericBench (Li et al., 2025), LongBench (Bai et al., 2024a, b), and LongEval (Dacheng Li* and Zhang, 2023). However, these benchmarks have two significant limitations that hinder their applicability to real-world scenarios. First, they assume that user queries are always positioned at the end of the input context. This architectural bias simplifies the evaluation process but fails to capture the complexity of real-world interactions, where queries may appear at the beginning, middle, or interspersed throughout the input. As a result, existing KV-compression approaches perform well on these benchmarks because they are optimized for scenarios where the user question is placed at the end of the input. However, these methods struggle to generalize to real-world human-LLM interactions. Second, most of these benchmarks are single-turn, focusing on isolated question-answering or summarization tasks. This approach overlooks the multi-turn nature of real-world conversations, where contextual dependencies across multiple turns are critical (LI et al., 2025).
To address these limitations, we propose a multi-turn benchmark with 11 datasets in Section 4 with diverse query positions and interaction patterns, providing a more realistic evaluation of LLMs for long-context tasks.
Table 2. LLM acceleration model comparisons, following (LI et al., 2025). P and D denote whether the model has optimization in the Prefilling and Decoding phases, respectively. $n$ is the token size of the input, $m$ is the generation token size, and $c$ and $k$ are constants with $c,k\ll n$ and $c,k\ll m$.
Methods P D KV Size Prefilling Decoding
LLMLingua (Pan et al., 2024) $\times$ $O(\alpha n)$ $O(\alpha^{2}n^{2})$ $O(\alpha nm)$
A-shape (Xiao et al., 2024) $\times$ $O(n)$ $O(kn)$ $O(nm)$
Tri-shape (LI et al., 2025) $\times$ $O(n)$ $O(kn)$ $O(nm)$
MInference (Jiang et al., 2024a) $\times$ $O(n)$ $O(kn)$ $O(nm)$
SLLM (Xiao et al., 2024) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
SnapKV (Li et al., 2024b) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
AdaKV (Feng et al., 2024) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
LoopServe $O(k)$ $O(kn)$ $O(k(m-c)+nc)$
3. Motivational Experiments and Insights
To better understand the fundamental challenges of accelerating LLM inference in multi-turn dialogues, we conduct a series of motivational experiments. These experiments, presented in Sections 3.1 and 3.2, provide key observations on the dynamic sparsity of attention patterns and the impact of query position on acceleration effectiveness. Based on these insights, we introduce the LoopServe system in Section 5, which is specifically designed to address these unique challenges in real-world multi-turn dialogue scenarios.
3.1. Key Point 1: Uncertain Attention Patterns
As discussed in Section 2, each attention head matrix is highly sparse. Current approaches (Jiang et al., 2024a; Xiao et al., 2024; Feng et al., 2024; LI et al., 2025; Li et al., 2024b) for accelerating the prefilling and decoding phases often rely on attention matrix sparsification and important KV token selection. However, these methods are built on a strong assumption: the attention sparsification in during the prefilling and important tokens in decoding follow fixed patterns, and their positions can be identified offline. However, the sparsity patterns of attention matrices are dynamic and unpredictable, varying across different inputs, heads, and layers. This variability challenges the effectiveness of fixed-pattern approaches. In this subsection, we first reveal the sparsity of each attention head and then demonstrate that the sparsity patterns vary across different attention heads and input sequences.
(a) All heads of Lamma 3.
(b) Average ratio among all heads.
Figure 1. Attention head sparsity. Only a small fraction of vertical and slash lines in each attention head of LLMs can recover most of the attention weights, revealing strong sparsity in the attention matrix.
3.1.1. Motivational Observation 1: only 90% of vertical lines and slash lines can collectively account for most (e.g., 90%) of the attention weight.
In the attention matrix, a vertical line is a single column, showing that one token is attended to by all other tokens in the sequence. This often happens with special tokens like separators or keywords. A slash line is a diagonal, showing that each token mainly focuses on its nearby tokens, which is a pattern of local attention. We conduct experiments on a SAMSum QA (Bai et al., 2024a) dataset $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{|\mathcal{D}|}$ using the Llama-3.1-8B-Instruct model with $n_{h}$ attention heads. For each user query $X_{i}$ of length $n_{i}$, each $k$-th attention matrix $\mathbf{A}^{k}_{i}\in\mathbb{R}^{n_{i}\times n_{i}}$ contains $n_{i}$ slash lines $\mathcal{S}^{k}_{i}$ and $n_{i}$ vertical lines $\mathcal{V}^{k}_{i}$. The total attention weight for each slash line $s^{k}\in\mathcal{S}^{k}_{i}$ is computed as $\sum_{(a,b)\in s^{k}}\mathbf{A}^{k}_{i}[a][b]$, and the total attention weight for each vertical line $v^{k}\in\mathcal{V}^{k}_{i}$ is $\sum_{(a,b)\in v^{k}}\mathbf{A}^{k}_{i}[a][b]$. Next, we only use the top $\eta\cdot 2n_{i}$ slash lines $\mathcal{\hat{S}}^{k}_{i}\subseteq\mathcal{S}^{k}_{i}$ and vertical lines $\mathcal{\hat{V}}^{k}_{i}\subseteq\mathcal{V}^{k}_{i}$. Therefore, for each $k$-the attention head, the ratio of the attention weight within these selected lines to the total attention weight can be computed as follows: $r^{k}_{i}=\frac{1}{|\mathcal{D}|}\sum_{X_{i}\in\mathcal{D}}\frac{\sum_{(a,b)\in\mathcal{\hat{S}}^{k}_{i}\cup\mathcal{\hat{V}}^{k}_{i}}\mathbf{A}^{k}_{i}[a][b]}{n_{i}}.$ Also, the average ratio among $n_{h}$ attention heads can be computed as $\frac{1}{n_{h}}\sum_{j=1}^{n_{h}}r^{k}_{i}$. As shown in Figure 1 (a), increasing $\eta$ raises the accumulated attention weight across different heads; a small percentage of slash and vertical lines captures most of the attention weight. Figure 1 (b) shows that both LLaMA and Mistral exhibit this sparsity: covering 90% of the total attention weight requires selecting only 10% of slash and vertical lines. This indicates that, due to the sparsity in attention heads, vertical and slash lines are fundamental for efficient selection.
3.1.2. Motivational Observation 2: The positions of the top vertical and slash lines within the same head vary across different user inputs.
Given the LLM model $M_{\theta}$ with $n_{h}$ attention heads, the dataset $\mathcal{D}=\{X_{i}\}_{i=1}^{|\mathcal{D}|}$, and the percentage $\eta$, we select the top-$\eta\cdot 2n_{i}$ slash lines $\mathcal{\hat{S}}^{k}_{i}\subseteq\mathcal{{S}}^{k}_{i}$ and vertical lines $\mathcal{\hat{V}}^{k}_{i}\subseteq\mathcal{V}^{k}_{i}$ for each user query $X_{i}$ under each $k$-th head. For two queries $X_{i}$ and $X_{j}$, we compute the line overlap of the selected vertical and slash lines as: $r^{k}_{i,j}=\frac{|\mathcal{\hat{S}}_{i}^{k}\cap\mathcal{\hat{S}}_{j}^{k}|+|\mathcal{\hat{V}}_{i}^{k}\cap\mathcal{\hat{V}}_{j}^{k}|}{|\mathcal{\hat{S}}_{i}^{k}\cup\mathcal{\hat{S}}_{j}^{k}|+|\mathcal{\hat{V}}_{i}^{k}\cup\mathcal{\hat{V}}_{j}^{k}|}.$ The average important line overlap ratio across $n_{h}$ heads can then be computed as $\frac{1}{n_{h}|\mathcal{D}|^{2}}\sum_{k=1}^{n_{h}}\sum_{X_{i},X_{j}\in\mathcal{D}}r^{k}_{i,j}$. In particular, we use SAMSum QA (Bai et al., 2024a) dataset along with Llama-3.1-8B-Instruct and Mistral-7B-Instruct-v0.3, and vary $\eta\in\{0.1,0.15,0.2,0.25,0.3\}$. As shown in Figure 2, the overlap of important lines across both LLMs is mostly below 0.5 for different inputs in most heads. This shows that different texts produce distinct line patterns, even within the same attention head. Therefore, it is not practical to identify important lines offline and use them directly for online inference.
(a) Llama-3.1-8B-Instruct.
(b) Mistral-7B-Instruct-v0.3.
Figure 2. Line overlap of the same head with different inputs. The important attention lines differ significantly across inputs, even for the same attention head, highlighting the dynamic and input-dependent nature of attention sparsity.
(a) Attention map.
(b) Important line overlap.
Figure 3. Different patterns in the same head. It illustrates that important attention lines differ even within different segments of the same input, emphasizing the locality and unpredictability of attention patterns.
3.1.3. Motivational Observation 3: Given an input $X_{i}=[C_{i}^{1},C^{2}_{i}]$, where $X_{i}$ consists of two sequential segments $C_{i}^{1}$ and $C_{i}^{2}$. The important positions of the top vertical and slash lines within the same head differ between $C^{1}_{i}$ and $C^{2}_{i}$. Both $C^{1}_{i}$ and $C^{2}_{i}$ exhibit their own local attention sparsity patterns.
As shown in Figure 3 (a), the prominent slash and vertical lines in the attention matrix of $C^{1}$ do not appear in $C^{2}$. Instead, the attention part of ${C^{2}}$ exhibits distinct local patterns. To qualitatively verify this, we use SAMSum QA (Bai et al., 2024a) data with Llama-3.1-8B-Instruct to conduct experiments. Given the dataset $\mathcal{D}=\{X_{i}\}_{i=1}^{|\mathcal{D}|}$, each $X_{i}$ is split into two sequential parts as $X_{i}=[C^{1}_{i},C^{2}_{i}]$. We denote the attention matrix of $C^{1}_{i}$ as $\mathbf{A}^{k}_{C^{1}_{i}}\in\mathbb{R}^{|C^{1}_{i}|\times|C^{1}_{i}|}$ and the attention matrix of $C^{2}_{i}$ as $\mathbf{A}^{k}_{C^{2}_{i}}\in\mathbb{R}^{|C^{2}_{i}|\times n_{i}}$. Next, we select the top-$\eta$ slash and vertical lines for $\mathbf{A}^{k}_{C^{1}_{i}}$ and $\mathbf{A}^{k}_{C^{2}_{i}}$, denoted as $L^{k}_{C^{1}_{i}}$ and $L^{k}_{C^{2}_{i}}$, respectively. The overlap rate of importat lines between $C^{1}_{i}$ and $C^{2}_{i}$ as $r^{k}_{C_{i}^{1}\to C_{i}^{2}}={\sum_{l\in L^{k}_{C^{1}}}\mathbb{I}(l\in L^{k}_{C^{2}})}/{|L^{k}_{C^{2}}|},$ where $\mathbb{I}(l\in L^{k}_{C^{2}})=1$ if the line $l\in L^{k}_{C^{2}}$, indicating that an important line $l$ of $C_{i}^{1}$ is still important in $C_{i}^{2}$. Then, given data $\mathcal{D}$ and the $k$-th attention head, the average overlapping ratio is $\frac{1}{\mathcal{D}}\sum_{X_{i}\in\mathcal{D}}{r^{k}_{C_{i}^{1}\to C_{i}^{2}}}$. As shown in Figure 3 (b), under different $\eta$, The overlap rate between contexts in a user query remains low and highly volatile, indicating distinct local attention patterns for different subsections of input $X_{i}$.
This observation directly highlights that it is challenging to use a subsection of the input to determine the patterns of the entire input, as distinct local patterns exist within different subsections. Consequently, existing prefilling acceleration approaches, such as Minference (Jiang et al., 2024a), which rely on the last window (e.g., $C^{2}_{i}$ )to determine the patterns for the whole input $X_{i}$ and select important attention blocks, fail to achieve consistent performance in real-world applications. Similarly, decoding acceleration approaches, such as SnapKV (Li et al., 2024b), H2O (Zhang et al., 2023b), and Keyformer (Adnan et al., 2024), which use the last observation window (e.g., last several tokens of $C_{i}^{1}$) to identify important tokens for generating outputs (e.g., the second context $C_{i}^{2}$), are also unreliable. As a result, these methods struggle to deliver good performance in practical scenarios.
3.2. Key Point 2: Question Position Matters.
Based on the above motivational experiments, both prefilling-based and decoding-phase acceleration approaches, that rely on offline sparse pattern discovery or fixed sparse patterns, should incapable of performing well in real-world use cases. However, the results reported in their papers are often comparable to original LLMs using the full attention matrix without any compression. Why? The reason is that existing long-context benchmarks, such as Longbench (Bai et al., 2024a, b), place each user question $q_{i}$ at the end of the input sequence $X_{i}=[C_{i},q_{i}]$. Then, the LLMs will answer the question $q_{i}$ based on the context $C_{i}$. In this setup, by relying only on the last observation window of the input (i.e., the user questions), existing acceleration approaches identify important tokens in the context that are highly correlated with the user’s specific questions. This partially alleviates the uncertainty caused by distinct local patterns observed in real-world scenarios.
3.2.1. Motivational Observation 4: We reveal that relying solely on the last observation window of the input cannot reliably identify the important tokens from the input $X_{i}$ for generating the output $Y_{i}$.
Existing approaches, such as H2O (Zhang et al., 2023b), SnapKV (Li et al., 2024b), and AdaKV (Feng et al., 2024), propose selecting the top-$B$ most important tokens $\hat{X}_{i}\subseteq X_{i}$ from the input $X_{i}$ before generating the output tokens. Given the input $X_{i}$, we denote the last observation window of $X_{i}$ with length $n_{s}$ as $X^{obs}_{i}=X_{i}[n_{i}-n_{s}:n_{i}]$. Under the LLM model $M_{\theta}$ with $n_{h}$ attention heads, these approaches select the top-$B$ tokens $\hat{X}_{i}\subseteq X_{i}$ where $|\hat{X}_{i}|=B$ as follows:
(4) $\hat{X}_{i}=\arg\max_{\hat{X}_{i}\subseteq X_{i}}\sum_{j=1}^{n_{h}}\sum_{a\in\hat{X}_{i}}\sum_{b\in X^{obs}_{i}}\mathbf{A}^{k}_{i}[a][b],$
Then, given the output $Y_{i}=[x_{n_{i}+1},x_{n_{i}+2},\cdots,x_{n_{i}+n_{o}}]$ corresponding to the input $X_{i}$, the ground-truth top-$B$ important tokens $X^{*}_{i}\subseteq X_{i}$ can be obtained as follows:
(5) $\hat{X}^{*}_{i}=\arg\max_{\hat{X}_{i}\subseteq X_{i}}\sum_{j=1}^{n_{h}}\sum_{a\in\hat{X}_{i}}\sum_{b\in Y_{i}}\mathbf{A}^{k}_{i}[a][b],$
The overlap between $\hat{X}_{i}$ and $\hat{X}^{*}_{i}$ can be computed as $r_{i}=\frac{|\hat{X}_{i}\cap\hat{X}^{*}_{i}|}{B}$ and the average rate regarding the data $\mathcal{D}$ is $\frac{\sum_{X_{i}\in\mathcal{D}}{r_{i}}}{|\mathcal{D}|}$. We set the budget $B=0.1|X_{i}|$, corresponding to 10% of the tokens in each input $X_{i}$. We use Llama-3.1-8B-Instruct and construct the dataset from LongEval’s topic retrieval set (Dacheng Li* and Zhang, 2023), where each instance contains three topic contexts and an instruction (framed as a question) to summarize two of them. For each instance $X_{i}=(C_{i},q_{i})$, we create three variants by placing the instruction or question $q_{i}$ at the beginning, middle, or end of the context $C_{i}$. As shown in Figure 4 (a), the average overlap of important tokens is highest when the question is at the end, but much lower when placed earlier, indicating that relying only on the last observation window misses relevant information unless the question appears last.
(a) Important token overlapping.
(b) Question in different position.
Figure 4. Question position matters. When user queries are not at the end of the input, both the identification of important input tokens and the task performance of KV compression methods decline, indicating their reliance on queries being at the end.
Also, as shown in Figure 4 (b), SnapKV (Li et al., 2024b) and AdaKV (Feng et al., 2024) perform similarly to the original LLM when the question is at the end of the input. However, their performance drops significantly when the question appears in the middle or at the beginning. It is because SnapKV (Li et al., 2024b) and AdaKV (Feng et al., 2024) rely on the last few tokens to select the most important tokens from the input. When the user questions are at the end of inputs, the last few tokens include the question itself, which helps identify tokens related to the question and generate high-quality outputs. In contrast, when the questions are placed in the beginning or middle, the last few tokens are unrelated to the questions and answers, making it difficult to identify the important tokens and leading to performance drop . This shows that current approaches depend heavily on the benchmark’s question placement, which does not reflect real-world input patterns. Consequently, they struggle to generalize to more realistic scenarios.
4. Multi-turn Long-Context Benchmarks
As discussed in Section 2.3, existing benchmarks mainly focus on single-turn tasks and place user queries at the end of the context, failing to reflect the complexity of real-world multi-turn dialogues. To address this, we propose a new multi-turn long-context benchmark that enables more comprehensive and realistic evaluation of LLMs in conversational scenarios, which is released in Huggingface.
Table 3. Multi-turn dataset statistics.
Type Dataset $|D|$Turn Avg Token Metric
QA NQA 500 3 30545.54 F1
Qasper 500 3 5434.41 F1
MFQA-en 500 3 7279.64 F1
HotpotQA 500 3 13847.19 F1
2WikiMQA 500 3 7471.16 F1
Musique 500 3 16587.56 F1
Summary MultiNews 500 2 2376.46 Rouge-L
GovReport 500 2 9324.43 Rouge-L
QMSum 500 2 12780.29 Rouge-L
Few-shot Learning TREC 199 2 2293.99 Accuracy
SAMSUM 199 2 3113.73 Accuracy
4.1. The Design of Multi-turn LongBench
Each $m$-turn long-context data can be denoted as $\mathcal{D}=\{\mathcal{I}_{i}=[(C_{i,1},q_{i,1},a_{i,1}),(C_{i,2},q_{i,2},a_{i,2}),\dots,(C_{i,m},q_{i,m},a_{i,m})]\}_{i=1}^{|\mathcal{D}|}$, where $C_{i,j}$ is the context at the $j$-th turn of the instance $\mathcal{I}_{i}$, which can be empty, $q_{i,j}$ is the corresponding user question, and $a_{i,j}$ denotes the generated answer of the LLMs for $q_{i,j}$. We design diverse formats of for each multi-turn long-context data instance as follows:
  • Diverse Query Positions: For the $j$-th turn, given a context $C_{i,j}=\{C^{1}_{i,j},C^{2}_{i,j},\dots,C^{p}_{i,j}\}$ consisting of $p$ distinct paragraphs (e.g., segments), the query $q_{i,j}$ can be positioned at various locations within $C_{i,j}$. Specifically, it can appear at the beginning of $C_{i,j}$, at the end of $C_{i,j}$, or between two segments $C^{k}_{i,j}$ and $C^{k+1}_{i,j}$. Such a way addresses the limitation of existing methods, which only place the query at the end of the context. This placement may not accurately reflect real-world scenarios.
  • Diverse Query Relevance: At the $j$-th turn, the answer $a_{i,j}$ to question $q_{i,j}$ is derived from the contexts $\{C_{i,j^{\prime}}\}_{j^{\prime}=1}^{j}$. In real user scenarios, the context sources for answering $q_{i,j}$ are diverse. Instead of restricting $q_{i,j}$ to rely solely on $C_{i,j}$, we design the answer $a_{i,j}$ to $q_{i,j}$ to come from a subset of contexts $C_{q_{i}}\subseteq\{C_{i,j^{\prime}}\}_{j^{\prime}=1}^{j}$, with the size of the subset varying as $|C_{q_{i}}|\in\{1,2,\dots,j\}$.
4.2. Multi-turn LongBench Generation
Based on the above format, we design multi-turn long-context benchmarks across various categories to comprehensively evaluate the capabilities of LLMs. The detailed information for each dataset is presented in Table 3, and the specific methodology for constructing our proposed long-context benchmarks is outlined as follows:
  • Question Answering (QA). These datasets are derived from the single-document QA and multi-document QA tasks in LongBench (Bai et al., 2024a, b), with each dataset comprising 500 instances. Each instance is structured into three turns. To construct these instances, we randomly select three question-answer pairs from a dataset in LongBench (Bai et al., 2024a, b) as the foundation for the three turns. The associated contexts are then systematically modified through splitting and recombination, with additional irrelevant contexts incorporated. This meticulous design ensures that each instance satisfies the requirements for diverse query positions and diverse query relevance, as outlined in Section 4.1.
  • Summarization. These datasets are derived from the summarization tasks within LongBench (Bai et al., 2024a, b). Each dataset comprises 500 instances, with each instance consisting of two turns. To enhance the diversity of the input, we randomly selected two instances from the original dataset and segmented their original contexts, subsequently recombining them into two turns. This process was carefully designed to ensure compliance with the diverse query relevance requirement outlined in Section 4.1. Finally, we annotated the source paragraphs for traceability and introduced additional noisy contexts to further enrich the complexity and challenge of the dataset.
  • Few-shot Learning. These datasets are derived from the few-shot learning tasks in LongBench (Bai et al., 2024a, b). To fulfill the requirements of diverse query positions and diverse query relevance as outlined in Section 4.1, we exclude instances containing fewer than four examples. For the remaining eligible instances, the examples are segmented and distributed across the first and second turns. The LLM is tasked with generating an initial response based on the examples provided in the first turn and subsequently refining its response in the second turn using the additional examples. Furthermore, the query is strategically positioned at the beginning, middle, and end of the examples to ensure diversity in query placement. To maintain the semantic integrity and structural completeness of the examples, regular expressions are employed for segmentation.
Input: The $m$-turn input $\mathcal{I}_{i}=\{X_{i,j}\}_{j=1}^{m}$, LLM $M_{\theta}$, threshold $\alpha$, re-selection interval $n_{d}$, and budget $B$
Output: Answers $\{y_{i,j}\}_{j=1}^{m}$
1 $X_{i}\leftarrow\emptyset$
2 for $j=1$ to $m$ do
3       $X_{i}=X_{i}\cup y_{i,j-1}\cup X_{i,j},\ \hat{X}_{i,j}=[y_{i,j-1},X_{i,j}]$
       // Step 1: Parallel Prefilling Line Selection
4       for $k=1$ to $n_{h}$ do
5             $\mathcal{\hat{V}}^{k}_{i,j},\mathcal{\hat{S}}^{k}_{i,j}$ = PrefillingLineSelection($M_{\theta}$, $X_{i}$, $\hat{X}_{i,j}$, $\alpha$)
6
      // Step 2: KV Compression for Decoding
7       =$\mathcal{L}=\{\mathcal{\hat{V}}^{k}_{i,j^{\prime}},\mathcal{\hat{S}}^{k}_{i,j^{\prime}}\}_{j^{\prime}=1,k=1}^{j,n_{h}}$
8
9      $y_{i,j}=\textsf{ProgressiveDecoding}(M_{\theta},{X}_{i},\mathcal{L},n_{d},B)$
10
Return  Answers $\{y_{i,j}\}_{j=1}^{m}$.
Algorithm 1 LoopServe framework overview
5. LoopLLM System
In this section, we first introduce an overview of the LoopServe framework and introduce its two key components: online pre-filling sparsification and progressive KV compression for decoding.
5.1. Framework Overview of LoopLLM System
The LoopServe framework addresses the inefficiencies of existing LLM inference systems in multi-turn dialogue scenarios, where traditional acceleration methods based on sparse attention or static KV compression use fixed patterns that cannot adapt to dynamic input and context changes. These methods often fail to maintain efficiency and conversational quality when handling long contexts or varied query positions. LoopServe overcomes these challenges with an adaptive dual-phase system that applies online attention sparsification in the prefilling phase and progressive KV compression in the decoding phase. It dynamically selects the most important parts of the attention matrix to reduce computation and iteratively compresses the KV cache by using generated output tokens to select relevant input tokens. For an $m$-turn input $\mathcal{I}_{i}=\{X_{i,j}\}_{j=1}^{m}$, where $X_{i,j}$ is the context or query in the $j$-th turn, LoopServe generates each answer $y_{i,j}$ using these two steps.
Step 1. Online Attention Head Sparsification in Prefilling. To reduce the quadratic complexity of the attention matrix, we dynamically select the most important slash lines and vertical lines that capture the key attention weights. Specifically, in Algorithm 1 line 3-5, for each new input $X_{i,j}$ and the all inputs $X_{i}=\left(\cup_{j^{\prime}=1}^{j-1}{(X_{i,j^{\prime}}\cup y_{i,j^{\prime}})}\right)\cup X_{i,j}$, we get $\hat{X}_{i,j}=[y_{i,j-1},X_{i,j}]$ as the new appended input in the $j$ turn. Then, for each $k$-th attention head, we first compute the attention matrix $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]\in\mathbb{R}^{\hat{n}_{i,j}\times n_{i}}$, and then select the slash lines $\mathcal{\hat{S}}^{k}_{i,j}$ and vertical lines $\mathcal{\hat{V}}^{k}_{i,j}$ that collectively recover at least $\alpha$ of the total attention weight in $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$.
Step 2. Progressive KV Compression in Decoding. To address the high complexity of decoding, LoopServe progressively selects and stores only the most relevant tokens needed for generating subsequent outputs. Specifically, as described in Algorithm 1 (lines 7–14), after every re-selection interval $n_{d}$ tokens, the framework uses the ProgressiveSelection function to compute a subset of input tokens $\hat{X}^{k}_{i}\subseteq X_{i}$ for each attention head $k$. these selected tokens $\hat{X}^{k}_{i}$ are important for output generation. At each decoding step, LoopServe leverages the compressed KV cache $\{\mathcal{\hat{S}}^{k}_{i,j^{\prime}},\mathcal{\hat{V}}^{k}_{i,j^{\prime}}\}_{j^{\prime}=1,k}^{j,n_{h}}$ to generate the output sequence $y_{i,j}$.
5.2. Online Attention Sparsification in Prefilling
As discussed in Section 2.1, the prefilling phase incurs quadratic time complexity in the input length $n$ because each attention matrix requires $O(n^{2})$ computation. Sparsifying the attention matrix is therefore crucial for efficiency. However, as shown in Key Point 1 in Section 3.1, attention matrix sparsity patterns are highly dynamic across inputs and even within the same attention head, making static or offline selection ineffective. To address this, we propose an online adaptive algorithm that, during prefilling, identifies and selects a subset of slash and vertical lines to recover at least an $\alpha$ fraction of the total attention weight for each head. This approach not only reduces computation but also enables efficient reuse of selected lines in subsequent dialogue turns, improving overall efficiency in multi-turn scenarios.
Definition 5.1 (Online Prefilling Sparsification Problem).
Given the input sequence $X_{i}=\left(\bigcup_{j^{\prime}=1}^{j-1}(X_{i,j^{\prime}}\cup y_{i,j^{\prime}})\right)\cup X_{i,j}$, where $y_{i,j^{\prime}}$ is the answer for $X_{i,j^{\prime}}$ and $X_{i,j}$ is the current turn’s input, consider the LLM model $M_{\theta}$ with $n_{h}$ attention heads. We denote the concatenation of the previous answer $y_{i,j-1}$ and the current user input $X_{i,j}$ as $\hat{X}_{i,j}=[y_{i,j-1},X_{i,j}]$, whose corresponding attention matrix requires sparsification. The $k$-th attention matrix between $X_{i}$ and $\hat{X}_{i,j}$ is denoted as $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]\in\mathbb{R}^{\hat{n}_{i,j}\times n_{i}}$. Let $\mathcal{S}^{k}_{i,j}$ (resp., $\mathcal{V}^{k}_{i,j}$ ) denote the set of all slash lines (resp., vertical lines) in $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$. The goal is to select a subset of slash lines $\hat{\mathcal{S}}^{k}_{i,j}\subseteq\mathcal{S}^{k}_{i,j}$ and a subset of vertical lines $\hat{\mathcal{V}}^{k}_{i,j}\subseteq\mathcal{V}^{k}_{i,j}$ such that together they recover at least an $\alpha$ fraction of the total attention weight in $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$, where $\alpha\in[0,1]$.
(6) $\min\sum_{s\in\mathcal{\hat{S}}^{k}_{i,j}}l_{s}+\sum_{v\in\mathcal{\hat{V}}^{k}_{i,j}}l_{v},$
$s.t.\quad$ $\sum_{(a,b)\in(\mathcal{\hat{S}}^{k}_{i,j}\cup\mathcal{\hat{V}}^{k}_{i,j})}\mathbf{A}^{k}_{i}[a][b]\geq\alpha\cdot\hat{n}_{i,j},$
where $\hat{n}_{i,j}$ is the total attention weight in the matrix $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$, and $l_{s}$ (resp., $l_{v}$) is the length of the slash line $s$ (resp., vertical line $v$).
Theorem 5.2.
The prefilling sparsification problem is NP-hard.
Proof.
The Online Prefilling Sparsification Problem (OPSP) can be proven NP-hard via a reduction from the Set Cover Problem, a well-known NP-hard problem. The Set Cover Problem is defined as follows: given a universe $U=\{u_{1},u_{2},\dots,u_{m}\}$, a collection of subsets $\mathcal{P}=\{P_{1},P_{2},\dots,P_{n}\}$, the objective is to find a subset $\mathcal{P}^{*}\subseteq\mathcal{P}$ such that $\bigcup_{P_{i}\in\mathcal{P}^{*}}P_{i}=U$, and the total cost $\sum_{P_{i}\in\mathcal{P}^{*}}|P_{i}|$ is minimized. We map the elements of the Set Cover Problem to the OPSP as follows. Each element $u\in U$ corresponds to an entry in the attention matrix $\mathbf{A}_{i}^{k}[\hat{X}_{i,j}]$ that needs to be covered. Each subset $P_{i}$ in $\mathcal{P}$ corresponds to a slash line $s$ or a vertical line $v$ in OPSP, which covers a subset of entries in the matrix. The cost of selecting subset $P_{i}$ is mapped to the cost $l_{s}$ (for slash lines $s$) or $l_{v}$ (for vertical lines $v$) in OPSP. The Set Cover Problem’s requirement to cover all elements in $U$ is equivalent to requiring $\alpha$ in OPSP. Therefore, if we can solve the OSOP optimally in polynomial time, we can solve the set cover problem in polynomial time. ∎
Input: The input $X_{i}$ and $\hat{X}_{i,j}$, $k$-th head of LLM $\mathcal{M}_{\theta}$, the parameter $\alpha$
Output: The selected slash lines $\mathcal{\hat{S}}_{i,j}^{k}$ and vectical lines $\mathcal{\hat{V}}_{i,j}^{k}$
1
[…]
3
4Compute Query $\mathbf{\tilde{Q}}^{k}_{i,j}$ for $\tilde{X}_{i,j}$
5
6Compute Key $\mathbf{{K}}^{k}_{i}$ for ${X}_{i}$
7
[…]
9
[…]
11
[…]
13
[…]
15 $ol_{s}=0,ol_{v}=0,sum=0$
16 while $sum<\alpha\cdot\textsf{Sum}(\mathbf{{A}}^{k}_{i}[\tilde{X}_{i,j}])$ do
17
18      $s=\mathcal{{S}}^{k}_{i,j}[0],v=\mathcal{V}^{k}_{i,j}[0]$
19
20      $\triangle w_{s}=w_{s}-ol_{v},\triangle w_{v}=w_{v}-ol_{s}$
21       if $\triangle w_{s}/(l_{s}-|\mathcal{\hat{V}}^{k}_{i,j}|)\geq\triangle w_{v}/(l_{v}-|\mathcal{\hat{S}}^{k}_{i,j}|)$ then
22             $\mathcal{\hat{S}}^{k}_{i,j}=\mathcal{\hat{S}}^{k}_{i,j}\cup\{s\}$, $ol_{s}=ol_{s}+w^{max}_{s}$
23             $sum=sum+w_{s}-ol_{v}$
24
25      else
26             $\mathcal{\hat{V}}^{k}_{i,j}=\mathcal{\hat{V}}^{k}_{i,j}\cup\{v\}$, $ol_{v}=ol_{v}+w^{max}_{v}$
27             $sum=sum+w_{v}-ol_{s}$
28
29
30
Return  $\mathcal{\hat{S}}^{k}_{i,j}$ and $\mathcal{\hat{V}}^{k}_{i,j}$.
Algorithm 2 Adaptive Prefilling Sparsification Framework
Algorithm. Algorithm 2 takes as input the concatenated sequence $\hat{X}_{i,j}=[y_{i,j-1},X_{i,j}]$, where $y_{i,j-1}$ is the previous answer and $X_{i,j}$ is the current user input. It also requires the $k$-th attention head of the LLM $M_{\theta}$ and a sparsity threshold parameter $\alpha$. The output consists of selected slash lines $\mathcal{\hat{S}}^{k}_{i,j}$ and selected vertical lines $\mathcal{\hat{V}}^{k}_{i,j}$. The first step involves sampling a subset $\tilde{X}_{i,j}$ of the concatenated input $\hat{X}_{i,j}$, which reduces the computational cost of subsequent operations. Next, Algorithm 2 computes the query matrix $\mathbf{\tilde{Q}}^{k}_{i,j}$ for the sampled input $\tilde{X}_{i,j}$ and the key matrix $\mathbf{K}^{k}_{i}$ for the entire input $X_{i}$. Next, all slash lines $\mathcal{S}^{k}_{i,j}$ and vertical lines $\mathcal{V}^{k}_{i,j}$ are summarized based on $\mathbf{A}^{k}_{i}[\tilde{X}_{i,j}]$ and sorted in descending order. Two empty sets, $\mathcal{\hat{S}}^{k}_{i,j}$ and $\mathcal{\hat{V}}^{k}_{i,j}$, are initialized to store the selected lines, and overlap weights $ol_{s}$ and $ol_{v}$ are set to zero. Then, Algorithm 2 iteratively selects lines until the total recovered attention weight sum meets or exceeds $\alpha\cdot\textsf{sum}(\mathbf{A}^{k}_{i}[\tilde{X}_{i,j}])$. At each step, the top slash line $s\in\mathcal{S}^{k}_{i,j}$ and the top vertical line $v\in\mathcal{V}^{k}_{i,j}$ are compared based on their marginal contributions, $\Delta w_{s}=w_{s}-ol_{v}$ and $\Delta w_{v}=w_{v}-ol_{s}$. Particularly, since one slash line only has one overlap point with one vertical line, and thus $\Delta w_{s}=w_{s}-ol_{v}$. The line with the greater contribution is added to its respective set ($\mathcal{\hat{S}}^{k}_{i,j}$ or $\mathcal{\hat{V}}^{k}_{i,j}$), and the overlap weights and total recovered weight are updated accordingly. The loop ends once the recovery condition is satisfied, and the algorithm returns the sets $\mathcal{\hat{S}}^{k}_{i,j}$ or $\mathcal{\hat{V}}^{k}_{i,j}$.
Time Complexity Analysis. It is $O(n_{i}|\tilde{X}_{i,j}|+n_{i}logn_{i}+n_{i})$. Firstly, it takes $O(n_{i}|\tilde{X}_{i,j}|)$ to compute the partial attention matrix $\mathbf{{A}}^{k}_{i}[\tilde{X}_{i,j}]$. Then, it takes $O(n_{i}|\tilde{X}_{i,j}|)$ to summarize the values for each slash line and vertical line, and takes $O(n_{i}\log n_{i})$ for descending sorts. Finally, the greedy selection loop runs in $O(n_{i})$.
5.3. Progressive KV Compression in Decoding
As discussed in Section 2.1.1, conventional decoding strategies often use fixed patterns or focus on the last few input tokens to select important information. However, as demonstrated in Key Point 2 in Section 3.2, these methods are ineffective when user queries appear earlier in the input sequence. To address this, we propose a progressive KV compression strategy that dynamically selects important input tokens during decoding based on the most recently generated output tokens. This ensures that the selected tokens remain relevant throughout the generation process. Our analysis shows that using recent output tokens to guide selection leads to greater overlap with truly important input tokens, enhancing both efficiency and accuracy.
To verify this, we consider the LLM $M_{\theta}$ with $n_{h}$ attention heads, the input sequence $X_{i}$, the generated output sequence $Y_{i}$, and a window size of $n_{w}$. Based on Equation (5), we compute the ground truth top-$B$ important tokens $\hat{X}^{*}_{i}\subseteq X_{i}$ using the output $Y_{i}$. We then use an observation window extracted from either $X_{i}$ or $Y_{i}$ to select the top-$B$ important tokens. Specifically, given the window size $n_{w}$, the $-\frac{|X_{i}|}{n_{w}}$-th to -1 observation windows from the input $X_{i}$ are defined as $X_{i}[n_{i}-m\cdot n_{w}:n_{i}-(m-1)\cdot n_{w}]$, $X_{i}[n_{i}-(m-1)\cdot n_{w}:n_{i}-(m-2)\cdot n_{w}]$, …, $X[n_{i}-n_{w}:n_{i}]$, and the 0-th to $\frac{|Y_{i}|}{n_{w}}$ observation windows from the output $Y_{i}$ are defined as $Y_{i}[0:n_{w}]$, $Y_{i}[n_{w}:2n_{w}]$, $Y_{i}[(m-1)\cdot n_{w}:m\cdot n_{w}]$. For each observation window $X^{obs}_{i}$ from $X_{i}$ or $Y_{i}$, we select the top-$B$ tokens $\hat{X}_{i}\in X_{i}$ following Equation (4) and compute the overlap rate between $\hat{X}_{i}$ and the ground truth $\hat{X}^{*}_{i}$ as $\frac{|\hat{X}_{i}\cap\hat{X}^{*}_{i}|}{B}$. We use the same LongEval’s topic retrieval set (Dacheng Li* and Zhang, 2023) in Section 3 and Llama to conduct experiments, and set $B=\{5\%,10\%,15\%\}\cdot|X_{i}|$. As shown in Figure 5 (a), the overlap rate between the important tokens selected based on input block windows (block index ¡ 0) is low. In contrast, the overlap rate is significantly higher when important tokens are selected based on output block windows (block index $\geq$ 0). It demonstrates that leveraging output tokens to identify important input tokens is a more effective strategy for selecting relevant tokens.
(a) Important token overlapping.
(b) Decoding blocks.
Figure 5. Progressive decoding. (a) It shows that selecting important tokens from output windows yields higher overlap than from input windows, supporting progressive KV selection. (b) Also, the overlap between selected tokens in output blocks decreases with block distance, highlighting the need for progressive, output-driven cache updates.
Input: Input $\mathcal{I}_{i}$, LLM $M_{\theta}$, all selected slash and vertical lines $\{\mathcal{\hat{V}}^{k}_{i,j^{\prime}},\mathcal{\hat{S}}^{k}_{i,j^{\prime}}\}_{j^{\prime}=1,k=1}^{j,n_{h}}$, re-selection interval $n_{d}$, and budget $B$
Output: Answer $y_{i,j}$
1 $n_{o}=0,y_{i,j}=\emptyset$
2 while LLM generation is not finshed do
3       if $n_{o}=16$ or $(n_{o}-16)\%n_{d}=0$ then
4             ${X}^{obs}_{i}=X_{i}[|X_{i}|-n_{d}:|X_{i}|]$
[…]
6
7      $n_{o}=n_{o}+1$
8       $x_{n_{i}+n_{o}}=\textsf{Decoding}(M_{\theta},\{\hat{X}^{k}_{i}\}_{k=1}^{n_{h}},\{\mathcal{\hat{V}}^{k}_{i,j^{\prime}},\mathcal{\hat{S}}^{k}_{i,j^{\prime}}\}_{j^{\prime}=1,k=1}^{j,n_{h}})$
9
10      $\hat{X}_{i}=\hat{X}_{i}\cup x_{n_{i}+n_{o}}$, ${X}_{i}={X}_{i}\cup x_{n_{i}+n_{o}}$
11
12      $y_{i,j}=y_{i,j}\cup x_{n_{i}+n_{o}}$
13
Return  Answer $y_{i,j}$
Algorithm 3 Progressive Decoding
We also observe that the overlap of important input tokens between output blocks decreases as the distance between the blocks increases. Specifically, given the output sequence $Y_{i}$, we divide it into $n_{b}$ blocks, i.e., $Y_{i}=[Y^{1}_{i},\cdots,Y^{n_{b}}_{i}]$, where $Y^{j}_{i}$ denotes the $j$-th output block with a size of $\frac{|Y_{i}|}{n_{b}}$. For each block $Y^{j}_{i}$, we compute the top-$B$ important input tokens based on the attention weights as $\hat{X}_{i,Y^{j}_{i}}=\arg\max_{\hat{X}_{i}\subseteq X_{i}}\sum_{k=1}^{n_{h}}\sum_{a\in\hat{X}_{i}}\sum_{b\in Y^{j}_{i}}\mathbf{A}^{k}_{i}[a][b]$, Next, we compare the overlap ${|\hat{X}_{i,Y^{j}_{i}}\bigcap\hat{X}_{i,Y^{j^{\prime}}_{i}}|}/{B}$ of important input tokens between every pair of output blocks $Y^{j}_{i}$ and $Y^{j^{\prime}}_{i}$. As shown in Figure 5 (b), for each block $Y^{j}_{i}$, the overlap of important input tokens between $Y^{j}_{i}$ and $Y^{j^{\prime}}_{i}$ (where $j^{\prime}>j$) gradually decreases as the block distance increases. Additionally, the overlap between $Y^{j}_{i}$ and its immediate successor $Y^{j+1}_{i}$ is higher compared to earlier blocks, such as $Y^{j-2}_{i}$. This indicates that tokens identified from $Y^{j}_{i}$ are highly relevant for the generation of $Y^{j+1}_{i}$. Therefore, when generating the output block $Y^{j+1}_{i}$, we leverage the preceding output block $Y^{j}_{i}$ to dynamically identify the top-$B$ important input tokens required for the generation of $Y^{j+1}_{i}$. This progressive approach ensures that the token selection remains relevant and efficient, reducing computational overhead while maintaining high accuracy.
Algorithm. Algorithm 3 firstly set the answer $y_{i,j}$ as empty and decoding step counter $n_{o}$ to 0 (line 1). During the decoding loop, if the current step reaches either the empirically predefined decoding size 16 or a re-selection interval $n_{d}$ (line 3), the algorithm extracts the most recent tokens from the input sequence $X_{i}$, i.e., $X^{obs}_{i}=X_{i}[|X_{i}|-n_{d}:|X_{i}|]$), and dynamically updates the compressed KV cache subset $\hat{X}^{k}_{i}$ for each $k$-th attention head (line 4-5). Specifically, for each head, we select top-$B$ tokens for each head as follows.
(7) $\hat{X}^{k}_{i}=\arg\max_{\hat{X}^{k}_{i}\subseteq X_{i},|\hat{X}^{k}_{i}|=B|}\sum_{a\in\hat{X}_{i}}\sum_{b\in X^{obs}_{i}}\mathbf{A}^{k}_{i}[a][b],$
The LLM generates the next token using the compressed KV cache and the updated input, which is appended to both the input and output sequences (line 6-9). This process iterates until the LLM completes generation, returning the final answer $y_{i,j}$.
6. EXPERIMENTS
Table 4. Effectiveness Evaluation. The bold number indicates the best performance among the baselines.
P Data Llama 3.1 Qwen2.5
Base Snap Ada SLLM A-S T-S Minf Ours Base Snap Ada SLLM A-S T-S Minf Ours
Begin QA MFQA-en 45.70 36.50 34.94 25.83 28.00 34.96 43.05 46.82 44.24 35.59 33.00 21.45 29.56 38.14 42.63 43.47
2WikiMQA 31.68 29.31 28.73 22.05 19.42 24.20 28.57 35.11 37.86 34.54 33.83 23.19 22.11 32.90 35.33 37.52
Musique 17.16 14.29 14.80 9.76 4.22 8.18 14.92 18.81 18.22 14.74 13.95 8.68 5.92 12.15 15.15 16.63
HotpotQA 36.67 33.87 32.98 22.52 12.72 21.29 34.38 39.00 42.93 38.90 37.56 21.42 17.22 31.03 38.67 42.50
NrtvQA 13.40 13.42 8.67 5.22 7.61 9.19 13.47 14.03 13.63 11.05 10.18 6.82 7.57 10.45 10.78 14.36
Qasper 25.33 20.77 17.03 16.58 20.64 23.37 22.79 24.64 26.08 21.79 20.13 17.96 19.67 23.44 24.92 25.37
SUM MultiNews 20.08 19.20 18.03 19.30 20.10 20.15 20.25 20.14 18.39 15.87 14.52 17.29 18.29 18.30 18.30 20.14
GovReport 25.25 18.36 16.53 17.98 24.16 24.13 24.55 24.90 20.93 16.80 15.86 15.90 22.28 21.90 20.86 23.60
QMSum 20.53 17.49 17.56 14.66 17.77 18.92 20.15 20.65 19.97 17.54 17.39 14.66 18.45 19.12 19.61 19.66
FS TREC 46.99 46.74 45.23 46.23 41.21 46.99 45.48 47.99 65.08 64.07 63.32 62.65 60.30 56.28 64.07 65.33
SAMSUM 17.32 16.75 16.48 12.95 17.39 17.35 17.46 18.12 16.19 13.97 13.15 11.22 15.32 15.92 15.96 17.15
Middle QA MFQA-en 46.73 37.30 35.19 26.87 28.26 33.39 43.84 44.73 44.00 34.21 31.68 22.73 29.06 34.31 41.42 41.95
2WikiMQA 34.10 31.71 29.90 25.59 18.77 27.26 32.50 34.25 27.80 22.32 22.69 14.19 20.18 26.15 25.40 27.01
Musique 16.30 13.68 14.05 8.59 3.17 9.51 15.39 17.48 10.05 7.37 7.15 3.32 4.92 8.38 8.76 9.13
HotpotQA 40.63 37.48 36.30 27.30 12.36 25.25 36.81 41.25 29.43 25.24 24.56 12.25 12.96 22.00 26.43 29.05
NrtvQA 15.29 13.06 10.64 7.26 7.14 10.10 13.98 15.25 14.82 11.88 11.83 9.04 7.68 11.24 11.75 15.21
Qasper 30.79 26.80 22.07 21.40 24.90 28.86 28.47 31.12 28.74 23.57 21.77 20.75 24.73 27.84 28.72 29.27
SUM MultiNews 20.59 19.78 18.12 19.91 20.49 20.52 20.79 20.66 18.37 15.54 14.28 17.53 18.09 18.26 18.19 18.44
GovReport 24.08 18.39 15.92 18.09 23.50 24.01 23.74 22.88 20.54 16.24 15.25 15.99 21.00 21.55 20.70 20.68
QMSum 20.51 17.90 17.84 15.08 17.62 17.93 20.20 20.41 20.04 17.79 17.29 15.06 18.57 18.67 19.46 19.81
FS TREC 50.00 50.00 48.74 50.25 50.00 50.76 52.27 56.03 64.07 62.06 60.81 63.32 63.82 40.71 64.07 65.33
SAMSUM 10.62 12.03 11.83 13.34 11.10 10.92 10.62 17.43 12.38 12.55 13.28 13.83 15.65 12.95 12.93 12.40
End QA MFQA-en 50.93 47.93 48.38 32.52 28.62 51.40 49.67 51.69 48.82 47.57 47.24 29.28 27.66 47.94 49.18 48.67
2WikiMQA 42.43 42.34 41.96 37.20 25.19 39.54 41.75 44.05 42.70 42.25 41.24 32.89 26.50 37.54 41.34 42.24
Musique 29.39 28.07 29.01 20.96 7.80 26.44 23.56 31.60 24.18 23.08 22.39 11.82 8.68 17.65 24.34 24.82
HotpotQA 53.62 52.38 53.59 43.83 25.04 51.71 51.97 55.05 53.09 51.03 50.71 35.64 24.63 43.67 53.01 53.13
NrtvQA 25.76 25.50 24.58 19.28 14.19 23.90 23.87 25.87 19.07 17.90 16.75 14.08 11.41 14.39 18.37 19.86
Qasper 37.97 35.95 33.98 28.01 27.37 38.67 36.50 38.27 33.55 31.82 30.21 24.28 25.15 33.72 34.18 33.18
SUM MultiNews 20.59 20.03 18.87 19.97 20.16 20.40 20.49 20.45 18.31 16.61 15.32 17.74 17.99 18.01 18.36 18.40
GovReport 23.90 20.10 18.54 18.04 23.17 23.35 23.92 24.27 21.21 18.06 17.09 16.95 21.53 21.82 21.28 21.10
QMSum 22.69 22.21 22.12 19.67 18.59 22.28 22.27 22.82 21.17 20.15 20.03 18.13 18.07 20.33 21.04 20.95
FS TREC 59.05 58.54 58.54 58.54 52.51 59.55 59.05 60.31 68.34 66.08 66.59 67.84 63.82 67.59 68.85 68.34
SAMSUM 18.78 23.88 20.63 19.84 18.45 17.75 17.62 23.53 39.46 39.58 38.77 38.71 39.13 39.20 39.57 39.85
We compare our LoopServe with six state-of-the-art KV cache algorithms on two representation LLM base models.
6.1. Experimental Settings
6.1.1. Datasets
We designed the multi-turn long-context benchmark to rigorously evaluate LLMs in realistic conversational scenarios, addressing the limitations of single-turn benchmarks. Each instance comprises multiple rounds of context, queries, and answers, introducing diverse query positions and dependencies that reflect real-world conversational complexity. The benchmark covers three key categories: Question Answering, Summarization, and Few-shot Learning, using carefully curated datasets to ensure both diversity and complexity. Detailed dataset statistics are in Table 3.
6.1.2. Baselines
In this study, we evaluate the performance of our proposed LoopServe algorithm against six competitive and state-of-the-art KV cache algorithms on Llama-3.1-8B-Instruct (Grattafiori et al., 2024) and Qwen2.5-7B-Instruct (Team, 2024), as well as the base model without any KV cache acceleration. The following KV cache methods are utilized for comparison:
  • SnapKV (Snap)  (Li et al., 2024b) optimizes KV caching through a snapshot mechanism, focusing on reducing redundant computations and enhancing memory utilization.
  • AdaKV (Ada) (Feng et al., 2024) dynamically reallocates cache budgets across attention heads, shifting resources from heads with concentrated attention to those with dispersed patterns.
  • StreamingLLM (SLLM)  (Xiao et al., 2024) maintains both attention sinks and a rolling window of recent tokens in its KV cache to enable stable infinite-length processing.
  • A-shape (A-S) (Xiao et al., 2024) sparses attention that resembles an A shape.
  • Tri-Shape (T-S)  (LI et al., 2025) creates a triangular sparse attention pattern for pre-filling.
  • Minference (Minf) (Jiang et al., 2024a) employs a minimal inference caching strategy to reduce computational overhead while maintaining effective KV cache management.
6.1.3. Evaluation Metrics.
We adopt evaluation metrics for each dataset according to its task type. The metrics for each dataset are listed in Table 3. The details of the three metrics are as follows.
  • F1: It is harmonic mean of precision and recall.
  • Accuracy: Proportion of correct predictions over all predictions:
[…]
6.1.4. Hyperparameter and Hardware Setting
All codes are executed on a Rocky Linux 8.10 machine with an 8-core Intel® Xeon® Gold 6542Y CPU, an NVIDIA H100 GPU with 80GB of memory, and 256GB of RAM. For baselines, we use their suggested default setting. For the main experiments in Section 6.2, for our LoopServe, we set $\alpha=0.955$, and $n_{d}=16$ as defaults, and we set the token selection budget $B=1024$ for all baselines and LoopServe.
6.2. Main Experiments
6.2.1. Effectiveness Evaluation
To evaluate the effectiveness of our proposed LoopServe and baselines, we conduct experiments on the proposed multi-turn long-context benchmark across three categories: Question Answering, Summarization, and Few-shot Learning. For each dataset, we compared LoopServe with six state-of-the-art KV cache acceleration baselines as well as the base LLMs, using the appropriate metrics (F1, Rouge-L, and Accuracy) for each task. As shown in Table 4, LoopServe achieves the best or comparable performance among KV compression baselines across a wide range of datasets and query positions. A key finding is that LoopServe maintains robust performance regardless of where the user query appears in the beginning, middle, or end of the input context. In contrast, existing baselines such as SnapKV and AdaKV exhibit a significant drop in performance when the query is not positioned at the end, highlighting their reliance on positional heuristics which do not generalize well to realistic dialogue settings. LoopServe’s adaptive approach, which dynamically selects relevant information both during prefilling and decoding, consistently leads to higher accuracy and quality, even as conversational context grows longer and more complex. Notably, these improvements are observed across both Llama-3.1 and Qwen2.5 models, indicating that LoopServe’s design generalizes well across different LLM architectures.
6.2.2. Efficiency Evaluation
In addition to effectiveness, we also evaluate the efficiency of LoopServe in terms of prefilling time and decoding latency. As shown in Figure 6, LoopServe achieves substantial reductions in both prefilling time and inference time compared to the base LLM. The combination of efficient online sparsification and adaptive KV compression enables LoopServe to provide both high-quality outputs and rapid response times. Specifically, LoopServe’s online prefilling sparsification mechanism significantly reduces computational overhead by identifying and selecting only the most critical components of the attention matrix. This targeted sparsification not only accelerates the prefilling stage but also reduces the memory footprint. During the decoding phase, LoopServe’s progressive KV compression dynamically maintains a compact and relevant KV cache, minimizing redundant computations and further improving inference speed.
(a) Prefilling time (time to first token). Higher is better.
(b) Decoding latency (time between tokens). Lower is better.
Figure 6. Efficiency experiment. Base uses naive attention, and experiences OOM at length of 16384. We use Llama-3.1-8B-Instruct for experiment.
6.2.3. Ablation Study
To better understand the contribution of each component within LoopServe, we conducted ablation studies by evaluating two variants: LoopServe-D, which uses only progressive KV compression, and LoopServe-P, which applies only online prefilling sparsification. We conduct experiments on three datasets, i.e., MultifieldQA (MF), 2WikiMQA (2WM), and Qasper (Qsp), with both Llama and Qwen backbones. As shown in Figure 7, LoopServe delivers the highest and comparable performance, demonstrating that the integration of both online prefilling sparsification and progressive KV compression is essential for optimal results. The ablation study further reveals that these two components are complementary: while each addresses a different bottleneck in LLM inference, their combination ensures robust adaptation to diverse input patterns and maximizes both efficiency and accuracy. These findings are stable across all task types, datasets, and query positions, as well as for both tested LLM architectures.
6.3. Parameter Sensitivity
To further evaluate the robustness and adaptability of LoopServe, we conducted a parameter sensitivity analysis focusing on three key hyperparameters: the attention sparsity threshold $\alpha$ used in the online prefilling phase, and the token budget $B$ and the decoding interval $n_{d}$ in the progressive KV compression phase.
(a) Llama-3.1-8B-Instruct.
(b) Qwen-2.5-7B-Instruct.
Figure 7. Ablation Study.
6.3.1. Threshold $\alpha$ in Equation (6).
The parameter $\alpha$ determines how much total attention weight is preserved by the selected vertical and slash lines during online prefilling sparsification. Higher $\alpha$ retains more information but increases computation, while lower $\alpha$ improves efficiency at the risk of losing important context. Specifically, we conduct experiments on 2WikiMQA and Qasper data with Llama and Qwen backbones, containing questions at the beginning (-B), middle (-M), and end (-E) positions. We vary $\alpha$ within $\alpha\in\{0.980,0.985,0.990,0.995,1.00\}$. As shown in Figure 8, LoopServe achieves a favorable balance in the range of $\alpha=0.99$ to $\alpha=1.00$, where the accuracy remains high and the computational gains are significant. Notably, setting $\alpha$ too low can cause a drop in generation quality, while excessively high values diminish the efficiency benefits. These results indicate that LoopServe is not overly sensitive to the precise value of $\alpha$ within a reasonable range, and users can adjust this parameter to trade off between speed and output quality according to their application needs.
(a) Llama-3.1-8B-Instruct.
(b) Qwen-2.5-7B-Instruct.
Figure 8. Parameter Sensitivity to $\alpha$
6.3.2. Budget $B$.
The token selection budget $B$ in LoopServe’s progressive KV compression impacts the balance between efficiency and output quality. Experiments conducted on MultiFieldQA and Qasper across three query positions (beginning, middle, and end). As shown in Table 9, increasing $B$ generally improves performance, as retaining more tokens captures a larger portion of relevant context. However, beyond a threshold (e.g., 1024 tokens), the performance gains diminish, while computational and memory costs increase. Smaller budgets (e.g., 256 or 512) result in lower accuracy, particularly for queries in the beginning or middle of the input, as key tokens may be excluded. In contrast, queries at the end are less sensitive to budget variations, as the most relevant tokens are naturally included in the cache. A budget of 1024 to 2048 tokens achieves the best balance, providing high performance across all query positions while maintaining computational efficiency.
(a) MultiFieldQA.
(b) Qasper.
Figure 9. Parameter Sensitivity to $B$.
(a) Llama-3.1-8B-Instruct.
(b) Qwen-2.5-7B-Instruct.
Figure 10. Parameter Sensitivity to $n_{d}$
6.3.3. The decoding interval $n_{d}$ in Algorithm 3.
The decoding interval $n_{d}$ determines how frequently LoopServe re-selects important input tokens based on the most recent output during progressive KV compression. A smaller $n_{d}$ allows the system to adapt more frequently to evolving output dependencies, which can further boost accuracy in challenging, dynamic dialogues. However, this comes with increased overhead due to more frequent KV cache updates. Specifically, we conduct experiments on MultiNews and GovReport data with Llama and Qwen backbones, containing questions at the beginning (-B), middle (-M), and end (-E) positions. As shown in Figure 10, our sensitivity analysis finds that moderate values of $n_{d}$ (e.g., 16 or 32) provide the best balance: they preserve the majority of the efficiency gains while maintaining robust generation quality. Extremely large $n_{d}$ values reduce the adaptivity of the KV cache and may result in decreased relevance of selected tokens, leading to lower performance on complex, multi-turn tasks.
7. Conclusion
In this paper, we propose LoopServe, an adaptive dual-phase LLM inference acceleration system designed for realistic multi-turn dialogues. By combining online attention sparsification and progressive KV compression, LoopServe addresses the limitations of static acceleration methods and adapts efficiently to dynamic conversational patterns. Our experiments on diverse, multi-turn benchmarks show that LoopServe significantly improves both inference speed and output quality compared to existing baselines, regardless of query position. This work provides a practical solution for efficient and effective LLM deployment in real-world dialogue scenarios.

### v2

@@
 LoopServe: An Adaptive Dual-phase LLM Inference Acceleration System for Multi-Turn Dialogues
-Abstract.
+Abstract
 Multi-turn dialogues are essential in many real-world applications of large language models, such as chatbots and virtual assistants.
@@
 Most current acceleration methods either compress the context or optimize key value caching, but they often rely on fixed or position-based heuristics that do not adapt well to the dynamic and unpredictable patterns found in actual multi-turn conversations.
+As a result, these models cannot accurately identify and prioritize the most relevant context, leading to degraded response quality.
 In this paper, we present LoopServe, an adaptive dual-phase inference acceleration framework for large language models in multi-turn dialogues.
@@
 Extensive experiments demonstrate that LoopServe consistently achieves superior effectiveness compared to existing baselines and significantly accelerates LLM inference across a wide range of long-context dialogue tasks.
-PVLDB Reference Format:
-PVLDB, 18(1): XXX-XXX, 2025.
-doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
-Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
-For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
-Copyright is held by the owner/author(s).
-Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097. doi:XX.XX/XXX.XX
-1. Introduction
+1 Introduction
 Multi-turn dialogues are at the core of numerous real-world applications, from customer service chatbots to virtual assistants and collaborative agents.
-These scenarios demand that large language models (LLMs) (Hadi et al., 2023; Deng et al., 2025; Li et al., 2024c; Zhou et al., 2024; van Renen et al., 2024) not only generate coherent responses but also maintain contextual consistency across lengthy, evolving conversations.
+These scenarios demand that large language models (LLMs) (Hadi et al., 2023; Deng et al., 2025; Li et al., 2024a; Zhou et al., 2024; van Renen et al., 2024) not only generate coherent responses but also maintain contextual consistency across lengthy, evolving conversations.
 As the number of dialogue turns increases, so does the computational workload.
 For instance, processing a multi-turn dialogue comprising 10,000 tokens with Llama-3.1-70B (Grattafiori et al., 2024) can demand trillions of floating-point operations (FLOPs), quickly challenging the limits of real-time inference.
-Despite the remarkable progress of LLMs such as GPT (Brown et al., 2020; Radford et al., 2018, 2019), Llama (Grattafiori et al., 2024; Touvron et al., 2023), and DeepSeek (DeepSeek-AI, 2024, 2025), their inefficiency in handling multi-turn dialogues remains largely unaddressed.
+Despite the remarkable progress of LLMs such as GPT (Brown et al., 2020; Radford et al., 2018; 2019), Llama (Grattafiori et al., 2024; Touvron et al., 2023), and DeepSeek (DeepSeek-AI, 2024; 2025), their inefficiency in handling multi-turn dialogues remains largely unaddressed.
 In a typical multi-turn setting, each new user turn expands the conversation history, requiring the LLM to process ever-growing input sequences.
@@
 This compounding effect of context accumulation makes real-time, cost-efficient inference increasingly difficult as the conversation progresses.
-This challenge is distinct from single-turn tasks, where the input context is static and the user query is typically placed at the end of the sequence.
-In contrast, multi-turn dialogues feature dynamic context accumulation and variable query positions: user questions may appear at the beginning, middle, or end of the input.
-This variability causes the attention patterns within the model to change unpredictably across turns and inputs, requiring the model to flexibly maintain and retrieve relevant information from any point in the conversation.
-The resulting attention matrices not only grow in size with each turn but also exhibit highly dynamic and input-dependent sparsity patterns, further exacerbating the inefficiency of existing inference methods.
-Recent research on LLM inference acceleration has primarily focused on two directions: context compression and Key-Value (KV) compression.
-Firstly, Context compression techniques, such as LLMLingua (Pan et al., 2024), CompAct (Yoon et al., 2024), and RAG-based approaches (Zhao et al., 2024; Jiang et al., 2024b; Wang et al., 2025; Chen et al., 2025; Edge et al., 2024; Wang et al., 2025), reduce input length by filtering, summarizing, or structuring the context.
-While effective in lowering computational costs, these strategies often sacrifice fine-grained details, which can degrade output fidelity.
-Secondly, KV compression approaches aim to reduce the computational burden of attention weight calculations during both the prefilling and decoding stages.
-In the prefilling stage, where the attention matrix is computed for all token pairs, methods (Jiang et al., 2024a; Lv et al., 2024; Lai et al., 2025), such as Minference (Jiang et al., 2024a), sparsify the attention matrix to reduce quadratic computation (See Table 2).
-In the decoding stage, KV caches store precomputed Key and Value vectors for reuse, minimizing redundant computations during autoregressive text generation.
-For example, H2O (Zhang et al., 2023b), PQCache (Zhang et al., 2025), SnapKV (Li et al., 2024b), and AdaKV (Feng et al., 2024) select or cache only the most relevant tokens based on attention weights or fixed token positions.
-However, these methods predominantly rely on static or position-based heuristics, which are ill-suited for the dynamic, input-dependent patterns observed in real-world multi-turn dialogues.
-Moreover, current evaluation benchmarks (Li et al., 2024c, 2025; Hsieh et al., 2024; Kim et al., 2025; An et al., 2023) for LLM acceleration misrepresent real-world dialogue scenarios.
-Most benchmarks (Bai et al., 2024a; Li et al., 2025; Dacheng Li* and Zhang, 2023) assume queries are always placed at the end of the input and focus on single-turn tasks, which oversimplifies the problem and favors acceleration methods that exploit positional biases.
-As a result, approaches (Zhang et al., 2023b; Li et al., 2024b; Feng et al., 2024) that perform well on these benchmarks often fail to generalize to realistic dialogue scenarios where queries may appear at arbitrary positions and contextual dependencies span multiple turns.
-To address these challenges, we conduct experiments and summarize three key findings: (1) The sparsity patterns of attention matrices are highly dynamic across inputs, heads, and layers.
-A small fraction of vertical and slash lines in the attention matrix accounts for most attention weights, but these patterns vary significantly and cannot be effectively captured by static or offline sparsification methods.
-(2) The position of questions in user queries significantly affects the performance of existing LLM acceleration techniques, which assume queries are always at the end of the input sequence.
-However, queries often appear in the middle or beginning of the context.
-(3) Token importance evolves dynamically during multi-turn dialogues as new outputs are generated.
-Fixed token selection strategies fail to capture these changing dependencies, leading to inefficiencies and reduced conversational quality.
-Based on these observations, we propose LoopServe, an adaptive dual-phase LLM inference acceleration framework specifically designed for multi-turn dialogues.
-LoopServe introduces two key innovations: online prefilling sparsification and progressive KV compression.
-In the prefilling phase, LoopServe dynamically identifies and selects the most critical components of the attention matrix, specifically the vertical and slash lines that account for the majority of attention weights.
-Unlike fixed sparsification methods, LoopServe adapts online, ensuring computational efficiency while maintaining high attention fidelity.
-In the decoding phase, LoopServe employs a progressive KV compression strategy.
-It dynamically selects and compresses relevant input tokens based on the most recently generated output tokens.
-This approach ensures the KV cache remains efficient and relevant throughout the decoding process, significantly reducing computational overhead without compromising output quality.
-Also, we propose a multi-turn long-context benchmark, including 11 datasets.
-This benchmark captures diverse query positions and multi-turn dependencies, providing a more realistic evaluation framework for multi-turn dialogue scenarios.
+Different from single-turn tasks, the context in multi-turn dialogues accumulates dynamically and queries may appear at the beginning, middle, or end of the input, causing attention patterns to shift unpredictably.
+The resulting attention matrices not only grow with each turn but also exhibit highly dynamic and input-dependent sparsity, exacerbating the inefficiency of current inference methods.
+Recent research proposes accelerating LLM inference by reducing the computational burden of attention weight calculations during both the prefilling and decoding stages.
+In the prefilling stage, where the attention matrix is computed for all token pairs, methods (Jiang et al., 2024a; Lv et al., 2024; Lai et al., 2025), such as Minference (Jiang et al., 2024a), use fixed pattern to sparsify the attention matrix to reduce quadratic computation.
+During decoding, KV caches store precomputed Key and Value vectors to reduce redundant computation.
+Methods like H2O (Zhang et al., 2023), SnapKV (Li et al., 2024b), and AdaKV (Feng et al., 2024) cache tokens selected based on tokens at the end of the query.
+However, these approaches rely on static or position-based heuristics and cannot adapt to the dynamic, input-dependent patterns of real multi-turn dialogues.
+Moreover, current evaluation benchmarks (Li et al., 2024a; 2025; Hsieh et al., 2024; Kim et al., 2025; An et al., 2023) for LLM acceleration misrepresent real-world dialogue scenarios.
+Most benchmarks (Bai et al., 2024a; Li et al., 2025; Dacheng Li* & Zhang, 2023) assume queries are always placed at the end of the input and focus on single-turn tasks, which oversimplifies the problem and favors acceleration methods that exploit positional biases.
+As a result, approaches (Zhang et al., 2023; Li et al., 2024b; Feng et al., 2024) that perform well on these benchmarks often fail to generalize to realistic dialogue scenarios where queries may appear at arbitrary positions and contextual dependencies span multiple turns.
+In this paper, we propose LoopServe, an adaptive dual-phase LLM inference acceleration framework specifically designed for multi-turn dialogues.
+LoopServe features two core innovations: online prefilling sparsification and progressive KV compression.
+In the prefilling phase, LoopServe dynamically identifies and selects the most critical components of the attention matrix, focusing on the vertical and slash line patterns that contribute most to attention weights.
+Unlike fixed sparsification methods, LoopServe adapts in real time to maintain both efficiency and high attention fidelity.
+During decoding, LoopServe applies progressive KV compression by dynamically selecting and compressing relevant input tokens based on the most recently generated outputs.
+This strategy keeps the KV cache efficient and relevant throughout decoding, significantly reducing computational overhead without compromising output quality.
+We also introduce a multi-turn long-context benchmark containing 11 datasets.
+This benchmark captures diverse query positions and multi-turn dependencies, offering a more realistic evaluation framework for dialogue scenarios.
 The contributions of this paper are summarized as follows:
-  • We empirically demonstrate that attention patterns and key token positions in multi-turn dialogues are highly dynamic and unpredictable, revealing the limitations of static sparsification and KV selection methods.
-  • We introduce LoopServe, a novel dual-phase LLM inference acceleration system that integrates online attention sparsification and progressive KV compression, substantially improving efficiency and robustness for multi-turn dialogue inference.
-  • We propose a new benchmark with 11 diverse multi-turn datasets.
-This benchmark incorporates diverse query positions and multi-turn dependencies, offering a comprehensive evaluation framework for LLMs in realistic long-context conversational tasks.
-  • Extensive experiments conducted on 11 long-context multi-turn datasets demonstrate that LoopServe consistently outperforms state-of-the-art LLM acceleration models.
-2. Preliminary and Related Work
-In this section, we first introduce the preliminary of LLMs in Section 2.1, and then introduce various techniques that enhance the efficiency of long-context LLM inference in Section 2.2.
-Finally, we analyze existing benchmarks for LLM acceleration in Section 2.3.
-2.1. Large Language Models
-LLMs, such as GPT (Brown et al., 2020), Llama (Grattafiori et al., 2024), and DeepSeek (DeepSeek-AI, 2024, 2025), demonstrate exceptional abilities in context understanding and reasoning, achieving significant success in various tasks, such as data cleaning (Feuer et al., 2024; Sun et al., 2024; Ni et al., 2024; Zhu et al., 2024c; Eltabakh et al., 2024; Shankar et al., 2024; Ang et al., 2024; Yan et al., 2024), data management (Li et al., 2024e; Zhang et al., 2024; Halevy et al., 2024; Lou et al., 2024; Huang et al., 2024; Le et al., 2024; Tang et al., 2024; Nagrecha and Kumar, 2023), data visualization (Arora et al., 2023; Wu et al., 2024), data security (Li et al., 2024a), recommendation system (Yao et al., 2023; Zheng et al., 2024), text to SQL (van Renen et al., 2024; Fan et al., 2024; Li et al., 2024d; Zhu et al., 2024a; Ren et al., 2024; Luoma and Kumar, 2025; Zhang et al., 2023a), and information extraction (Zhu et al., 2024b; Xue et al., 2024; Singh et al., 2023; Hu et al., 2024; Tang et al., 2024; Dargahi Nobari and Rafiei, 2024; Chen and Koudas, 2024; Omar et al., 2025; Yu et al., 2025; Huo et al., 2024).
-Their impressive performance stems from extensive training datasets and a transformer architecture, which effectively captures long-range dependencies and enables advanced autoregressive generation.
-2.1.1. Transformer Decoder
-The effectiveness of Transformer decoders stems from the Multi-Head Self-Attention (MHSA) mechanism, which efficiently captures both local and global dependencies among input tokens.
-For clarity, we analyze a single representative Transformer block, as the operational principles are consistent across all blocks in the architecture.
-Given an input token sequence ${X}=[x_{1},x_{2},\cdots,x_{n}]$ with embeddings $\mathbf{X}\in\mathbb{R}^{n\times d}$, the self-attention mechanism computes weighted aggregations over the token sequence, where the weights are derived from inter-token similarity scores.
-Specifically, the MHSA mechanism generates query vectors $\mathbf{Q}^{i}\in\mathbb{R}^{n\times d_{k}}$, key vectors $\mathbf{K}^{i}\in\mathbb{R}^{n\times d_{k}}$, and value vectors $\mathbf{V}^{i}\in\mathbb{R}^{n\times d_{v}}$ for the $i$-th attention head through learned linear transformations:
-(1) $\mathbf{Q}^{i}=\mathbf{X}\mathbf{W}_{Q^{i}},\quad\mathbf{K}^{i}=\mathbf{X}\mathbf{W}_{K^{i}},\quad\mathbf{V}^{i}=\mathbf{X}\mathbf{W}_{V^{i}},$
-where $\mathbf{W}_{Q^{i}}\in\mathbb{R}^{d_{x}\times d_{k}}$, $\mathbf{W}_{K_{i}}\in\mathbb{R}^{d_{x}\times d_{k}}$ and $\mathbf{W}_{V_{i}}\in\mathbb{R}^{d_{x}\times d_{v}}$ are learnable parameters.
-The self-attention operation is subsequently applied to each triple $(\mathbf{Q}^{i},\mathbf{K}^{i},\mathbf{V}^{i})$, yielding the output of the $i$-th attention head $\mathbf{Z}^{i}$ as follows:
-(2) $\mathbf{Z}^{i}=\textsf{Attention}(\mathbf{Q}^{i},\mathbf{K}^{i},\mathbf{V}^{i})=\textsf{Softmax}\left(\frac{\mathbf{Q}^{i}(\mathbf{K}^{i})^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}^{i},$
-where the scaling factor $\sqrt{d_{k}}$ ensures numerical stability (Vaswani, 2017).
-To capture diverse semantic relationships, $h$ parallel attention heads process $\mathbf{X}$ concurrently.
-Then, $\mathbf{Z}=\textsf{Concat}(\mathbf{Z}^{1},\mathbf{Z}^{2},\dots,\mathbf{Z}^{h})\mathbf{W}_{O},$ where Concat is the concatenation operation, and $\mathbf{W}_{O}\in\mathbb{R}^{h\cdot d_{v}\times d_{o}}$ is the trainable projection matrix.
-Following the self-attention mechanism, the output is passed through a Feed Forward Network, whose result serves as the input to the next Transformer block.
-2.1.2. Auto-regressive Generation Mechanism
-LLMs utilize an autoregressive generation paradigm, where text is produced sequentially with each token conditioned on the tokens generated before it.
-This step-by-step approach maintains semantic coherence and ensures contextual relevance across the output.
-Formally, given an input sequence $X=[x_{1},x_{2},\cdots,x_{n}]$, LLM predicts the subsequent token $x_{n+1}$ by estimating the conditional probability distribution:
-(3) $P(x_{n+1}|x_{1},x_{2},\cdots,x_{n})=\textsf{Softmax}(\mathbf{h}_{n}\mathbf{W}_{\text{out}}+\mathbf{b}_{\text{out}}),$
-where $\mathbf{h}_{n}\in\mathbb{R}^{d_{h}}$ represents the hidden state encoding for sequence $X$ at step $n$, $\mathbf{W}_{\text{out}}\in\mathbb{R}^{d_{h}\times n_{vocal}}$ denotes the output projection matrix mapping to vocabulary size $n_{vocal}$.
-The softmax function transforms logits into a probability distribution across the vocabulary space.
-During generation, the model samples the next token $x_{n+1}$ from the computed probability distribution: $x_{n+1}\sim P(x_{n+1}|x_{1},x_{2},\cdots,x_{n})$.
-The newly generated token $x_{n+1}$ is appended to the sequence $X$ to obtain resulting $X=[x_{1},\cdots,x_{n},x_{n+1}]$.
-This process iterates until either a designated end-of-sequence (EOS) token is produced or a predetermined maximum sequence length is reached.
-2.2. Efficient Long-Context Inference
-The impressive abilities of LLMs across various tasks weaken significantly when handling long contexts.
-This happens mainly because of basic limits in computation and memory.
-The drop in performance, especially efficiency, is due to the quadratic complexity of self-attention mechanisms in Equation (2), which require calculating interactions between every pair of tokens in the input sequence.
-Specifically, for an LLM architecture comprising $L$ layers with $h$ attention heads per layer, processing an input sequence of length $n$ requires computational operations scaling at $O(Lhn^{2})$.
-This quadratic growth in both computational demand and memory utilization renders the handling of extensive sequences prohibitively expensive (Li et al., 2024c).
-To overcome these efficiency challenges, researchers have created specialized techniques to speed up long-context LLM processing.
-These methods typically fall into two main categories: context compression and KV-based optimization.
-Table 1. Summary of important notations.
+  • We empirically reveal that attention patterns and key token positions in multi-turn dialogues are highly dynamic, limiting static sparsification and KV selection.
+  • We present LoopServe, a dual-phase LLM acceleration framework with online attention sparsification and progressive KV compression, improving multi-turn inference efficiency.
+  • We introduce a benchmark of 11 long-context multi-turn datasets with varied query positions and dependencies for realistic evaluation.
+  • Experiments on 11 multi-turn datasets demonstrate the superior performance of LoopServe.
+2 Preliminary and Related Work
+2.1 Large Language Models
+LLMs like GPT (Brown et al., 2020), Llama (Grattafiori et al., 2024), and DeepSeek (DeepSeek-AI, 2024; 2025) excel at context understanding and reasoning, enabled by large-scale training and the Transformer architecture (Vaswani, 2017).
+Transformers are effective due to Multi-Head Self-Attention (MHSA), which captures both local and global token dependencies.
+Given an input sequence $X=[x_{1},x_{2},\cdots,x_{n}]$ with embeddings $\mathbf{X}\in\mathbb{R}^{n\times d}$, the MHSA computes query vectors $\mathbf{Q}^{i}\in\mathbb{R}^{n\times d_{k}}$, key vectors $\mathbf{K}^{i}\in\mathbb{R}^{n\times d_{k}}$, and value vectors $\mathbf{V}^{i}\in\mathbb{R}^{n\times d_{v}}$ for the $i$-th attention head as $\mathbf{Q}^{i}=\mathbf{X}\mathbf{W}_{Q^{i}},\mathbf{K}^{i}=\mathbf{X}\mathbf{W}_{K^{i}},\mathbf{V}^{i}=\mathbf{X}\mathbf{W}_{V^{i}}$, where $\mathbf{W}_{Q^{i}}$, $\mathbf{W}_{K^{i}}$, and $\mathbf{W}_{V^{i}}$ are learnable matrices.
+Each $i$-th attention head $\mathbf{Z}^{i}$ as : $\mathbf{Z}^{i}=\textsf{Attention}(\mathbf{Q}^{i},\mathbf{K}^{i},\mathbf{V}^{i})=\textsf{Softmax}\left(\frac{\mathbf{Q}^{i}(\mathbf{K}^{i})^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}^{i}.$ Next, outputs from $h$ heads are concatenated and projected: $\mathbf{Z}=\textsf{Concat}(\mathbf{Z}^{1},\mathbf{Z}^{2},\dots,\mathbf{Z}^{h})\mathbf{W}_{O},$ where $\mathbf{W}_{O}$ is a learned projection.
+For text generation, LLMs use an autoregressive process: given $X=[x_{1},\ldots,x_{n}]$, the model predicts the next token $x_{n+1}$ by modeling $P(x_{n+1}|x_{1},x_{2},\cdots,x_{n})=\textsf{Softmax}(\mathbf{h}_{n}\mathbf{W}_{\text{out}}+\mathbf{b}_{\text{out}})$, where $\mathbf{h}_{n}$ is the state at step $n$.
+The next token $x_{n+1}$ is sampled from this distribution and appended to the sequence.
+Generation continues until an end-of-sequence token or a maximum length is reached.
+2.2 Efficient Long-Context Inference
+The performance of LLMs degrades with long input contexts, due to the quadratic complexity of self-attention, which scales as $O(Lhn^{2})$ for $L$ layers, $h$ heads, and sequence length $n$ (Li et al., 2024a).
+It makes long-sequence processing prohibitively expensive.
+Context Compression Methods.
+Context compression methods reduce the effective sequence length, transforming lengthy inputs into more manageable representations.
+Filtering-based approaches such as LLMLingua (Jiang et al., 2023), LLMLingua-v2 (Pan et al., 2024), and CompAct (Yoon et al., 2024) focus on identifying and preserving high-relevance content, allowing models to process only critical information.
+In contrast, RAG-based (retrieval-augmented generation) methods (Zhao et al., 2024; Jiang et al., 2024b; Wang et al., 2025; Chen et al., 2025; Edge et al., 2024) construct knowledge graphs or extract semantic triples from the input, synthesizing them into condensed forms for LLMs.
+These strategies substantially decrease computational and memory costs, but may sacrifice fine-grained details, potentially affecting output quality.
+Also, to reduce computational burden, KV-based approaches minimize the number of attention weight calculations during both prefilling and decoding, summarized in Table 3 in the Appendix.
+Prefilling-stage Optimization.
+Self-attention requires $O(n^{2})$ computation for an input of length $n$.
+Recent methods such as Minference (Jiang et al., 2024a), FlexPrefill (Lai et al., 2025), and CritiPrefill (Lv et al., 2024) use binary masks $\mathbf{M}\in\{0,1\}^{n\times n}$ to zero out less important attention weights: $\min\,\left|\mathbf{{A}}^{j}_{i}-\mathbf{\hat{A}}^{j}_{i}\right|,s.t.\ ,\mathbf{\hat{A}}^{j}_{i}=\textsf{Softmax}\left(\frac{\mathbf{Q}^{j}(\mathbf{K}^{j})^{\top}}{\sqrt{d_{k}}}-c(1-\mathbf{M})\right)$, where $c$ is a large constant that suppresses masked entries.
+This reduces complexity to $O(\alpha n^{2})$ per head, with $\alpha\ll 1$.
+However, Minference (Jiang et al., 2024a) and CritiPrefill (Lv et al., 2024) rely on fixed attention patterns or block selection, while FlexPrefill (Lai et al., 2025) adjusts sparsity globally with simple heuristics.
+However, our experiments (Section 3) show that attention patterns are highly input-dependent and dynamic, so these static or coarse methods struggle to adapt in multi-turn scenarios.
+Decoding-stage Optimization.
+During autoregressive generation, KV cache methods such as H2O (Zhang et al., 2023), SnapKV (Li et al., 2024b), AdaKV (Feng et al., 2024), and others (Ge et al., 2024; Li et al., 2024b) select important tokens to store, reducing redundancy.
+These approaches assume that critical tokens are near the end of the input, performing well on benchmarks like LongBench (Bai et al., 2024a; b), where queries are always placed last.
+However, as analyzed in Section 3.2, their effectiveness drops when queries appear elsewhere, underscoring the need for adaptive, context-aware KV selection in real dialogue.
+3 Motivational Experiments and Insights
+To clarify the core challenges in accelerating LLM inference for multi-turn dialogues, we conduct motivational experiments in Section 3.1 and 3.2.
+They reveal how attention patterns are dynamically sparse and how query position influences acceleration effectiveness.
+Building on these findings, we introduce the LoopServe system, specifically designed to address these real-world challenges.
+3.1 Key Point 1: Uncertain Attention Patterns
+As investigated previously, attention head matrices are highly sparse Jiang et al.
+(2024a).
+Existing acceleration methods (Xiao et al., 2024; Feng et al., 2024; LI et al., 2025; Li et al., 2024b) often rely on sparsifying attention matrices or selecting important KV tokens based on the assumption that attention patterns are fixed and can be identified offline.
+In reality, these patterns are highly variable across inputs, heads, and layers, limiting the effectiveness of such approaches, as reveled as follows.
+(a) Vertical and slash lines.
+(b) All heads of Llama 3.1.
+(c) Average ratio on all heads.
+(d) Overlap among different inputs.
+Figure 1: Attention head sparsity is shown in (b) and (c).
+Motivational Observation 1: Only 10% of vertical and slash lines can collectively account for most (e.g., 90%) of the attention weight.
+As shown in Figure 1 (a), in an attention matrix, a vertical line (column) indicates a single token attended by all others, common for special tokens like separators or keywords.
+A slash line (diagonal) shows each token mostly focusing on its nearby tokens, reflecting local attention patterns.
+We analyze this using the SAMSum QA dataset (Bai et al., 2024a) and the Llama-3.1-8B-Instruct with $n_{h}$ heads.
+For each query $X_{i}$ of length $n_{i}$, each $k$-th attention matrix $\mathbf{A}^{k}_{i}$ contains $n_{i}$ vertical lines $\mathcal{V}^{k}_{i}$ and $n_{i}$ slash lines $\mathcal{S}^{k}_{i}$.
+The total attention weight for a slash line $s^{k}$ is $\sum_{(a,b)\in s^{k}}\mathbf{A}^{k}_{i}[a][b]$, and for a vertical line $v^{k}$, it is $\sum_{(a,b)\in v^{k}}\mathbf{A}^{k}_{i}[a][b]$.
+We select the top $\eta\cdot 2n_{i}$ slash and vertical lines ($\mathcal{\hat{S}}^{k}_{i}$, $\mathcal{\hat{V}}^{k}_{i}$) based on their total weights.
+For each head $k$, we compute the ratio of weight within these lines: $r^{k}_{i}=\frac{1}{|\mathcal{D}|}\sum_{X_{i}\in\mathcal{D}}\frac{\sum_{(a,b)\in\mathcal{\hat{S}}^{k}_{i}\cup\mathcal{\hat{V}}^{k}_{i}}\mathbf{A}^{k}_{i}[a][b]}{n_{i}}.$ Averaging over all $n_{h}$ heads yields the mean ratio.
+As shown in Figure 1 (b), higher $\eta$ increases cumulative attention weight.
+Figure 1 (c) shows that for both Llama and Mistral, just 10% of slash and vertical lines account for 90% of total attention, indicating highly concentrated attention and enabling efficient selection by focusing on these sparse lines.
+Motivational Observation 2: The positions of the top vertical and slash lines within the same head vary across different user inputs.
+For a model $M_{\theta}$ with $n_{h}$ attention heads and dataset $\mathcal{D}={X_{i}}$, we select the top $\eta\cdot 2n_{i}$ vertical and slash lines ($\mathcal{\hat{V}}^{k}_{i}$, $\mathcal{\hat{S}}^{k}_{i}$) for each input $X_{i}$ and attention head $k$.
+For any pair of queries $X_{i}$, $X_{j}$, the overlap of their selected lines under head $k$ is: $r^{k}_{i,j}=\frac{|\mathcal{\hat{S}}_{i}^{k}\cap\mathcal{\hat{S}}_{j}^{k}|+|\mathcal{\hat{V}}_{i}^{k}\cap\mathcal{\hat{V}}_{j}^{k}|}{|\mathcal{\hat{S}}_{i}^{k}\cup\mathcal{\hat{S}}_{j}^{k}|+|\mathcal{\hat{V}}_{i}^{k}\cup\mathcal{\hat{V}}_{j}^{k}|}.$ Averaging over all heads and input pairs gives the mean overlap ratio: $\frac{1}{n_{h}|\mathcal{D}|^{2}}\sum_{k=1}^{n_{h}}\sum_{X_{i},X_{j}\in\mathcal{D}}r^{k}_{i,j}$.
+Using the SAMSum QA dataset (Bai et al., 2024a) and models Llama-3.1-8B-Instruct and Mistral-7B-Instruct-v0.3, with $\eta$ ranging from 0.1 to 0.3, Figure 1 (d) and Figure 7 in Appendix show that for most heads, the overlap remains below 0.5.
+This indicates that the most important lines differ significantly depending on the input, even within the same head.
+As a result, important lines cannot be reliably determined offline for use during online inference.
+Motivational Observation 3: For an input $X_{i}=[C_{i}^{1},C^{2}_{i}]$ split into two segments, the top vertical and slash lines within the same head differ between $C^{1}_{i}$ and $C^{2}_{i}$.
+Each segment shows its own local attention sparsity pattern.
+As illustrated in Figure 2 (a), the key vertical and slash lines in $C^{1}_{i}$’s attention matrix are largely absent in $C^{2}_{i}$, which displays distinct local patterns.
+To verify this, we use the SAMSum QA (Bai et al., 2024a) dataset and Llama-3.1-8B-Instruct.
+For each $X_{i}$ in dataset $\mathcal{D}$, we split it into $C^{1}_{i},C^{2}_{i}]$ and extract the attention matrices $\mathbf{A}^{k}_{C^{1}_{i}}$ and $\mathbf{A}^{k}_{C^{2}_{i}}$ for each head $k$.
+After selecting the top-$\eta$ important slash and vertical lines for each ($L^{k}_{C^{1}_{i}}$, $L^{k}_{C^{2}_{i}}$) for ( $\mathbf{A}^{k}_{C^{1}_{i}}$, $\mathbf{A}^{k}_{C^{2}_{i}}$), we compute the overlap rate: $r^{k}_{C_{i}^{1}\to C_{i}^{2}}={\sum_{l\in L^{k}_{C^{1}}}\mathbb{I}(l\in L^{k}_{C_{i}^{2}})}/{|L^{k}_{C_{i}^{2}}|},$ where $\mathbb{I}$ indicates whether a line from $C^{1}_{i}$ is also important in $C^{2}_{i}$.
+Averaging across data gives the mean overlap line rate for each head.
+Figure 2 (b) shows that, for different $\eta$, the overlap in important lines between $C^{1}_{i}$ and $C^{2}_{i}$ is consistently low and unstable.
+This confirms that each segment exhibits unique local attention patterns.
+This finding indicates that using only a segment (such as the last window or last few tokens) to predict important attention patterns for the whole input is unreliable.
+As a result, acceleration methods like Minference (Jiang et al., 2024a), SnapKV (Li et al., 2024b), H2O (Zhang et al., 2023), and Keyformer (Adnan et al., 2024), which rely on such assumptions, struggle to deliver consistent performance in real-world scenarios.
+3.2 Key Point 2: Question Position Matters.
+Motivational experiments indicate that both prefilling based methods and decoding phase acceleration methods, which depend on offline sparse pattern discovery or fixed sparse patterns, tend to underperform in practical scenarios.
+However, their reported outcomes on existing benchmarks are often similar to those of large language models that use full attention.
+Why does this discrepancy occur?
+The main reason is that benchmarks like Longbench (Bai et al., 2024a; b) always place the user question $q_{i}$ at the end of the input $X_{i}=[C_{i},q_{i}]$, so the LLM answers $q_{i}$ based on context $C_{i}$.
+In this setup, acceleration methods only need to focus on the last observation window (near the question), which makes it easier to identify context tokens relevant to the question, thus partially mitigating the unpredictability found in real-world input patterns.
+(a) Attention map.
+(b) Overlap of different segments.
+(c) Important token overlap.
+(d) Question in different position.
+Figure 2: Different attention sparsity patterns and query position impact on performance.
+Motivational Observation 4: Relying only on the last observation window cannot reliably identify important input tokens for generating the output.
+Recent methods like H2O (Zhang et al., 2023), SnapKV (Li et al., 2024b), and AdaKV (Feng et al., 2024) select the top-$B$ important tokens $\hat{X}_{i}$ from $X_{i}$ based on attention between each input token and the last observation window $X^{obs}_{i}$ (the last $n_{s}$ tokens), formally as $\hat{X}_{i}=\arg\max_{\hat{X}_{i}\subseteq X_{i}}\sum_{j=1}^{n_{h}}\sum_{a\in\hat{X}_{i}}\sum_{b\in X^{obs}_{i}}\mathbf{A}^{k}_{i}[a][b]$.
+However, the true top-$B$ tokens $\hat{X}^{*}_{i}$ for generating the output $Y_{i}$ should be selected based on their attention to the output tokens: $\hat{X}^{*}_{i}=\arg\max_{\hat{X}_{i}\subseteq X_{i}}\sum_{j=1}^{n_{h}}\sum_{a\in\hat{X}_{i}}\sum_{b\in Y_{i}}\mathbf{A}^{k}_{i}[a][b]$.
+We measure the overlap $r_{i}$ between $\hat{X}_{i}$ and $\hat{X}^{*}_{i}$ ($B$ is set to 10% of $|X_{i}|$), using Llama-3.1-8B-Instruct and LongEval’s topic retrieval set, where the instruction $q_{i}$ is placed at the beginning, middle, or end of $C_{i}$.
+As shown in Figure 2 (c), the average overlap of important tokens is highest when the question is at the end, but much lower when it appears in middle or the beginning.
+This demonstrates that focusing only on the last part of the input misses relevant information unless the question is placed last.
+Similarly, as shown in Figure 2 (d), SnapKV and AdaKV match the original model only when the question is at the end.
+Their performance drops sharply when the question appears earlier, since they rely on the last tokens for context selection.
+This shows that current methods are overly dependent on input order and do not generalize well when question positions vary.
+4 Multi-turn Long-Context Benchmarks
+Existing benchmarks, such as NumericBench (Li et al., 2025), LongBench (Bai et al., 2024a; b), and LongEval (Dacheng Li* & Zhang, 2023) focus on single-turn tasks and place user queries only at the context end, which do not reflect the complexity of real-world, multi-turn conversations (LI et al., 2025).
+We introduce a benchmark of 11 long-context multi-turn datasets with varied query positions and dependencies.
+Specifically, each $m$-turn instance is defined as $\mathcal{D}={\mathcal{I}_{i}=[(C_{i,1},q_{i,1},a_{i,1}),\dots,(C_{i,m},q_{i,m},a_{i,m})]}_{i=1}^{|\mathcal{D}|}$, where $C_{i,j}$ is the context at turn $j$ (possibly empty), $q_{i,j}$ is the user query, and $a_{i,j}$ is the LLM-generated answer.
+We ensure diversity by: (1) Query Position: For each turn, the query $q_{i,j}$ can appear at the beginning, end, or between paragraphs of $C_{i,j}$, reflecting more realistic query placements.
+(2) Query Relevance: Answers $a_{i,j}$ may depend on any subset of current and previous contexts $\{C_{i,1},\ldots,C_{i,j}\}$, with variable subset sizes, simulating diverse real-world dependencies.
+We construct multi-turn benchmarks for several tasks, including question answering, summarization, and few-shot learning.
+For construction procedures and detailed benchmark statistics, please refer to Appendix A.5.
+5 LoopServe System
+Figure 3: Framework overview of LoopServe.
+As shown in Figure 3, we propose LoopServe, an adaptive dual-phase system that performs online attention sparsification during the prefilling phase and progressive KV compression during the decoding phase.
+For an $m$-turn input $\mathcal{I}_{i}=\{X_{i,j}\}_{j=1}^{m}$, where $X_{i,j}$ is the context or query in the $j$-th turn, LoopServe generates each answer $y_{i,j}$ using these two steps.
+Step 1.
+Online Attention Head Sparsification in Prefilling.
+In Algorithm 1 (line 3-5) in Appendix, for each new input $X_{i,j}$ and all inputs $X_{i}=\left(\cup_{j^{\prime}=1}^{j-1}{(X_{i,j^{\prime}}\cup y_{i,j^{\prime}})}\right)\cup X_{i,j}$, we get $\hat{X}_{i,j}=[y_{i,j-1},X_{i,j}]$ as the new appended input in the $j$ turn.
+Then, for each $k$-th attention head, we first compute the attention matrix $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]\in\mathbb{R}^{\hat{n}_{i,j}\times n_{i}}$, and then select the slash lines $\mathcal{\hat{S}}^{k}_{i,j}$ and vertical lines $\mathcal{\hat{V}}^{k}_{i,j}$ that collectively recover at least $\alpha$ of the total attention weight in $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$.
+Step 2.
+Progressive KV Compression in Decoding.
+As described in Algorithm 1 (lines 7–14) in Appendix, after every re-selection interval $n_{d}$ tokens, the framework uses the ProgressiveSelection Algorithm 3 to compute a subset of input tokens $\hat{X}^{k}_{i}\subseteq X_{i}$ for each attention head $k$.
+these selected tokens $\hat{X}^{k}_{i}$ are important for output generation.
+At each decoding step, LoopServe leverages the compressed KV cache $\{\mathcal{\hat{S}}^{k}_{i,j^{\prime}},\mathcal{\hat{V}}^{k}_{i,j^{\prime}}\}_{j^{\prime}=1,k}^{j,n_{h}}$ to generate the output sequence $y_{i,j}$.
+5.1 Online Attention Sparsification in Prefilling
+As shown in Key Point 1 in Section 3.1, attention sparsity patterns are highly dynamic, making static or offline selection ineffective.
+To address this, we propose an online adaptive algorithm that, during prefilling, selects a subset of slash and vertical lines to recover at least an $\alpha$ fraction of the total attention weight for each head, which can be reused in later dialogue turns.
+Definition 1 (Online Prefilling Sparsification Problem).
+Given the LLM model $M_{\theta}$ with $n_{h}$ attention heads, the input $X_{i}=\left(\bigcup_{j^{\prime}=1}^{j-1}(X_{i,j^{\prime}}\cup y_{i,j^{\prime}})\right)\cup X_{i,j}$, where $y_{i,j^{\prime}}$ is the answer for $X_{i,j^{\prime}}$ and $X_{i,j}$ is the current turn’s input.
+We denote the concatenation of the previous answer $y_{i,j-1}$ and the current user input $X_{i,j}$ as $\hat{X}_{i,j}=[y_{i,j-1},X_{i,j}]$, whose corresponding attention matrix requires sparsification.
+The $k$-th attention matrix between $X_{i}$ and $\hat{X}_{i,j}$ is denoted as $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]\in\mathbb{R}^{\hat{n}_{i,j}\times n_{i}}$.
+Let $\mathcal{S}^{k}_{i,j}$ (resp., $\mathcal{V}^{k}_{i,j}$ ) denote the set of all slash lines (resp., vertical lines) in $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$.
+The goal is to select a subset of slash lines $\hat{\mathcal{S}}^{k}_{i,j}\subseteq\mathcal{S}^{k}_{i,j}$ and a subset of vertical lines $\hat{\mathcal{V}}^{k}_{i,j}\subseteq\mathcal{V}^{k}_{i,j}$ such that together they recover at least an $\alpha$ fraction of the total attention weight in $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$, where $\alpha\in[0,1]$.
+$\min\sum_{s\in\mathcal{\hat{S}}^{k}_{i,j}}l_{s}+\sum_{v\in\mathcal{\hat{V}}^{k}_{i,j}}l_{v},\quad s.t.\sum_{(a,b)\in(\mathcal{\hat{S}}^{k}_{i,j}\cup\mathcal{\hat{V}}^{k}_{i,j})}\mathbf{A}^{k}_{i}[a][b]\geq\alpha\cdot\hat{n}_{i,j},$ (1)
+where $\hat{n}_{i,j}$ is the total attention weight of the matrix $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$, and $l_{s}$ (resp., $l_{v}$) is the length of the slash line $s$ (resp., vertical line $v$).
+Theorem 1.
+The prefilling sparsification problem is NP-hard. The proof is detailed in Appendix A.6.
+Algorithm.
+Algorithm 2 in Appendix A.7 takes as input the concatenated sequence $\hat{X}{i,j}=[y_{i,j-1},X_{i,j}]$, where $y_{i,j-1}$ is the previous answer and $X_{i,j}$ is the current user input.
+It also requires the $k$-th attention head of the LLM $M_{\theta}$ and a sparsity threshold parameter $\alpha$.
+The output consists of the selected slash lines $\mathcal{\hat{S}}^{k}_{i,j}$ and selected vertical lines $\mathcal{\hat{V}}^{k}_{i,j}$.
+The algorithm begins by sampling a subset $\tilde{X}_{i,j}$ from the concatenated input $\hat{X}_{i,j}$ to reduce computational cost.
+It then computes the query matrix $\mathbf{\tilde{Q}}^{k}_{i,j}$ for $\tilde{X}_{i,j}$ and the key matrix $\mathbf{K}^{k}_{i}$ for the full input $X_{i}$.
+Next, all slash lines $\mathcal{S}^{k}{i,j}$ and vertical lines $\mathcal{V}^{k}_{i,j}$ are summarized based on $\mathbf{A}^{k}_{i}[\tilde{X}{i,j}]$ and sorted in descending order.
+Two empty sets, $\mathcal{\hat{S}}^{k}_{i,j}$ and $\mathcal{\hat{V}}^{k}_{i,j}$, are initialized to store the selected lines, while the overlap weights $ol_{s}$ and $ol_{v}$ are initialized to zero.
+Algorithm 2 then iteratively selects lines until the total recovered attention weight sum meets or exceeds $\alpha\cdot\textsf{sum}(\mathbf{A}^{k}_{i}[\tilde{X}_{i,j}])$.
+At each iteration, it compares the top slash line $s\in\mathcal{S}^{k}_{i,j}$ and the top vertical line $v\in\mathcal{V}^{k}_{i,j}$ based on their marginal contributions, $\Delta w_{s}=w_{s}-ol_{v}$ and $\Delta w_{v}=w_{v}-ol_{s}$.
+Since each slash line overlaps with only one vertical line, $\Delta w_{s}=w_{s}-ol_{v}$.
+The line with the greater marginal contribution is added to its respective set ($\mathcal{\hat{S}}^{k}_{i,j}$ or $\mathcal{\hat{V}}^{k}_{i,j}$), and the overlap weights and total recovered weight are updated accordingly.
+The loop terminates once the recovery condition is satisfied, and the algorithm returns the sets $\mathcal{\hat{S}}^{k}_{i,j}$ and $\mathcal{\hat{V}}^{k}_{i,j}$.
+The time complexity is $O(n_{i}|\tilde{X}_{i,j}|+n_{i}\log n_{i}+n_{i})$, as detailed in Appendix A.7.1.
+5.2 Progressive KV Compression in Decoding
+As shown in Section 3.2, fixed or last-token-focused decoding struggles when queries are not at the end of the input.
+To overcome this, we propose a progressive KV compression that selects important input tokens based on recent outputs, leading to greater overlap with truly important input tokens.
+To verify this, we consider the LLM $M_{\theta}$ with $n_{h}$ attention heads, an input sequence $X_{i}$, a generated output sequence $Y_{i}$, and a window size $n_{w}$.
+As in Section 3.2, we compute the ground-truth top-$B$ important tokens $\hat{X}^{*}_{i}\subseteq X_{i}$ using the output $Y_{i}$.
+We then use observation windows, extracted from either $X_{i}$ or $Y_{i}$, to select the top-$B$ important tokens.
+Specifically, given the window size $n_{w}$, the $\frac{|X_{i}|}{n_{w}}-$th to 1-1 observation windows from $X_{i}$ are defined as $X_{i}[n_{i}-m\cdot n_{w}:n_{i}-(m-1)\cdot n_{w}]$, $X_{i}[n_{i}-(m-1)\cdot n_{w}:n_{i}-(m-2)\cdot n_{w}]$, …, $X_{i}[n_{i}-n_{w}:n_{i}]$.
+The 0-th to $\frac{|Y_{i}|}{n_{w}}$ observation windows from $Y_{i}$ are $Y_{i}[0:n_{w}]$, $Y_{i}[n_{w}:2n_{w}]$, …, $Y_{i}[(m-1)\cdot n_{w}:m\cdot n_{w}]$.
+For each observation window $X^{obs}_{i}$ from $X_{i}$ or $Y_{i}$, we select the top-$B$ tokens $\hat{X}_{i}\subseteq X_{i}$ following Section 3.2, and compute the overlap rate between $\hat{X}_{i}$ and the ground truth $\hat{X}^{*}_{i}$ as $\frac{|\hat{X}_{i}\cap\hat{X}^{*}_{i}|}{B}$.
+(a) Important token overlapping.
+(b) Decoding blocks.
+Figure 4: Progressive decoding.
+We use LongEval (Dacheng Li* & Zhang, 2023) as in Section 3, experimenting with Llama-3.1-8B-Instruct and setting $B\in\{5\%,10\%,15\%\}\cdot|X_{i}|$.
+Figure 4(a) shows that important tokens selected from input blocks (block index <0<0) have low overlap rates, while those selected from output blocks ($\geq 0$) have much higher overlap.
+This indicates that using output tokens is more effective for identifying relevant input tokens.
+We also observe that the overlap of important input tokens between output blocks decreases as the distance between blocks increases, which indicates that we can use the most recent output tokens to select important input tokens for the next output block.
+Specifically, given the output sequence $Y_{i}$, we divide it into $n_{b}$ blocks, i.e., $Y_{i}=[Y^{1}_{i},\cdots,Y^{n_{b}}_{i}]$, where $Y^{j}_{i}$ denotes the $j$-th output block of size $\frac{|Y_{i}|}{n_{b}}$.
+For each block $Y^{j}_{i}$, we compute the top-$B$ important input tokens as: $\hat{X}_{i,Y^{j}_{i}}=\arg\max_{\hat{X}_{i}\subseteq X_{i}}\sum_{k=1}^{n_{h}}\sum_{a\in\hat{X}_{i}}\sum_{b\in Y^{j}_{i}}\mathbf{A}^{k}_{i}[a][b]$.
+Next, we compare the overlap $\frac{|\hat{X}_{i,Y^{j}_{i}}\cap\hat{X}_{i,Y^{j^{\prime}}_{i}}|}{B}$ of important input tokens between every pair of output blocks $Y^{j}_{i}$ and $Y^{j^{\prime}}_{i}$.
+As shown in Figure 4 (b), for each block $Y^{j}_{i}$, the overlap of important input tokens between $Y^{j}_{i}$ and $Y^{j^{\prime}}_{i}$ (where $j^{\prime}>j$) gradually decreases as the block distance increases.
+Notably, the overlap between $Y^{j}_{i}$ and its immediate successor $Y^{j+1}_{i}$ is higher compared to earlier blocks such as $Y^{j-2}_{i}$.
+This indicates that tokens identified from $Y^{j}_{i}$ are highly relevant for the generation of $Y^{j+1}_{i}$.
+Therefore, when generating $Y^{j+1}_{i}$, we can use the preceding block $Y^{j}_{i}$ to dynamically identify the top-$B$ important input tokens.
+Based on this, we propose the following progressive KV compression algorithm.
+Algorithm.
+Algorithm 3 firstly set the answer $y_{i,j}$ as empty and decoding step counter $n_{o}$ to 0 (line 1).
+During the decoding loop, if the current step reaches either the empirically predefined decoding size 16 or a re-selection interval $n_{d}$ (line 3), the algorithm extracts the most recent tokens from the input sequence $X_{i}$, i.e., $X^{obs}_{i}=X_{i}[|X_{i}|-n_{d}:|X_{i}|]$), and dynamically updates the compressed KV cache subset $\hat{X}^{k}_{i}$ for each $k$-th attention head (line 4-5).
+Specifically, for each head, we select top-$B$ tokens for each head as: $\hat{X}^{k}_{i}=\arg\max_{\hat{X}^{k}_{i}\subseteq X_{i},|\hat{X}^{k}_{i}|=B|}\sum_{a\in\hat{X}_{i}}\sum_{b\in X^{obs}_{i}}\mathbf{A}^{k}_{i}[a][b],$ The LLM generates the next token using the compressed KV cache and the updated input, which is appended to both the input and output sequences (line 6-9).
+This process iterates until the LLM completes generation, returning the final answer $y_{i,j}$.
+6 EXPERIMENTS
+6.1 Experimental Settings
+Datasets, Tasks, and Evaluation Metrics.
+We design multi-turn long-context benchmarks.
+Each instance contains multiple rounds with diverse query positions and dependencies.
+It covers Question Answering, Summarization, and Few-shot Learning.
+Dataset statistics and each corresponding metric (e.g., F1, Accuracy, and Rouge-L) are in Table 4 in Appendix A.5.2.
+Baselines.
+We compare our LoopServe with six state-of-the-art KV cache algorithms on two representation LLM base models, including Llama-3.1-8B-Instruct (Grattafiori et al., 2024) and Qwen2.5-7B-Instruct (Team, 2024).
+The KV cache methods include SnapKV (Snap) Li et al.
+(2024b), AdaKV (Ada) Feng et al.
+(2024), StreamingLLM (SLLM) Xiao et al.
+(2024), A-shape (A-S) Xiao et al.
+(2024), Tri-Shape (T-S) LI et al.
+(2025), and Minference (Minf) Jiang et al.
+(2024a).
+Hyperparameter and Hardware Setting.
+All codes are executed on a Rocky Linux 8.10 machine with an 8-core Intel® Xeon® Gold 6542Y CPU, an NVIDIA H100 GPU with 80GB of memory, and 256GB of RAM.
+For baselines, we use their suggested setting.
+For main experiments in Section 6.2, for our LoopServe, we set $\alpha=0.955$, and $n_{d}=16$ as defaults, and we set the token budget $B=1024$ following Li et al.
+(2024b)  Feng et al.
+(2024) for all baselines and LoopServe.
+Table 1: Effectiveness Evaluation. The bold number indicates the best performance.
+P Data Llama-3.1 Qwen2.5
+Base Snap Ada SLLM A-S T-S Minf Ours Base Snap Ada SLLM A-S T-S Minf Ours
+Begin
+QA
+MFQA-en 45.70 36.50 34.94 25.83 28.00 34.96 43.05 46.82 44.24 35.59 33.00 21.45 29.56 38.14 42.63 43.47
+2WikiMQA 31.68 29.31 28.73 22.05 19.42 24.20 28.57 35.11 37.86 34.54 33.83 23.19 22.11 32.90 35.33 37.52
+Musique 17.16 14.29 14.80 9.76 4.22 8.18 14.92 18.81 18.22 14.74 13.95 8.68 5.92 12.15 15.15 16.63
+HotpotQA 36.67 33.87 32.98 22.52 12.72 21.29 34.38 39.00 42.93 38.90 37.56 21.42 17.22 31.03 38.67 42.50
+NrtvQA 13.40 13.42 8.67 5.22 7.61 9.19 13.47 14.03 13.63 11.05 10.18 6.82 7.57 10.45 10.78 14.36
+Qasper 25.33 20.77 17.03 16.58 20.64 23.37 22.79 24.64 26.08 21.79 20.13 17.96 19.67 23.44 24.92 25.37
+SUM
+MultiNews 20.08 19.20 18.03 19.30 20.10 20.15 20.25 20.14 18.39 15.87 14.52 17.29 18.29 18.30 18.30 20.14
+GovReport 25.25 18.36 16.53 17.98 24.16 24.13 24.55 24.90 20.93 16.80 15.86 15.90 22.28 21.90 20.86 23.60
+QMSum 20.53 17.49 17.56 14.66 17.77 18.92 20.15 20.65 19.97 17.54 17.39 14.66 18.45 19.12 19.61 19.66
+FS
+TREC 46.99 46.74 45.23 46.23 41.21 46.99 45.48 47.99 65.08 64.07 63.32 62.65 60.30 56.28 64.07 65.33
+SAMSUM 17.32 16.75 16.48 12.95 17.39 17.35 17.46 18.12 16.19 13.97 13.15 11.22 15.32 15.92 15.96 17.15
+Middle
+QA
+MFQA-en 46.73 37.30 35.19 26.87 28.26 33.39 43.84 44.73 44.00 34.21 31.68 22.73 29.06 34.31 41.42 41.95
+2WikiMQA 34.10 31.71 29.90 25.59 18.77 27.26 32.50 34.25 27.80 22.32 22.69 14.19 20.18 26.15 25.40 27.01
+Musique 16.30 13.68 14.05 8.59 3.17 9.51 15.39 17.48 10.05 7.37 7.15 3.32 4.92 8.38 8.76 9.13
+HotpotQA 40.63 37.48 36.30 27.30 12.36 25.25 36.81 41.25 29.43 25.24 24.56 12.25 12.96 22.00 26.43 29.05
+NrtvQA 15.29 13.06 10.64 7.26 7.14 10.10 13.98 15.25 14.82 11.88 11.83 9.04 7.68 11.24 11.75 15.21
+Qasper 30.79 26.80 22.07 21.40 24.90 28.86 28.47 31.12 28.74 23.57 21.77 20.75 24.73 27.84 28.72 29.27
+SUM
+MultiNews 20.59 19.78 18.12 19.91 20.49 20.52 20.79 20.66 18.37 15.54 14.28 17.53 18.09 18.26 18.19 18.44
+GovReport 24.08 18.39 15.92 18.09 23.50 24.01 23.74 22.88 20.54 16.24 15.25 15.99 21.00 21.55 20.70 20.68
+QMSum 20.51 17.90 17.84 15.08 17.62 17.93 20.20 20.41 20.04 17.79 17.29 15.06 18.57 18.67 19.46 19.81
+FS
+TREC 50.00 50.00 48.74 50.25 50.00 50.76 52.27 56.03 64.07 62.06 60.81 63.32 63.82 40.71 64.07 65.33
+SAMSUM 10.62 12.03 11.83 13.34 11.10 10.92 10.62 17.43 12.38 12.55 13.28 13.83 15.65 12.95 12.93 12.40
+End
+QA
+MFQA-en 50.93 47.93 48.38 32.52 28.62 51.40 49.67 51.69 48.82 47.57 47.24 29.28 27.66 47.94 49.18 48.67
+2WikiMQA 42.43 42.34 41.96 37.20 25.19 39.54 41.75 44.05 42.70 42.25 41.24 32.89 26.50 37.54 41.34 42.24
+Musique 29.39 28.07 29.01 20.96 7.80 26.44 23.56 31.60 24.18 23.08 22.39 11.82 8.68 17.65 24.34 24.82
+HotpotQA 53.62 52.38 53.59 43.83 25.04 51.71 51.97 55.05 53.09 51.03 50.71 35.64 24.63 43.67 53.01 53.13
+NrtvQA 25.76 25.50 24.58 19.28 14.19 23.90 23.87 25.87 19.07 17.90 16.75 14.08 11.41 14.39 18.37 19.86
+Qasper 37.97 35.95 33.98 28.01 27.37 38.67 36.50 38.27 33.55 31.82 30.21 24.28 25.15 33.72 34.18 33.18
+SUM
+MultiNews 20.59 20.03 18.87 19.97 20.16 20.40 20.49 20.45 18.31 16.61 15.32 17.74 17.99 18.01 18.36 18.40
+GovReport 23.90 20.10 18.54 18.04 23.17 23.35 23.92 24.27 21.21 18.06 17.09 16.95 21.53 21.82 21.28 21.10
+QMSum 22.69 22.21 22.12 19.67 18.59 22.28 22.27 22.82 21.17 20.15 20.03 18.13 18.07 20.33 21.04 20.95
+FS
+TREC 59.05 58.54 58.54 58.54 52.51 59.55 59.05 60.31 68.34 66.08 66.59 67.84 63.82 67.59 68.85 68.34
+SAMSUM 18.78 23.88 20.63 19.84 18.45 17.75 17.62 23.53 39.46 39.58 38.77 38.71 39.13 39.20 39.57 39.85
+6.2 Main Experiments
+Effectiveness Evaluation.
+To evaluate LoopServe and baselines, we conduct experiments on the proposed 11 multi-turn long-context datasets across three tasks: QA, Summarization (SUM), and Few-shot Learning (FS).
+For each dataset, we compare LoopServe with six state-of-the-art KV cache acceleration baselines and two base LLMs, using F1, Rouge-L, or Accuracy as appropriate.
+As shown in Table 1, LoopServe achieves the best or comparable results across most datasets and query positions.
+Notably, LoopServe maintains strong performance regardless of query location, while baselines like SnapKV and AdaKV perform well only when the query is at the end.
+This highlights their reliance on positional heuristics, which limits generalization.
+In contrast, LoopServe’s adaptive approach consistently yields higher accuracy and quality, even as context length increases.
+These gains hold for both Llama-3.1 and Qwen2.5, showing LoopServe generalizes well across LLMs.
+Efficiency Evaluation.
+Beyond effectiveness, we also assess LoopServe’s generation efficiency.
+As shown in Figure 5 (a), LoopServe delivers the highest efficiency.
+This is achieved through efficient online sparsification, which selects only the most critical attention components, and adaptive KV compression, which maintains a compact, relevant cache.
+Together, these mechanisms reduce computation and memory usage, enabling fast and high-quality generation.
+Ablation Study.
+We explore LoopServe-D (progressive KV compression only) and LoopServe-P (online prefilling sparsification only) on three datasets (MF, 2WM, Qsp) using Llama and Qwen.
+As shown in Figure 5 (b) and Figure 7 in Appendix, LoopServe achieves the best performance, indicating both components are essential and complementary.
+This advantage holds across tasks, datasets, query positions, and model architectures.
+The ablation study reveals that these two components are complementary: while each addresses a different bottleneck in LLM inference, their combination ensures robust adaptation to diverse input patterns and maximizes both efficiency and accuracy.
+6.3 Parameter Sensitivity
+We analyze all hyperparameters in LoopServe: attention sparsity threshold $\alpha$ in online prefilling, token budget $B$, and decoding interval $n_{d}$ in progressive KV compression.
+Due to space limit, the analysis of the parameter $n_{d}$ is presented in Appendix A.8.1.
+Threshold $\alpha$ in  Equation 1.
+The parameter $\alpha$ controls how much total attention weight is preserved in prefilling.
+Higher $\alpha$ keeps more information but increases computation; lower $\alpha$ boosts efficiency but may lose context.
+We evaluate LoopServe on 2WikiMQA and Qasper, with Llama and Qwen backbones, across questions at the beginning (-B), middle (-M), and end (-E) positions, varying $\alpha\in\{0.980,0.985,0.990,0.995,1.00\}$.
+As shown in Figure 5 (c) and Figure 8 (a) in Appendix, LoopServe get the best accuracy and efficient for $\alpha$ between 0.99 and 1.00.
+Setting $\alpha$ too low hurts quality, while values close to 1.00 reduce efficiency gains.
+Overall, LoopServe is not overly sensitive to $\alpha$ within this range, allowing users to balance speed and quality.
+Budget $B$.
+The token selection budget $B$ in LoopServe’s progressive KV compression controls the trade-off between efficiency and output quality.
+We evaluate this on MultiFieldQA and Qasper with queries at the beginning, middle, and end.
+As shown in Figure 5 (d) and Figure 8 (b) in Appendix, increasing $B$ improves accuracy by preserving more relevant tokens, but gains are limited beyond 1024 tokens while computation and memory costs rise.
+Smaller budgets (256 or 512) reduce accuracy, especially for queries at the beginning or middle, as important tokens may be missed.
+End-position queries are less affected since key tokens are already cached.
+Overall, a budget of 1024–2048 tokens offers the best balance of performance and efficiency across all query positions.
+(a) Generation latency.
+(b) Llama-3.1-8B-Instruct.
+(c) Llama-3.1-8B-Instruct.
+(d) MultiFieldQA.
+Figure 5: Efficiency in (a), ablation study in (b), and parameter Sensitivity in (c) and (d).
+7 Conclusion
+In this paper, we propose LoopServe, an adaptive dual-phase LLM inference acceleration system designed for realistic multi-turn dialogues.
+By combining online attention sparsification and progressive KV compression, LoopServe addresses the limitations of static acceleration methods and adapts efficiently to dynamic conversational patterns.
+Our experiments on diverse, multi-turn benchmarks show that LoopServe significantly improves both inference speed and output quality compared to existing baselines, regardless of query position.
+This work provides a practical solution for efficient and effective LLM deployment in real-world dialogue scenarios.
+Appendix A Appendix
+A.1 Important Notations Table
+Table 2: Summary of important notations.
 Symbol Definition
@@
 $\hat{X}^{k}_{i}$ Selected important tokens for attention head $k$
-2.2.1. Context Compression Methods
-Context compression strategies aim to reduce effective sequence length by systematically eliminating non-essential content or distilling key information from verbose inputs.
-These approaches effectively transform lengthy inputs into more computationally manageable representations.
-Firstly, the filtering-based methods, such as LLMLingua (Jiang et al., 2023), LLMLingua-v2 (Pan et al., 2024), and CompAct (Yoon et al., 2024), selectively identify and preserve high-relevance content, enabling models to process critical information without engaging the entire sequence at maximum resolution.
-Alternatively, RAG-based techniques (Zhao et al., 2024; Jiang et al., 2024b; Wang et al., 2025; Chen et al., 2025; Edge et al., 2024) employ a two-stage process: initially constructing knowledge graphs through extraction of essential semantic triples from input text, followed by synthesizing this structured knowledge into condensed representations that serve as reduced-dimension inputs for LLM processing.
-While these compression strategies substantially reduce computational and memory requirements during inference, they inevitably sacrifice some degree of granular detail, potentially compromising the fidelity and precision of LLM-generated outputs.
-2.2.2. KV-based Optimization
-To mitigate the computational burden, KV-based approaches focus on reducing the number of attention weight calculations during both prefilling and decoding stages, as summarized in Table 2.
-Prefilling-stage Optimization.
-Given a input $X_{i}$ with $n_{i}$ length, the prefilling stage requires $O(n^{2})$ computations to compute each attention matrix across all token pairs, which is a significant computational bottleneck.
-Recent approaches such as Minference (Jiang et al., 2024a), FlexPrefill (Lai et al., 2025), and CritiPrefill (Lv et al., 2024) address this by selectively discarding negligible attention values, effectively sparsifying each attention head matrix:
-$\min\,\left|\mathbf{{A}}^{j}_{i}-\mathbf{\hat{A}}^{j}_{i}\right|,s.t.\ ,\mathbf{\hat{A}}^{j}_{i}=\textsf{Softmax}\left(\frac{\mathbf{Q}^{j}(\mathbf{K}^{j})^{\top}}{\sqrt{d_{k}}}-c(1-\mathbf{M})\right),$
-where $\mathbf{M}\in\{0,1\}^{n\times n}$ represents a binary sparsity mask, where $\mathbf{M}[a][b]=1$ indicates that the attention weight $\mathbf{A}^{j}_{i}[a][b]$ should be preserved.
-The constant $c=10^{8}$ ensures that masked positions ($\mathbf{M}[a][b]=0$) receive negligible attention weight after softmax normalization.
-This optimization reduces the computational complexity from $O(n^{2})$ to $O(\alpha n^{2})$ per head, where the sparsity factor $\alpha\ll 1$.
-Specifically, Minference (Jiang et al., 2024a) improves efficiency by selecting key vertical and slash attention patterns or important blocks, while FlexPrefill (Lai et al., 2025) enhances this approach by dynamically adjusting the sparsity rate based on attention recovery rates.
-In contrast, CritiPrefill (Lv et al., 2024) partitions keys and queries into blocks and simply selects the top-$k$ most important key blocks for computation.
-However, as shown by our motivational experiments (Section 3), the sparsity and importance patterns in attention matrices are highly dynamic and input-dependent, varying significantly across turns, heads, and layers.
-Consequently, fixed-pattern selection approaches often fail to achieve robust performance in real multi-turn scenarios.
-Decoding-stage Optimization.
-During autoregressive text generation, LLMs utilize KV caches to store and reuse partial previously computed Key and Value embeddings, thereby reducing computational redundancy.
-Firstly, token-level selection approaches (Ge et al., 2024; Li et al., 2024b), such as , H2O (Zhang et al., 2023b), SnapKV (Li et al., 2024b), and AdaKV (Feng et al., 2024), aim to select important tokens from inputs before output generation.
-These methods assume that critical input tokens required for generating outputs can be identified beforehand and that these tokens significantly overlap with those in the last observation window of the inputs.
-As analyzed in Section 3, these approaches perform well on existing benchmarks like LongBench (Bai et al., 2024a, b), because such benchmarks place user queries at the end of the input sequence.
-This assumption holds in existing benchmarks, where user queries are always placed at the end of the input sequence.
-In such settings, token-level selection methods perform well: by focusing on the most recent tokens (which include the user query), these methods can accurately preserve the necessary context for high-quality output, while significantly reducing memory and computation.
-However, this benchmark design does not match real-world multi-turn dialogues, where user queries can appear anywhere in the input.
-When queries are not always at the end, token-level selection methods like SnapKV (Li et al., 2024b) and AdaKV (Feng et al., 2024) often miss the critical tokens needed for accurate output, resulting in lower performance and poor generalization.
-As shown in Section 3.2, their effectiveness drops significantly when queries are in the middle or at the front of the context.
-This highlights the need for adaptive, context-aware KV selection methods that can handle dynamic query positions and dependencies in practical dialogue scenarios.
-2.3. Long-Context Benchmarks
-Recently, various long-context benchmarks (Li et al., 2024c, 2025; Hsieh et al., 2024; Kim et al., 2025; An et al., 2023; Qin et al., 2024) have been proposed to evaluate LLMs, such as NumericBench (Li et al., 2025), LongBench (Bai et al., 2024a, b), and LongEval (Dacheng Li* and Zhang, 2023).
-However, these benchmarks have two significant limitations that hinder their applicability to real-world scenarios.
-First, they assume that user queries are always positioned at the end of the input context.
-This architectural bias simplifies the evaluation process but fails to capture the complexity of real-world interactions, where queries may appear at the beginning, middle, or interspersed throughout the input.
-As a result, existing KV-compression approaches perform well on these benchmarks because they are optimized for scenarios where the user question is placed at the end of the input.
-However, these methods struggle to generalize to real-world human-LLM interactions.
-Second, most of these benchmarks are single-turn, focusing on isolated question-answering or summarization tasks.
-This approach overlooks the multi-turn nature of real-world conversations, where contextual dependencies across multiple turns are critical (LI et al., 2025).
-To address these limitations, we propose a multi-turn benchmark with 11 datasets in Section 4 with diverse query positions and interaction patterns, providing a more realistic evaluation of LLMs for long-context tasks.
-Table 2.
-LLM acceleration model comparisons, following (LI et al., 2025).
+Table 2 provides detailed definitions of important notations appearing in this paper.
+A.2 Summary Table of KV-based approaches
+Table 3: LLM acceleration model comparisons, following LI et al.
+(2025).
 P and D denote whether the model has optimization in the Prefilling and Decoding phases, respectively.
@@
 Methods P D KV Size Prefilling Decoding
-LLMLingua (Pan et al., 2024) $\times$ $O(\alpha n)$ $O(\alpha^{2}n^{2})$ $O(\alpha nm)$
-A-shape (Xiao et al., 2024) $\times$ $O(n)$ $O(kn)$ $O(nm)$
-Tri-shape (LI et al., 2025) $\times$ $O(n)$ $O(kn)$ $O(nm)$
-MInference (Jiang et al., 2024a) $\times$ $O(n)$ $O(kn)$ $O(nm)$
-SLLM (Xiao et al., 2024) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
-SnapKV (Li et al., 2024b) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
-AdaKV (Feng et al., 2024) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
-LoopServe $O(k)$ $O(kn)$ $O(k(m-c)+nc)$
-3. Motivational Experiments and Insights
-To better understand the fundamental challenges of accelerating LLM inference in multi-turn dialogues, we conduct a series of motivational experiments.
-These experiments, presented in Sections 3.1 and 3.2, provide key observations on the dynamic sparsity of attention patterns and the impact of query position on acceleration effectiveness.
-Based on these insights, we introduce the LoopServe system in Section 5, which is specifically designed to address these unique challenges in real-world multi-turn dialogue scenarios.
-3.1. Key Point 1: Uncertain Attention Patterns
-As discussed in Section 2, each attention head matrix is highly sparse.
-Current approaches (Jiang et al., 2024a; Xiao et al., 2024; Feng et al., 2024; LI et al., 2025; Li et al., 2024b) for accelerating the prefilling and decoding phases often rely on attention matrix sparsification and important KV token selection.
-However, these methods are built on a strong assumption: the attention sparsification in during the prefilling and important tokens in decoding follow fixed patterns, and their positions can be identified offline.
-However, the sparsity patterns of attention matrices are dynamic and unpredictable, varying across different inputs, heads, and layers.
-This variability challenges the effectiveness of fixed-pattern approaches.
-In this subsection, we first reveal the sparsity of each attention head and then demonstrate that the sparsity patterns vary across different attention heads and input sequences.
-(a) All heads of Lamma 3.
-(b) Average ratio among all heads.
-Figure 1.
-Attention head sparsity.
-Only a small fraction of vertical and slash lines in each attention head of LLMs can recover most of the attention weights, revealing strong sparsity in the attention matrix.
-3.1.1. Motivational Observation 1: only 90% of vertical lines and slash lines can collectively account for most (e.g., 90%) of the attention weight.
-In the attention matrix, a vertical line is a single column, showing that one token is attended to by all other tokens in the sequence.
-This often happens with special tokens like separators or keywords.
-A slash line is a diagonal, showing that each token mainly focuses on its nearby tokens, which is a pattern of local attention.
-We conduct experiments on a SAMSum QA (Bai et al., 2024a) dataset $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{|\mathcal{D}|}$ using the Llama-3.1-8B-Instruct model with $n_{h}$ attention heads.
-For each user query $X_{i}$ of length $n_{i}$, each $k$-th attention matrix $\mathbf{A}^{k}_{i}\in\mathbb{R}^{n_{i}\times n_{i}}$ contains $n_{i}$ slash lines $\mathcal{S}^{k}_{i}$ and $n_{i}$ vertical lines $\mathcal{V}^{k}_{i}$.
-The total attention weight for each slash line $s^{k}\in\mathcal{S}^{k}_{i}$ is computed as $\sum_{(a,b)\in s^{k}}\mathbf{A}^{k}_{i}[a][b]$, and the total attention weight for each vertical line $v^{k}\in\mathcal{V}^{k}_{i}$ is $\sum_{(a,b)\in v^{k}}\mathbf{A}^{k}_{i}[a][b]$.
-Next, we only use the top $\eta\cdot 2n_{i}$ slash lines $\mathcal{\hat{S}}^{k}_{i}\subseteq\mathcal{S}^{k}_{i}$ and vertical lines $\mathcal{\hat{V}}^{k}_{i}\subseteq\mathcal{V}^{k}_{i}$.
-Therefore, for each $k$-the attention head, the ratio of the attention weight within these selected lines to the total attention weight can be computed as follows: $r^{k}_{i}=\frac{1}{|\mathcal{D}|}\sum_{X_{i}\in\mathcal{D}}\frac{\sum_{(a,b)\in\mathcal{\hat{S}}^{k}_{i}\cup\mathcal{\hat{V}}^{k}_{i}}\mathbf{A}^{k}_{i}[a][b]}{n_{i}}.$ Also, the average ratio among $n_{h}$ attention heads can be computed as $\frac{1}{n_{h}}\sum_{j=1}^{n_{h}}r^{k}_{i}$.
-As shown in Figure 1 (a), increasing $\eta$ raises the accumulated attention weight across different heads; a small percentage of slash and vertical lines captures most of the attention weight.
-Figure 1 (b) shows that both LLaMA and Mistral exhibit this sparsity: covering 90% of the total attention weight requires selecting only 10% of slash and vertical lines.
-This indicates that, due to the sparsity in attention heads, vertical and slash lines are fundamental for efficient selection.
-3.1.2. Motivational Observation 2: The positions of the top vertical and slash lines within the same head vary across different user inputs.
-Given the LLM model $M_{\theta}$ with $n_{h}$ attention heads, the dataset $\mathcal{D}=\{X_{i}\}_{i=1}^{|\mathcal{D}|}$, and the percentage $\eta$, we select the top-$\eta\cdot 2n_{i}$ slash lines $\mathcal{\hat{S}}^{k}_{i}\subseteq\mathcal{{S}}^{k}_{i}$ and vertical lines $\mathcal{\hat{V}}^{k}_{i}\subseteq\mathcal{V}^{k}_{i}$ for each user query $X_{i}$ under each $k$-th head.
-For two queries $X_{i}$ and $X_{j}$, we compute the line overlap of the selected vertical and slash lines as: $r^{k}_{i,j}=\frac{|\mathcal{\hat{S}}_{i}^{k}\cap\mathcal{\hat{S}}_{j}^{k}|+|\mathcal{\hat{V}}_{i}^{k}\cap\mathcal{\hat{V}}_{j}^{k}|}{|\mathcal{\hat{S}}_{i}^{k}\cup\mathcal{\hat{S}}_{j}^{k}|+|\mathcal{\hat{V}}_{i}^{k}\cup\mathcal{\hat{V}}_{j}^{k}|}.$ The average important line overlap ratio across $n_{h}$ heads can then be computed as $\frac{1}{n_{h}|\mathcal{D}|^{2}}\sum_{k=1}^{n_{h}}\sum_{X_{i},X_{j}\in\mathcal{D}}r^{k}_{i,j}$.
-In particular, we use SAMSum QA (Bai et al., 2024a) dataset along with Llama-3.1-8B-Instruct and Mistral-7B-Instruct-v0.3, and vary $\eta\in\{0.1,0.15,0.2,0.25,0.3\}$.
-As shown in Figure 2, the overlap of important lines across both LLMs is mostly below 0.5 for different inputs in most heads.
-This shows that different texts produce distinct line patterns, even within the same attention head.
-Therefore, it is not practical to identify important lines offline and use them directly for online inference.
-(a) Llama-3.1-8B-Instruct.
-(b) Mistral-7B-Instruct-v0.3.
-Figure 2.
-Line overlap of the same head with different inputs.
-The important attention lines differ significantly across inputs, even for the same attention head, highlighting the dynamic and input-dependent nature of attention sparsity.
-(a) Attention map.
-(b) Important line overlap.
-Figure 3.
-Different patterns in the same head.
-It illustrates that important attention lines differ even within different segments of the same input, emphasizing the locality and unpredictability of attention patterns.
-3.1.3.
-Motivational Observation 3: Given an input $X_{i}=[C_{i}^{1},C^{2}_{i}]$, where $X_{i}$ consists of two sequential segments $C_{i}^{1}$ and $C_{i}^{2}$.
-The important positions of the top vertical and slash lines within the same head differ between $C^{1}_{i}$ and $C^{2}_{i}$.
-Both $C^{1}_{i}$ and $C^{2}_{i}$ exhibit their own local attention sparsity patterns.
-As shown in Figure 3 (a), the prominent slash and vertical lines in the attention matrix of $C^{1}$ do not appear in $C^{2}$.
-Instead, the attention part of ${C^{2}}$ exhibits distinct local patterns.
-To qualitatively verify this, we use SAMSum QA (Bai et al., 2024a) data with Llama-3.1-8B-Instruct to conduct experiments.
-Given the dataset $\mathcal{D}=\{X_{i}\}_{i=1}^{|\mathcal{D}|}$, each $X_{i}$ is split into two sequential parts as $X_{i}=[C^{1}_{i},C^{2}_{i}]$.
-We denote the attention matrix of $C^{1}_{i}$ as $\mathbf{A}^{k}_{C^{1}_{i}}\in\mathbb{R}^{|C^{1}_{i}|\times|C^{1}_{i}|}$ and the attention matrix of $C^{2}_{i}$ as $\mathbf{A}^{k}_{C^{2}_{i}}\in\mathbb{R}^{|C^{2}_{i}|\times n_{i}}$.
-Next, we select the top-$\eta$ slash and vertical lines for $\mathbf{A}^{k}_{C^{1}_{i}}$ and $\mathbf{A}^{k}_{C^{2}_{i}}$, denoted as $L^{k}_{C^{1}_{i}}$ and $L^{k}_{C^{2}_{i}}$, respectively.
-The overlap rate of importat lines between $C^{1}_{i}$ and $C^{2}_{i}$ as $r^{k}_{C_{i}^{1}\to C_{i}^{2}}={\sum_{l\in L^{k}_{C^{1}}}\mathbb{I}(l\in L^{k}_{C^{2}})}/{|L^{k}_{C^{2}}|},$ where $\mathbb{I}(l\in L^{k}_{C^{2}})=1$ if the line $l\in L^{k}_{C^{2}}$, indicating that an important line $l$ of $C_{i}^{1}$ is still important in $C_{i}^{2}$.
-Then, given data $\mathcal{D}$ and the $k$-th attention head, the average overlapping ratio is $\frac{1}{\mathcal{D}}\sum_{X_{i}\in\mathcal{D}}{r^{k}_{C_{i}^{1}\to C_{i}^{2}}}$.
-As shown in Figure 3 (b), under different $\eta$, The overlap rate between contexts in a user query remains low and highly volatile, indicating distinct local attention patterns for different subsections of input $X_{i}$.
-This observation directly highlights that it is challenging to use a subsection of the input to determine the patterns of the entire input, as distinct local patterns exist within different subsections.
-Consequently, existing prefilling acceleration approaches, such as Minference (Jiang et al., 2024a), which rely on the last window (e.g., $C^{2}_{i}$ )to determine the patterns for the whole input $X_{i}$ and select important attention blocks, fail to achieve consistent performance in real-world applications.
-Similarly, decoding acceleration approaches, such as SnapKV (Li et al., 2024b), H2O (Zhang et al., 2023b), and Keyformer (Adnan et al., 2024), which use the last observation window (e.g., last several tokens of $C_{i}^{1}$) to identify important tokens for generating outputs (e.g., the second context $C_{i}^{2}$), are also unreliable.
-As a result, these methods struggle to deliver good performance in practical scenarios.
-3.2. Key Point 2: Question Position Matters.
-Based on the above motivational experiments, both prefilling-based and decoding-phase acceleration approaches, that rely on offline sparse pattern discovery or fixed sparse patterns, should incapable of performing well in real-world use cases.
-However, the results reported in their papers are often comparable to original LLMs using the full attention matrix without any compression.
-Why?
-The reason is that existing long-context benchmarks, such as Longbench (Bai et al., 2024a, b), place each user question $q_{i}$ at the end of the input sequence $X_{i}=[C_{i},q_{i}]$.
-Then, the LLMs will answer the question $q_{i}$ based on the context $C_{i}$.
-In this setup, by relying only on the last observation window of the input (i.e., the user questions), existing acceleration approaches identify important tokens in the context that are highly correlated with the user’s specific questions.
-This partially alleviates the uncertainty caused by distinct local patterns observed in real-world scenarios.
-3.2.1.
-Motivational Observation 4: We reveal that relying solely on the last observation window of the input cannot reliably identify the important tokens from the input $X_{i}$ for generating the output $Y_{i}$.
-Existing approaches, such as H2O (Zhang et al., 2023b), SnapKV (Li et al., 2024b), and AdaKV (Feng et al., 2024), propose selecting the top-$B$ most important tokens $\hat{X}_{i}\subseteq X_{i}$ from the input $X_{i}$ before generating the output tokens.
-Given the input $X_{i}$, we denote the last observation window of $X_{i}$ with length $n_{s}$ as $X^{obs}_{i}=X_{i}[n_{i}-n_{s}:n_{i}]$.
-Under the LLM model $M_{\theta}$ with $n_{h}$ attention heads, these approaches select the top-$B$ tokens $\hat{X}_{i}\subseteq X_{i}$ where $|\hat{X}_{i}|=B$ as follows:
-(4) $\hat{X}_{i}=\arg\max_{\hat{X}_{i}\subseteq X_{i}}\sum_{j=1}^{n_{h}}\sum_{a\in\hat{X}_{i}}\sum_{b\in X^{obs}_{i}}\mathbf{A}^{k}_{i}[a][b],$
-Then, given the output $Y_{i}=[x_{n_{i}+1},x_{n_{i}+2},\cdots,x_{n_{i}+n_{o}}]$ corresponding to the input $X_{i}$, the ground-truth top-$B$ important tokens $X^{*}_{i}\subseteq X_{i}$ can be obtained as follows:
-(5) $\hat{X}^{*}_{i}=\arg\max_{\hat{X}_{i}\subseteq X_{i}}\sum_{j=1}^{n_{h}}\sum_{a\in\hat{X}_{i}}\sum_{b\in Y_{i}}\mathbf{A}^{k}_{i}[a][b],$
-The overlap between $\hat{X}_{i}$ and $\hat{X}^{*}_{i}$ can be computed as $r_{i}=\frac{|\hat{X}_{i}\cap\hat{X}^{*}_{i}|}{B}$ and the average rate regarding the data $\mathcal{D}$ is $\frac{\sum_{X_{i}\in\mathcal{D}}{r_{i}}}{|\mathcal{D}|}$.
-We set the budget $B=0.1|X_{i}|$, corresponding to 10% of the tokens in each input $X_{i}$.
-We use Llama-3.1-8B-Instruct and construct the dataset from LongEval’s topic retrieval set (Dacheng Li* and Zhang, 2023), where each instance contains three topic contexts and an instruction (framed as a question) to summarize two of them.
-For each instance $X_{i}=(C_{i},q_{i})$, we create three variants by placing the instruction or question $q_{i}$ at the beginning, middle, or end of the context $C_{i}$.
-As shown in Figure 4 (a), the average overlap of important tokens is highest when the question is at the end, but much lower when placed earlier, indicating that relying only on the last observation window misses relevant information unless the question appears last.
-(a) Important token overlapping.
-(b) Question in different position.
-Figure 4.
-Question position matters.
-When user queries are not at the end of the input, both the identification of important input tokens and the task performance of KV compression methods decline, indicating their reliance on queries being at the end.
-Also, as shown in Figure 4 (b), SnapKV (Li et al., 2024b) and AdaKV (Feng et al., 2024) perform similarly to the original LLM when the question is at the end of the input.
-However, their performance drops significantly when the question appears in the middle or at the beginning.
-It is because SnapKV (Li et al., 2024b) and AdaKV (Feng et al., 2024) rely on the last few tokens to select the most important tokens from the input.
-When the user questions are at the end of inputs, the last few tokens include the question itself, which helps identify tokens related to the question and generate high-quality outputs.
-In contrast, when the questions are placed in the beginning or middle, the last few tokens are unrelated to the questions and answers, making it difficult to identify the important tokens and leading to performance drop .
-This shows that current approaches depend heavily on the benchmark’s question placement, which does not reflect real-world input patterns.
-Consequently, they struggle to generalize to more realistic scenarios.
-4. Multi-turn Long-Context Benchmarks
-As discussed in Section 2.3, existing benchmarks mainly focus on single-turn tasks and place user queries at the end of the context, failing to reflect the complexity of real-world multi-turn dialogues.
-To address this, we propose a new multi-turn long-context benchmark that enables more comprehensive and realistic evaluation of LLMs in conversational scenarios, which is released in Huggingface.
-Table 3. Multi-turn dataset statistics.
-Type Dataset $|D|$Turn Avg Token Metric
+LLMLingua Pan et al. (2024) $\times$ $O(\alpha n)$ $O(\alpha^{2}n^{2})$ $O(\alpha nm)$
+A-shape Xiao et al. (2024) $\times$ $O(n)$ $O(kn)$ $O(nm)$
+Tri-shape LI et al. (2025) $\times$ $O(n)$ $O(kn)$ $O(nm)$
+MInference Jiang et al. (2024a) $\times$ $O(n)$ $O(kn)$ $O(nm)$
+SLLM Xiao et al. (2024) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
+SnapKV Li et al. (2024b) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
+AdaKV Feng et al. (2024) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
+LoopServe $O(k)$ $O(kn)$ O(k(mc)+nc)O(k(m-c)+nc)
+Table 3 summarizes the time complexity for each KV-based approach.
+A.3 Supplementary Figure for Motivational Observation 2
+Figure 6: Overlap rate of each head regarding different inputs of Mistral-7B-Instruct-v0.3
+Figure 7: Ablation Study of LoopServe on Qwen2.5-7B-Instruct.
+Figure 7 shows that for most heads, the overlap of Mistral-7B-Instruct-v0.3 remains below 0.5. Please refer to the detailed analysis in Section 3.1 motivational experiment 2.
+A.4 Supplementary Results for Ablation Study
+Figure 7 shows that LoopServe applied on Qwen2.5-7B-Instruct achieves the best performance, indicating the significance of both components.
+Please refer to the detailed analysis in ablation study in Section 6.2.
+A.5 Long-context Multi-turn LongBench
+Recently, various long-context benchmarks, such as NumericBench Li et al.
+(2025), LongBench Bai et al.
+(2024a), and LongEval Dacheng Li* & Zhang (2023) have been proposed to evaluate LLMs.
+However, these benchmarks have two main limitations: (1) They assume user queries always appear at the end of the input, which does not reflect real-world scenarios with queries at arbitrary positions.
+This bias favors KV-compression methods optimized for end-positioned queries, limiting their generalizability.
+(2) Most benchmarks are single-turn, overlooking the multi-turn dependencies crucial for realistic conversations LI et al.
+(2025).
+To overcome these issues, we propose a multi-turn benchmark spanning 11 datasets with diverse query positions and interaction patterns, enabling more realistic long-context LLM evaluation.
+A.5.1 The Design of Multi-turn LongBench
+We represent each $m$-turn long-context data instance in our dataset using a structured format.
+Specifically, each data instance $\mathcal{I}_{i}$ consists of $m$ turns, where each turn contains a triplet of context, question, and answer.
+The complete dataset can be formally denoted as:
+$\mathcal{D}=\{\mathcal{I}_{i}=[(C_{i,1},q_{i,1},a_{i,1}),(C_{i,2},q_{i,2},a_{i,2}),\dots,(C_{i,m},q_{i,m},a_{i,m})]\}_{i=1}^{|\mathcal{D}|},$
+where $C_{i,j}$ is the context at the $j$-th turn of the instance $\mathcal{I}_{i}$, which can be empty, $q_{i,j}$ is the corresponding user question, and $a_{i,j}$ denotes the generated answer of the LLMs for $q_{i,j}$.
+We design diverse formats for each multi-turn long-context data instance as follows:
+  • Diverse Query Positions: For the $j$-th turn, given a context $C_{i,j}=\{C^{1}_{i,j},C^{2}_{i,j},\dots,C^{p}_{i,j}\}$ consisting of $p$ distinct paragraphs (e.g., segments), the query $q_{i,j}$ can be positioned at various locations within $C_{i,j}$.
+Specifically, it can appear at the beginning of $C_{i,j}$, at the end of $C_{i,j}$, or between two segments $C^{k}_{i,j}$ and $C^{k+1}_{i,j}$.
+Such a way addresses the limitation of existing methods, which only place the query at the end of the context.
+This placement may not accurately reflect real-world scenarios.
+  • Diverse Query Relevance: At the $j$-th turn, the answer $a_{i,j}$ to question $q_{i,j}$ is derived from the contexts $\{C_{i,j^{\prime}}\}_{j^{\prime}=1}^{j}$.
+In real user scenarios, the context sources for answering $q_{i,j}$ are diverse.
+Instead of restricting $q_{i,j}$ to rely solely on $C_{i,j}$, we design the answer $a_{i,j}$ to $q_{i,j}$ to come from a subset of contexts $C_{q_{i}}\subseteq\{C_{i,j^{\prime}}\}_{j^{\prime}=1}^{j}$, with the size of the subset varying as $|C_{q_{i}}|\in\{1,2,\dots,j\}$.
+A.5.2 Multi-turn LongBench Generation
+Table 4: Multi-turn dataset statistics.
+Type Dataset $|D|$ #Turn Avg Token Metric
 QA NQA 500 3 30545.54 F1
@@
 QMSum 500 2 12780.29 Rouge-L
-Few-shot Learning TREC 199 2 2293.99 Accuracy
+Few-shot
+Learning
+TREC 199 2 2293.99 Accuracy
 SAMSUM 199 2 3113.73 Accuracy
-4.1. The Design of Multi-turn LongBench
-Each $m$-turn long-context data can be denoted as $\mathcal{D}=\{\mathcal{I}_{i}=[(C_{i,1},q_{i,1},a_{i,1}),(C_{i,2},q_{i,2},a_{i,2}),\dots,(C_{i,m},q_{i,m},a_{i,m})]\}_{i=1}^{|\mathcal{D}|}$, where $C_{i,j}$ is the context at the $j$-th turn of the instance $\mathcal{I}_{i}$, which can be empty, $q_{i,j}$ is the corresponding user question, and $a_{i,j}$ denotes the generated answer of the LLMs for $q_{i,j}$.
-We design diverse formats of for each multi-turn long-context data instance as follows:
-  • Diverse Query Positions: For the $j$-th turn, given a context $C_{i,j}=\{C^{1}_{i,j},C^{2}_{i,j},\dots,C^{p}_{i,j}\}$ consisting of $p$ distinct paragraphs (e.g., segments), the query $q_{i,j}$ can be positioned at various locations within $C_{i,j}$.
-Specifically, it can appear at the beginning of $C_{i,j}$, at the end of $C_{i,j}$, or between two segments $C^{k}_{i,j}$ and $C^{k+1}_{i,j}$.
-Such a way addresses the limitation of existing methods, which only place the query at the end of the context.
-This placement may not accurately reflect real-world scenarios.
-  • Diverse Query Relevance: At the $j$-th turn, the answer $a_{i,j}$ to question $q_{i,j}$ is derived from the contexts $\{C_{i,j^{\prime}}\}_{j^{\prime}=1}^{j}$.
-In real user scenarios, the context sources for answering $q_{i,j}$ are diverse.
-Instead of restricting $q_{i,j}$ to rely solely on $C_{i,j}$, we design the answer $a_{i,j}$ to $q_{i,j}$ to come from a subset of contexts $C_{q_{i}}\subseteq\{C_{i,j^{\prime}}\}_{j^{\prime}=1}^{j}$, with the size of the subset varying as $|C_{q_{i}}|\in\{1,2,\dots,j\}$.
-4.2. Multi-turn LongBench Generation
-Based on the above format, we design multi-turn long-context benchmarks across various categories to comprehensively evaluate the capabilities of LLMs.
-The detailed information for each dataset is presented in Table 3, and the specific methodology for constructing our proposed long-context benchmarks is outlined as follows:
+Based on the above format, we design multi-turn long-context benchmarks across various categories. Dataset details are in Table 4. Construction methodology follows:
   • Question Answering (QA).
-These datasets are derived from the single-document QA and multi-document QA tasks in LongBench (Bai et al., 2024a, b), with each dataset comprising 500 instances.
+These datasets are derived from the single-document QA and multi-document QA tasks in LongBench Bai et al.
+(2024a; b), with each dataset comprising 500 instances.
 Each instance is structured into three turns.
-To construct these instances, we randomly select three question-answer pairs from a dataset in LongBench (Bai et al., 2024a, b) as the foundation for the three turns.
+To construct these instances, we randomly select three question-answer pairs from a dataset in LongBench Bai et al.
+(2024a; b) as the foundation for the three turns.
 The associated contexts are then systematically modified through splitting and recombination, with additional irrelevant contexts incorporated.
-This meticulous design ensures that each instance satisfies the requirements for diverse query positions and diverse query relevance, as outlined in Section 4.1.
+This meticulous design ensures that each instance satisfies the requirements for diverse query positions and diverse query relevance, as outlined in Appendix A.5.1.
   • Summarization.
-These datasets are derived from the summarization tasks within LongBench (Bai et al., 2024a, b).
+These datasets are derived from the summarization tasks within LongBench Bai et al.
+(2024a; b).
 Each dataset comprises 500 instances, with each instance consisting of two turns.
 To enhance the diversity of the input, we randomly selected two instances from the original dataset and segmented their original contexts, subsequently recombining them into two turns.
-This process was carefully designed to ensure compliance with the diverse query relevance requirement outlined in Section 4.1.
+This process was carefully designed to ensure compliance with the diverse query relevance requirement outlined in Appendix A.5.1.
 Finally, we annotated the source paragraphs for traceability and introduced additional noisy contexts to further enrich the complexity and challenge of the dataset.
   • Few-shot Learning.
-These datasets are derived from the few-shot learning tasks in LongBench (Bai et al., 2024a, b).
-To fulfill the requirements of diverse query positions and diverse query relevance as outlined in Section 4.1, we exclude instances containing fewer than four examples.
+These datasets are derived from the few-shot learning tasks in LongBench Bai et al.
+(2024a; b).
+To fulfill the requirements of diverse query positions and diverse query relevance as outlined in Appendix A.5.1, we exclude instances containing fewer than four examples.
 For the remaining eligible instances, the examples are segmented and distributed across the first and second turns.
@@
 To maintain the semantic integrity and structural completeness of the examples, regular expressions are employed for segmentation.
-Input: The $m$-turn input $\mathcal{I}_{i}=\{X_{i,j}\}_{j=1}^{m}$, LLM $M_{\theta}$, threshold $\alpha$, re-selection interval $n_{d}$, and budget $B$
-Output: Answers $\{y_{i,j}\}_{j=1}^{m}$
-1 $X_{i}\leftarrow\emptyset$
-2 for $j=1$ to $m$ do
-3       $X_{i}=X_{i}\cup y_{i,j-1}\cup X_{i,j},\ \hat{X}_{i,j}=[y_{i,j-1},X_{i,j}]$
-       // Step 1: Parallel Prefilling Line Selection
-4       for $k=1$ to $n_{h}$ do
-5             $\mathcal{\hat{V}}^{k}_{i,j},\mathcal{\hat{S}}^{k}_{i,j}$ = PrefillingLineSelection($M_{\theta}$, $X_{i}$, $\hat{X}_{i,j}$, $\alpha$)
-6
-      // Step 2: KV Compression for Decoding
-7       =$\mathcal{L}=\{\mathcal{\hat{V}}^{k}_{i,j^{\prime}},\mathcal{\hat{S}}^{k}_{i,j^{\prime}}\}_{j^{\prime}=1,k=1}^{j,n_{h}}$
-8
-9      $y_{i,j}=\textsf{ProgressiveDecoding}(M_{\theta},{X}_{i},\mathcal{L},n_{d},B)$
-10
-Return  Answers $\{y_{i,j}\}_{j=1}^{m}$.
-Algorithm 1 LoopServe framework overview
-5. LoopLLM System
-In this section, we first introduce an overview of the LoopServe framework and introduce its two key components: online pre-filling sparsification and progressive KV compression for decoding.
-5.1. Framework Overview of LoopLLM System
-The LoopServe framework addresses the inefficiencies of existing LLM inference systems in multi-turn dialogue scenarios, where traditional acceleration methods based on sparse attention or static KV compression use fixed patterns that cannot adapt to dynamic input and context changes.
-These methods often fail to maintain efficiency and conversational quality when handling long contexts or varied query positions.
-LoopServe overcomes these challenges with an adaptive dual-phase system that applies online attention sparsification in the prefilling phase and progressive KV compression in the decoding phase.
-It dynamically selects the most important parts of the attention matrix to reduce computation and iteratively compresses the KV cache by using generated output tokens to select relevant input tokens.
-For an $m$-turn input $\mathcal{I}_{i}=\{X_{i,j}\}_{j=1}^{m}$, where $X_{i,j}$ is the context or query in the $j$-th turn, LoopServe generates each answer $y_{i,j}$ using these two steps.
-Step 1.
-Online Attention Head Sparsification in Prefilling.
-To reduce the quadratic complexity of the attention matrix, we dynamically select the most important slash lines and vertical lines that capture the key attention weights.
-Specifically, in Algorithm 1 line 3-5, for each new input $X_{i,j}$ and the all inputs $X_{i}=\left(\cup_{j^{\prime}=1}^{j-1}{(X_{i,j^{\prime}}\cup y_{i,j^{\prime}})}\right)\cup X_{i,j}$, we get $\hat{X}_{i,j}=[y_{i,j-1},X_{i,j}]$ as the new appended input in the $j$ turn.
-Then, for each $k$-th attention head, we first compute the attention matrix $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]\in\mathbb{R}^{\hat{n}_{i,j}\times n_{i}}$, and then select the slash lines $\mathcal{\hat{S}}^{k}_{i,j}$ and vertical lines $\mathcal{\hat{V}}^{k}_{i,j}$ that collectively recover at least $\alpha$ of the total attention weight in $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$.
-Step 2.
-Progressive KV Compression in Decoding.
-To address the high complexity of decoding, LoopServe progressively selects and stores only the most relevant tokens needed for generating subsequent outputs.
-Specifically, as described in Algorithm 1 (lines 7–14), after every re-selection interval $n_{d}$ tokens, the framework uses the ProgressiveSelection function to compute a subset of input tokens $\hat{X}^{k}_{i}\subseteq X_{i}$ for each attention head $k$.
-these selected tokens $\hat{X}^{k}_{i}$ are important for output generation.
-At each decoding step, LoopServe leverages the compressed KV cache $\{\mathcal{\hat{S}}^{k}_{i,j^{\prime}},\mathcal{\hat{V}}^{k}_{i,j^{\prime}}\}_{j^{\prime}=1,k}^{j,n_{h}}$ to generate the output sequence $y_{i,j}$.
-5.2. Online Attention Sparsification in Prefilling
-As discussed in Section 2.1, the prefilling phase incurs quadratic time complexity in the input length $n$ because each attention matrix requires $O(n^{2})$ computation.
-Sparsifying the attention matrix is therefore crucial for efficiency.
-However, as shown in Key Point 1 in Section 3.1, attention matrix sparsity patterns are highly dynamic across inputs and even within the same attention head, making static or offline selection ineffective.
-To address this, we propose an online adaptive algorithm that, during prefilling, identifies and selects a subset of slash and vertical lines to recover at least an $\alpha$ fraction of the total attention weight for each head.
-This approach not only reduces computation but also enables efficient reuse of selected lines in subsequent dialogue turns, improving overall efficiency in multi-turn scenarios.
-Definition 5.1 (Online Prefilling Sparsification Problem).
-Given the input sequence $X_{i}=\left(\bigcup_{j^{\prime}=1}^{j-1}(X_{i,j^{\prime}}\cup y_{i,j^{\prime}})\right)\cup X_{i,j}$, where $y_{i,j^{\prime}}$ is the answer for $X_{i,j^{\prime}}$ and $X_{i,j}$ is the current turn’s input, consider the LLM model $M_{\theta}$ with $n_{h}$ attention heads.
-We denote the concatenation of the previous answer $y_{i,j-1}$ and the current user input $X_{i,j}$ as $\hat{X}_{i,j}=[y_{i,j-1},X_{i,j}]$, whose corresponding attention matrix requires sparsification.
-The $k$-th attention matrix between $X_{i}$ and $\hat{X}_{i,j}$ is denoted as $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]\in\mathbb{R}^{\hat{n}_{i,j}\times n_{i}}$.
-Let $\mathcal{S}^{k}_{i,j}$ (resp., $\mathcal{V}^{k}_{i,j}$ ) denote the set of all slash lines (resp., vertical lines) in $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$.
-The goal is to select a subset of slash lines $\hat{\mathcal{S}}^{k}_{i,j}\subseteq\mathcal{S}^{k}_{i,j}$ and a subset of vertical lines $\hat{\mathcal{V}}^{k}_{i,j}\subseteq\mathcal{V}^{k}_{i,j}$ such that together they recover at least an $\alpha$ fraction of the total attention weight in $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$, where $\alpha\in[0,1]$.
-(6) $\min\sum_{s\in\mathcal{\hat{S}}^{k}_{i,j}}l_{s}+\sum_{v\in\mathcal{\hat{V}}^{k}_{i,j}}l_{v},$
-$s.t.\quad$ $\sum_{(a,b)\in(\mathcal{\hat{S}}^{k}_{i,j}\cup\mathcal{\hat{V}}^{k}_{i,j})}\mathbf{A}^{k}_{i}[a][b]\geq\alpha\cdot\hat{n}_{i,j},$
-where $\hat{n}_{i,j}$ is the total attention weight in the matrix $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$, and $l_{s}$ (resp., $l_{v}$) is the length of the slash line $s$ (resp., vertical line $v$).
-Theorem 5.2.
-The prefilling sparsification problem is NP-hard.
+A.6 Proof of Theorem 1
 Proof.
@@
 ∎
+A.7 Algorithms of LoopServe System
+Algorithm 1, Algorithm 2, and Algorithm 3 of LoopServe System in section 5 are listed below.
+1
+Input: The $m$-turn input $\mathcal{I}_{i}=\{X_{i,j}\}_{j=1}^{m}$, LLM $M_{\theta}$, threshold $\alpha$, re-selection interval $n_{d}$, and budget $B$
+Output: Answers $\{y_{i,j}\}_{j=1}^{m}$
+2 $X_{i}\leftarrow\emptyset$
+3 for $j=1$ to $m$ do
+4 $X_{i}=X_{i}\cup y_{i,j-1}\cup X_{i,j},\ \hat{X}_{i,j}=[y_{i,j-1},X_{i,j}]$
+ // Step 1: Parallel Prefilling Line Selection
+5 for $k=1$ to $n_{h}$ do
+6    $\mathcal{\hat{V}}^{k}_{i,j},\mathcal{\hat{S}}^{k}_{i,j}$ = PrefillingLineSelection($M_{\theta}$, $X_{i}$, $\hat{X}_{i,j}$, $\alpha$)
+7
+ // Step 2: KV Compression for Decoding
+8 $\mathcal{L}=\{\mathcal{\hat{V}}^{k}_{i,j^{\prime}},\mathcal{\hat{S}}^{k}_{i,j^{\prime}}\}_{j^{\prime}=1,k=1}^{j,n_{h}}$
+9
+10 $y_{i,j}=\textsf{ProgressiveDecoding}(M_{\theta},{X}_{i},\mathcal{L},n_{d},B)$
+11
+Return Answers $\{y_{i,j}\}_{j=1}^{m}$.
+Algorithm 1 LoopServe framework overview
+1
 Input: The input $X_{i}$ and $\hat{X}_{i,j}$, $k$-th head of LLM $\mathcal{M}_{\theta}$, the parameter $\alpha$
 Output: The selected slash lines $\mathcal{\hat{S}}_{i,j}^{k}$ and vectical lines $\mathcal{\hat{V}}_{i,j}^{k}$
-1
+2
 […]
-3
-4Compute Query $\mathbf{\tilde{Q}}^{k}_{i,j}$ for $\tilde{X}_{i,j}$
-5
-6Compute Key $\mathbf{{K}}^{k}_{i}$ for ${X}_{i}$
-7
+4
+5Compute Query $\mathbf{\tilde{Q}}^{k}_{i,j}$ for $\tilde{X}_{i,j}$
+6
+7Compute Key $\mathbf{{K}}^{k}_{i}$ for ${X}_{i}$
+8
 […]
-9
+10
 […]
-11
+12
 […]
-13
+14
 […]
@@
 17 while $sum<\alpha\cdot\textsf{Sum}(\mathbf{{A}}^{k}_{i}[\tilde{X}_{i,j}])$ do
-17
+18
 19 $s=\mathcal{{S}}^{k}_{i,j}[0],v=\mathcal{V}^{k}_{i,j}[0]$
-19
+20
 21 $\triangle w_{s}=w_{s}-ol_{v},\triangle w_{v}=w_{v}-ol_{s}$
@@
 24    $sum=sum+w_{s}-ol_{v}$
-24
+25
 26 else
@@
 28    $sum=sum+w_{v}-ol_{s}$
-28
 29
 30
+31
 Return $\mathcal{\hat{S}}^{k}_{i,j}$ and $\mathcal{\hat{V}}^{k}_{i,j}$.
 Algorithm 2 Adaptive Prefilling Sparsification Framework
-Algorithm.
-Algorithm 2 takes as input the concatenated sequence $\hat{X}_{i,j}=[y_{i,j-1},X_{i,j}]$, where $y_{i,j-1}$ is the previous answer and $X_{i,j}$ is the current user input.
-It also requires the $k$-th attention head of the LLM $M_{\theta}$ and a sparsity threshold parameter $\alpha$.
-The output consists of selected slash lines $\mathcal{\hat{S}}^{k}_{i,j}$ and selected vertical lines $\mathcal{\hat{V}}^{k}_{i,j}$.
-The first step involves sampling a subset $\tilde{X}_{i,j}$ of the concatenated input $\hat{X}_{i,j}$, which reduces the computational cost of subsequent operations.
-Next, Algorithm 2 computes the query matrix $\mathbf{\tilde{Q}}^{k}_{i,j}$ for the sampled input $\tilde{X}_{i,j}$ and the key matrix $\mathbf{K}^{k}_{i}$ for the entire input $X_{i}$.
-Next, all slash lines $\mathcal{S}^{k}_{i,j}$ and vertical lines $\mathcal{V}^{k}_{i,j}$ are summarized based on $\mathbf{A}^{k}_{i}[\tilde{X}_{i,j}]$ and sorted in descending order.
-Two empty sets, $\mathcal{\hat{S}}^{k}_{i,j}$ and $\mathcal{\hat{V}}^{k}_{i,j}$, are initialized to store the selected lines, and overlap weights $ol_{s}$ and $ol_{v}$ are set to zero.
-Then, Algorithm 2 iteratively selects lines until the total recovered attention weight sum meets or exceeds $\alpha\cdot\textsf{sum}(\mathbf{A}^{k}_{i}[\tilde{X}_{i,j}])$.
-At each step, the top slash line $s\in\mathcal{S}^{k}_{i,j}$ and the top vertical line $v\in\mathcal{V}^{k}_{i,j}$ are compared based on their marginal contributions, $\Delta w_{s}=w_{s}-ol_{v}$ and $\Delta w_{v}=w_{v}-ol_{s}$.
-Particularly, since one slash line only has one overlap point with one vertical line, and thus $\Delta w_{s}=w_{s}-ol_{v}$.
-The line with the greater contribution is added to its respective set ($\mathcal{\hat{S}}^{k}_{i,j}$ or $\mathcal{\hat{V}}^{k}_{i,j}$), and the overlap weights and total recovered weight are updated accordingly.
-The loop ends once the recovery condition is satisfied, and the algorithm returns the sets $\mathcal{\hat{S}}^{k}_{i,j}$ or $\mathcal{\hat{V}}^{k}_{i,j}$.
-Time Complexity Analysis.
-It is $O(n_{i}|\tilde{X}_{i,j}|+n_{i}logn_{i}+n_{i})$.
-Firstly, it takes $O(n_{i}|\tilde{X}_{i,j}|)$ to compute the partial attention matrix $\mathbf{{A}}^{k}_{i}[\tilde{X}_{i,j}]$.
-Then, it takes $O(n_{i}|\tilde{X}_{i,j}|)$ to summarize the values for each slash line and vertical line, and takes $O(n_{i}\log n_{i})$ for descending sorts.
-Finally, the greedy selection loop runs in $O(n_{i})$.
-5.3. Progressive KV Compression in Decoding
-As discussed in Section 2.1.1, conventional decoding strategies often use fixed patterns or focus on the last few input tokens to select important information.
-However, as demonstrated in Key Point 2 in Section 3.2, these methods are ineffective when user queries appear earlier in the input sequence.
-To address this, we propose a progressive KV compression strategy that dynamically selects important input tokens during decoding based on the most recently generated output tokens.
-This ensures that the selected tokens remain relevant throughout the generation process.
-Our analysis shows that using recent output tokens to guide selection leads to greater overlap with truly important input tokens, enhancing both efficiency and accuracy.
-To verify this, we consider the LLM $M_{\theta}$ with $n_{h}$ attention heads, the input sequence $X_{i}$, the generated output sequence $Y_{i}$, and a window size of $n_{w}$.
-Based on Equation (5), we compute the ground truth top-$B$ important tokens $\hat{X}^{*}_{i}\subseteq X_{i}$ using the output $Y_{i}$.
-We then use an observation window extracted from either $X_{i}$ or $Y_{i}$ to select the top-$B$ important tokens.
-Specifically, given the window size $n_{w}$, the $-\frac{|X_{i}|}{n_{w}}$-th to -1 observation windows from the input $X_{i}$ are defined as $X_{i}[n_{i}-m\cdot n_{w}:n_{i}-(m-1)\cdot n_{w}]$, $X_{i}[n_{i}-(m-1)\cdot n_{w}:n_{i}-(m-2)\cdot n_{w}]$, …, $X[n_{i}-n_{w}:n_{i}]$, and the 0-th to $\frac{|Y_{i}|}{n_{w}}$ observation windows from the output $Y_{i}$ are defined as $Y_{i}[0:n_{w}]$, $Y_{i}[n_{w}:2n_{w}]$, $Y_{i}[(m-1)\cdot n_{w}:m\cdot n_{w}]$.
-For each observation window $X^{obs}_{i}$ from $X_{i}$ or $Y_{i}$, we select the top-$B$ tokens $\hat{X}_{i}\in X_{i}$ following Equation (4) and compute the overlap rate between $\hat{X}_{i}$ and the ground truth $\hat{X}^{*}_{i}$ as $\frac{|\hat{X}_{i}\cap\hat{X}^{*}_{i}|}{B}$.
-We use the same LongEval’s topic retrieval set (Dacheng Li* and Zhang, 2023) in Section 3 and Llama to conduct experiments, and set $B=\{5\%,10\%,15\%\}\cdot|X_{i}|$.
-As shown in Figure 5 (a), the overlap rate between the important tokens selected based on input block windows (block index ¡ 0) is low.
-In contrast, the overlap rate is significantly higher when important tokens are selected based on output block windows (block index $\geq$ 0).
-It demonstrates that leveraging output tokens to identify important input tokens is a more effective strategy for selecting relevant tokens.
-(a) Important token overlapping.
-(b) Decoding blocks.
-Figure 5.
-Progressive decoding.
-(a) It shows that selecting important tokens from output windows yields higher overlap than from input windows, supporting progressive KV selection.
-(b) Also, the overlap between selected tokens in output blocks decreases with block distance, highlighting the need for progressive, output-driven cache updates.
+1
 Input: Input $\mathcal{I}_{i}$, LLM $M_{\theta}$, all selected slash and vertical lines $\{\mathcal{\hat{V}}^{k}_{i,j^{\prime}},\mathcal{\hat{S}}^{k}_{i,j^{\prime}}\}_{j^{\prime}=1,k=1}^{j,n_{h}}$, re-selection interval $n_{d}$, and budget $B$
@@
 5    ${X}^{obs}_{i}=X_{i}[|X_{i}|-n_{d}:|X_{i}|]$
-[…]
-6
+6    foreach $k=1$ to $n_{h}$ do $\hat{X}^{k}_{i}=\arg\max_{\hat{X}^{k}_{i}\subseteq X_{i},|\hat{X}^{k}_{i}|=B|}\sum_{a\in\hat{X}_{i}}\sum_{b\in X^{obs}_{i}}\mathbf{A}^{k}_{i}[a][b]$ ;
+7
 8 $n_{o}=n_{o}+1$
 9 $x_{n_{i}+n_{o}}=\textsf{Decoding}(M_{\theta},\{\hat{X}^{k}_{i}\}_{k=1}^{n_{h}},\{\mathcal{\hat{V}}^{k}_{i,j^{\prime}},\mathcal{\hat{S}}^{k}_{i,j^{\prime}}\}_{j^{\prime}=1,k=1}^{j,n_{h}})$
-9
+10
 11 $\hat{X}_{i}=\hat{X}_{i}\cup x_{n_{i}+n_{o}}$, ${X}_{i}={X}_{i}\cup x_{n_{i}+n_{o}}$
-11
+12
 13 $y_{i,j}=y_{i,j}\cup x_{n_{i}+n_{o}}$
-13
+14
 Return Answer $y_{i,j}$
 Algorithm 3 Progressive Decoding
-We also observe that the overlap of important input tokens between output blocks decreases as the distance between the blocks increases.
-Specifically, given the output sequence $Y_{i}$, we divide it into $n_{b}$ blocks, i.e., $Y_{i}=[Y^{1}_{i},\cdots,Y^{n_{b}}_{i}]$, where $Y^{j}_{i}$ denotes the $j$-th output block with a size of $\frac{|Y_{i}|}{n_{b}}$.
-For each block $Y^{j}_{i}$, we compute the top-$B$ important input tokens based on the attention weights as $\hat{X}_{i,Y^{j}_{i}}=\arg\max_{\hat{X}_{i}\subseteq X_{i}}\sum_{k=1}^{n_{h}}\sum_{a\in\hat{X}_{i}}\sum_{b\in Y^{j}_{i}}\mathbf{A}^{k}_{i}[a][b]$, Next, we compare the overlap ${|\hat{X}_{i,Y^{j}_{i}}\bigcap\hat{X}_{i,Y^{j^{\prime}}_{i}}|}/{B}$ of important input tokens between every pair of output blocks $Y^{j}_{i}$ and $Y^{j^{\prime}}_{i}$.
-As shown in Figure 5 (b), for each block $Y^{j}_{i}$, the overlap of important input tokens between $Y^{j}_{i}$ and $Y^{j^{\prime}}_{i}$ (where $j^{\prime}>j$) gradually decreases as the block distance increases.
-Additionally, the overlap between $Y^{j}_{i}$ and its immediate successor $Y^{j+1}_{i}$ is higher compared to earlier blocks, such as $Y^{j-2}_{i}$.
-This indicates that tokens identified from $Y^{j}_{i}$ are highly relevant for the generation of $Y^{j+1}_{i}$.
-Therefore, when generating the output block $Y^{j+1}_{i}$, we leverage the preceding output block $Y^{j}_{i}$ to dynamically identify the top-$B$ important input tokens required for the generation of $Y^{j+1}_{i}$.
-This progressive approach ensures that the token selection remains relevant and efficient, reducing computational overhead while maintaining high accuracy.
-Algorithm.
-Algorithm 3 firstly set the answer $y_{i,j}$ as empty and decoding step counter $n_{o}$ to 0 (line 1).
-During the decoding loop, if the current step reaches either the empirically predefined decoding size 16 or a re-selection interval $n_{d}$ (line 3), the algorithm extracts the most recent tokens from the input sequence $X_{i}$, i.e., $X^{obs}_{i}=X_{i}[|X_{i}|-n_{d}:|X_{i}|]$), and dynamically updates the compressed KV cache subset $\hat{X}^{k}_{i}$ for each $k$-th attention head (line 4-5).
-Specifically, for each head, we select top-$B$ tokens for each head as follows.
-(7) $\hat{X}^{k}_{i}=\arg\max_{\hat{X}^{k}_{i}\subseteq X_{i},|\hat{X}^{k}_{i}|=B|}\sum_{a\in\hat{X}_{i}}\sum_{b\in X^{obs}_{i}}\mathbf{A}^{k}_{i}[a][b],$
-The LLM generates the next token using the compressed KV cache and the updated input, which is appended to both the input and output sequences (line 6-9).
-This process iterates until the LLM completes generation, returning the final answer $y_{i,j}$.
-6. EXPERIMENTS
-Table 4. Effectiveness Evaluation. The bold number indicates the best performance among the baselines.
-P Data Llama 3.1 Qwen2.5
-Base Snap Ada SLLM A-S T-S Minf Ours Base Snap Ada SLLM A-S T-S Minf Ours
-Begin QA MFQA-en 45.70 36.50 34.94 25.83 28.00 34.96 43.05 46.82 44.24 35.59 33.00 21.45 29.56 38.14 42.63 43.47
-2WikiMQA 31.68 29.31 28.73 22.05 19.42 24.20 28.57 35.11 37.86 34.54 33.83 23.19 22.11 32.90 35.33 37.52
-Musique 17.16 14.29 14.80 9.76 4.22 8.18 14.92 18.81 18.22 14.74 13.95 8.68 5.92 12.15 15.15 16.63
-HotpotQA 36.67 33.87 32.98 22.52 12.72 21.29 34.38 39.00 42.93 38.90 37.56 21.42 17.22 31.03 38.67 42.50
-NrtvQA 13.40 13.42 8.67 5.22 7.61 9.19 13.47 14.03 13.63 11.05 10.18 6.82 7.57 10.45 10.78 14.36
-Qasper 25.33 20.77 17.03 16.58 20.64 23.37 22.79 24.64 26.08 21.79 20.13 17.96 19.67 23.44 24.92 25.37
-SUM MultiNews 20.08 19.20 18.03 19.30 20.10 20.15 20.25 20.14 18.39 15.87 14.52 17.29 18.29 18.30 18.30 20.14
-GovReport 25.25 18.36 16.53 17.98 24.16 24.13 24.55 24.90 20.93 16.80 15.86 15.90 22.28 21.90 20.86 23.60
-QMSum 20.53 17.49 17.56 14.66 17.77 18.92 20.15 20.65 19.97 17.54 17.39 14.66 18.45 19.12 19.61 19.66
-FS TREC 46.99 46.74 45.23 46.23 41.21 46.99 45.48 47.99 65.08 64.07 63.32 62.65 60.30 56.28 64.07 65.33
-SAMSUM 17.32 16.75 16.48 12.95 17.39 17.35 17.46 18.12 16.19 13.97 13.15 11.22 15.32 15.92 15.96 17.15
-Middle QA MFQA-en 46.73 37.30 35.19 26.87 28.26 33.39 43.84 44.73 44.00 34.21 31.68 22.73 29.06 34.31 41.42 41.95
-2WikiMQA 34.10 31.71 29.90 25.59 18.77 27.26 32.50 34.25 27.80 22.32 22.69 14.19 20.18 26.15 25.40 27.01
-Musique 16.30 13.68 14.05 8.59 3.17 9.51 15.39 17.48 10.05 7.37 7.15 3.32 4.92 8.38 8.76 9.13
-HotpotQA 40.63 37.48 36.30 27.30 12.36 25.25 36.81 41.25 29.43 25.24 24.56 12.25 12.96 22.00 26.43 29.05
-NrtvQA 15.29 13.06 10.64 7.26 7.14 10.10 13.98 15.25 14.82 11.88 11.83 9.04 7.68 11.24 11.75 15.21
-Qasper 30.79 26.80 22.07 21.40 24.90 28.86 28.47 31.12 28.74 23.57 21.77 20.75 24.73 27.84 28.72 29.27
-SUM MultiNews 20.59 19.78 18.12 19.91 20.49 20.52 20.79 20.66 18.37 15.54 14.28 17.53 18.09 18.26 18.19 18.44
-GovReport 24.08 18.39 15.92 18.09 23.50 24.01 23.74 22.88 20.54 16.24 15.25 15.99 21.00 21.55 20.70 20.68
-QMSum 20.51 17.90 17.84 15.08 17.62 17.93 20.20 20.41 20.04 17.79 17.29 15.06 18.57 18.67 19.46 19.81
-FS TREC 50.00 50.00 48.74 50.25 50.00 50.76 52.27 56.03 64.07 62.06 60.81 63.32 63.82 40.71 64.07 65.33
-SAMSUM 10.62 12.03 11.83 13.34 11.10 10.92 10.62 17.43 12.38 12.55 13.28 13.83 15.65 12.95 12.93 12.40
-End QA MFQA-en 50.93 47.93 48.38 32.52 28.62 51.40 49.67 51.69 48.82 47.57 47.24 29.28 27.66 47.94 49.18 48.67
-2WikiMQA 42.43 42.34 41.96 37.20 25.19 39.54 41.75 44.05 42.70 42.25 41.24 32.89 26.50 37.54 41.34 42.24
-Musique 29.39 28.07 29.01 20.96 7.80 26.44 23.56 31.60 24.18 23.08 22.39 11.82 8.68 17.65 24.34 24.82
-HotpotQA 53.62 52.38 53.59 43.83 25.04 51.71 51.97 55.05 53.09 51.03 50.71 35.64 24.63 43.67 53.01 53.13
-NrtvQA 25.76 25.50 24.58 19.28 14.19 23.90 23.87 25.87 19.07 17.90 16.75 14.08 11.41 14.39 18.37 19.86
-Qasper 37.97 35.95 33.98 28.01 27.37 38.67 36.50 38.27 33.55 31.82 30.21 24.28 25.15 33.72 34.18 33.18
-SUM MultiNews 20.59 20.03 18.87 19.97 20.16 20.40 20.49 20.45 18.31 16.61 15.32 17.74 17.99 18.01 18.36 18.40
-GovReport 23.90 20.10 18.54 18.04 23.17 23.35 23.92 24.27 21.21 18.06 17.09 16.95 21.53 21.82 21.28 21.10
-QMSum 22.69 22.21 22.12 19.67 18.59 22.28 22.27 22.82 21.17 20.15 20.03 18.13 18.07 20.33 21.04 20.95
-FS TREC 59.05 58.54 58.54 58.54 52.51 59.55 59.05 60.31 68.34 66.08 66.59 67.84 63.82 67.59 68.85 68.34
-SAMSUM 18.78 23.88 20.63 19.84 18.45 17.75 17.62 23.53 39.46 39.58 38.77 38.71 39.13 39.20 39.57 39.85
-We compare our LoopServe with six state-of-the-art KV cache algorithms on two representation LLM base models.
-6.1. Experimental Settings
-6.1.1. Datasets
-We designed the multi-turn long-context benchmark to rigorously evaluate LLMs in realistic conversational scenarios, addressing the limitations of single-turn benchmarks.
-Each instance comprises multiple rounds of context, queries, and answers, introducing diverse query positions and dependencies that reflect real-world conversational complexity.
-The benchmark covers three key categories: Question Answering, Summarization, and Few-shot Learning, using carefully curated datasets to ensure both diversity and complexity.
-Detailed dataset statistics are in Table 3.
-6.1.2. Baselines
-In this study, we evaluate the performance of our proposed LoopServe algorithm against six competitive and state-of-the-art KV cache algorithms on Llama-3.1-8B-Instruct (Grattafiori et al., 2024) and Qwen2.5-7B-Instruct (Team, 2024), as well as the base model without any KV cache acceleration.
-The following KV cache methods are utilized for comparison:
-  • SnapKV (Snap)  (Li et al., 2024b) optimizes KV caching through a snapshot mechanism, focusing on reducing redundant computations and enhancing memory utilization.
-  • AdaKV (Ada) (Feng et al., 2024) dynamically reallocates cache budgets across attention heads, shifting resources from heads with concentrated attention to those with dispersed patterns.
-  • StreamingLLM (SLLM)  (Xiao et al., 2024) maintains both attention sinks and a rolling window of recent tokens in its KV cache to enable stable infinite-length processing.
-  • A-shape (A-S) (Xiao et al., 2024) sparses attention that resembles an A shape.
-  • Tri-Shape (T-S)  (LI et al., 2025) creates a triangular sparse attention pattern for pre-filling.
-  • Minference (Minf) (Jiang et al., 2024a) employs a minimal inference caching strategy to reduce computational overhead while maintaining effective KV cache management.
-6.1.3. Evaluation Metrics.
-We adopt evaluation metrics for each dataset according to its task type. The metrics for each dataset are listed in Table 3. The details of the three metrics are as follows.
-  • F1: It is harmonic mean of precision and recall.
-  • Accuracy: Proportion of correct predictions over all predictions:
-[…]
-6.1.4. Hyperparameter and Hardware Setting
-All codes are executed on a Rocky Linux 8.10 machine with an 8-core Intel® Xeon® Gold 6542Y CPU, an NVIDIA H100 GPU with 80GB of memory, and 256GB of RAM.
-For baselines, we use their suggested default setting.
-For the main experiments in Section 6.2, for our LoopServe, we set $\alpha=0.955$, and $n_{d}=16$ as defaults, and we set the token selection budget $B=1024$ for all baselines and LoopServe.
-6.2. Main Experiments
-6.2.1. Effectiveness Evaluation
-To evaluate the effectiveness of our proposed LoopServe and baselines, we conduct experiments on the proposed multi-turn long-context benchmark across three categories: Question Answering, Summarization, and Few-shot Learning.
-For each dataset, we compared LoopServe with six state-of-the-art KV cache acceleration baselines as well as the base LLMs, using the appropriate metrics (F1, Rouge-L, and Accuracy) for each task.
-As shown in Table 4, LoopServe achieves the best or comparable performance among KV compression baselines across a wide range of datasets and query positions.
-A key finding is that LoopServe maintains robust performance regardless of where the user query appears in the beginning, middle, or end of the input context.
-In contrast, existing baselines such as SnapKV and AdaKV exhibit a significant drop in performance when the query is not positioned at the end, highlighting their reliance on positional heuristics which do not generalize well to realistic dialogue settings.
-LoopServe’s adaptive approach, which dynamically selects relevant information both during prefilling and decoding, consistently leads to higher accuracy and quality, even as conversational context grows longer and more complex.
-Notably, these improvements are observed across both Llama-3.1 and Qwen2.5 models, indicating that LoopServe’s design generalizes well across different LLM architectures.
-6.2.2. Efficiency Evaluation
-In addition to effectiveness, we also evaluate the efficiency of LoopServe in terms of prefilling time and decoding latency.
-As shown in Figure 6, LoopServe achieves substantial reductions in both prefilling time and inference time compared to the base LLM.
-The combination of efficient online sparsification and adaptive KV compression enables LoopServe to provide both high-quality outputs and rapid response times.
-Specifically, LoopServe’s online prefilling sparsification mechanism significantly reduces computational overhead by identifying and selecting only the most critical components of the attention matrix.
-This targeted sparsification not only accelerates the prefilling stage but also reduces the memory footprint.
-During the decoding phase, LoopServe’s progressive KV compression dynamically maintains a compact and relevant KV cache, minimizing redundant computations and further improving inference speed.
-(a) Prefilling time (time to first token). Higher is better.
-(b) Decoding latency (time between tokens). Lower is better.
-Figure 6. Efficiency experiment. Base uses naive attention, and experiences OOM at length of 16384. We use Llama-3.1-8B-Instruct for experiment.
-6.2.3. Ablation Study
-To better understand the contribution of each component within LoopServe, we conducted ablation studies by evaluating two variants: LoopServe-D, which uses only progressive KV compression, and LoopServe-P, which applies only online prefilling sparsification.
-We conduct experiments on three datasets, i.e., MultifieldQA (MF), 2WikiMQA (2WM), and Qasper (Qsp), with both Llama and Qwen backbones.
-As shown in Figure 7, LoopServe delivers the highest and comparable performance, demonstrating that the integration of both online prefilling sparsification and progressive KV compression is essential for optimal results.
-The ablation study further reveals that these two components are complementary: while each addresses a different bottleneck in LLM inference, their combination ensures robust adaptation to diverse input patterns and maximizes both efficiency and accuracy.
-These findings are stable across all task types, datasets, and query positions, as well as for both tested LLM architectures.
-6.3. Parameter Sensitivity
-To further evaluate the robustness and adaptability of LoopServe, we conducted a parameter sensitivity analysis focusing on three key hyperparameters: the attention sparsity threshold $\alpha$ used in the online prefilling phase, and the token budget $B$ and the decoding interval $n_{d}$ in the progressive KV compression phase.
-(a) Llama-3.1-8B-Instruct.
-(b) Qwen-2.5-7B-Instruct.
-Figure 7. Ablation Study.
-6.3.1. Threshold $\alpha$ in Equation (6).
-The parameter $\alpha$ determines how much total attention weight is preserved by the selected vertical and slash lines during online prefilling sparsification.
-Higher $\alpha$ retains more information but increases computation, while lower $\alpha$ improves efficiency at the risk of losing important context.
-Specifically, we conduct experiments on 2WikiMQA and Qasper data with Llama and Qwen backbones, containing questions at the beginning (-B), middle (-M), and end (-E) positions.
-We vary $\alpha$ within $\alpha\in\{0.980,0.985,0.990,0.995,1.00\}$.
-As shown in Figure 8, LoopServe achieves a favorable balance in the range of $\alpha=0.99$ to $\alpha=1.00$, where the accuracy remains high and the computational gains are significant.
-Notably, setting $\alpha$ too low can cause a drop in generation quality, while excessively high values diminish the efficiency benefits.
-These results indicate that LoopServe is not overly sensitive to the precise value of $\alpha$ within a reasonable range, and users can adjust this parameter to trade off between speed and output quality according to their application needs.
-(a) Llama-3.1-8B-Instruct.
-(b) Qwen-2.5-7B-Instruct.
-Figure 8. Parameter Sensitivity to $\alpha$
-6.3.2. Budget $B$.
-The token selection budget $B$ in LoopServe’s progressive KV compression impacts the balance between efficiency and output quality.
-Experiments conducted on MultiFieldQA and Qasper across three query positions (beginning, middle, and end).
-As shown in Table 9, increasing $B$ generally improves performance, as retaining more tokens captures a larger portion of relevant context.
-However, beyond a threshold (e.g., 1024 tokens), the performance gains diminish, while computational and memory costs increase.
-Smaller budgets (e.g., 256 or 512) result in lower accuracy, particularly for queries in the beginning or middle of the input, as key tokens may be excluded.
-In contrast, queries at the end are less sensitive to budget variations, as the most relevant tokens are naturally included in the cache.
-A budget of 1024 to 2048 tokens achieves the best balance, providing high performance across all query positions while maintaining computational efficiency.
-(a) MultiFieldQA.
+A.7.1 Time Complexity Analysis of Algorithm 1
+It is $O(n_{i}|\tilde{X}_{i,j}|+n_{i}logn_{i}+n_{i})$.
+Firstly, it takes $O(n_{i}|\tilde{X}_{i,j}|)$ to compute the partial attention matrix $\mathbf{{A}}^{k}_{i}[\tilde{X}_{i,j}]$.
+Then, it takes $O(n_{i}|\tilde{X}_{i,j}|)$ to summarize the values for each slash line and vertical line, and takes $O(n_{i}\log n_{i})$ for descending sorts.
+Finally, the greedy selection loop runs in $O(n_{i})$.
+A.8 Supplementary Experimental Figure for Parameter Sensitivity
+(a) Qwen-2.5-7B-Instruct.
 (b) Qasper.
-Figure 9. Parameter Sensitivity to $B$.
-(a) Llama-3.1-8B-Instruct.
-(b) Qwen-2.5-7B-Instruct.
-Figure 10. Parameter Sensitivity to $n_{d}$
-6.3.3. The decoding interval $n_{d}$ in Algorithm 3.
-The decoding interval $n_{d}$ determines how frequently LoopServe re-selects important input tokens based on the most recent output during progressive KV compression.
-A smaller $n_{d}$ allows the system to adapt more frequently to evolving output dependencies, which can further boost accuracy in challenging, dynamic dialogues.
-However, this comes with increased overhead due to more frequent KV cache updates.
-Specifically, we conduct experiments on MultiNews and GovReport data with Llama and Qwen backbones, containing questions at the beginning (-B), middle (-M), and end (-E) positions.
-As shown in Figure 10, our sensitivity analysis finds that moderate values of $n_{d}$ (e.g., 16 or 32) provide the best balance: they preserve the majority of the efficiency gains while maintaining robust generation quality.
-Extremely large $n_{d}$ values reduce the adaptivity of the KV cache and may result in decreased relevance of selected tokens, leading to lower performance on complex, multi-turn tasks.
-7. Conclusion
-In this paper, we propose LoopServe, an adaptive dual-phase LLM inference acceleration system designed for realistic multi-turn dialogues.
-By combining online attention sparsification and progressive KV compression, LoopServe addresses the limitations of static acceleration methods and adapts efficiently to dynamic conversational patterns.
-Our experiments on diverse, multi-turn benchmarks show that LoopServe significantly improves both inference speed and output quality compared to existing baselines, regardless of query position.
-This work provides a practical solution for efficient and effective LLM deployment in real-world dialogue scenarios.
+(c) Llama-3.1-8B-Instruct.
+(d) Qwen-2.5-7B-Instruct.
+Figure 8: Parameter sensitivity to $\alpha$ in (a), budget $B$ in (b), and decoding interval $n_{d}$ in (c) and (d).
+A.8.1 The decoding interval $n_{d}$ in Algorithm 3
+The decoding interval $n_{d}$ controls how often LoopServe re-selects important input tokens during progressive KV compression.
+Smaller $n_{d}$ values enable frequent adaptation to changing output dependencies, improving accuracy in dynamic dialogues but increasing overhead from more KV cache updates.
+We evaluate this on MultiNews and GovReport with Llama-3.1-8B-Instruct.
+and Qwen-2.5-7B-Instruct., covering questions at the beginning, middle, and end.
+As shown in Figure 8 (c) and (d), moderate $n_{d}$ values (e.g., 16 or 32) strike the best balance, maintaining efficiency and robust generation quality.
+Very large $n_{d}$ reduces adaptivity, leading to lower performance on complex, multi-turn tasks.
+A.9 The Usage of LLMs for Paper Writing
+We use GPT-4o and DeepSeek-R1 to polish our paper.